### v1

[…]
1 Introduction
Cannings models, first introduced in [cannings74, cannings1975], form a class of discrete-time population models in which each generation consists of a fixed number of individuals, and reproduction is governed by an exchangeable offspring distribution. This framework includes many classical models such as the Wright–Fisher model and (discrete-time) Moran model, and has played a central role in mathematical population genetics.
In this paper, we study the genealogical trees arising from Cannings models, referred to as Cannings trees. While the scaling limits of various genealogical trees—most notably Galton–Watson trees—have been extensively studied (e.g., [1993Aldous, marckert2003depth, le2006random]), the scaling behavior of Cannings trees remains largely unexplored. This motivates our present work.
To obtain a richer class of limiting results, we extend the classical Cannings framework to a class of inhomogeneous Cannings models, in which the population size may vary across generations. We assume that the rescaled generation-size profile converge to a continuous function $\ell$. Our main result (Theorem 1.1) shows that, under suitable moment assumptions, the rescaled height and contour functions of Cannings trees converge to a time change of Brownian motion conditioned to have local time profile $\ell$.
1.1 Setup and main results
Inhomogeneous Cannings model
An inhomogeneous Cannings model describes the evolution of a finite population over discrete generations $s\in\mathbb{N}_{+}=\{1,2,\ldots\}$, with population sizes across generations given by a sequence $q=\big(q(s)\big)_{s\in\mathbb{N}_{+}}$ taking values in $\mathbb{N}=\{0,1,\ldots\}$. We assume that $q$ satisfies the following extinction closure condition:
$q(s)=0\quad\text{for all }s\geq h(q),\quad\text{where }h(q):=\inf\big\{i\in\mathbb{N}_{+}:q(i)=0\big\}\in\mathbb{N}\cup\{\infty\}.$
We refer to such a sequence $q$ as a discrete profile.
The offspring distribution in generation $s$ is described by a random vector
$\nu^{s}=\big(\nu_{1}^{s},\ldots,\nu_{q(s)}^{s}\big),$
where $\nu_{i}^{s}\in\mathbb{N}$ denotes the number of offspring produced by the $i$-th individual in generation $s$. The sequence $(\nu^{s})_{s\in\mathbb{N}_{+}}$ is assumed to satisfy the following conditions:
  • Mass conservation:
    $\sum_{i=1}^{q(s)}\nu_{i}^{s}=q(s+1);$
  • Exchangeability: for all permutations $\pi\in S_{q(s)}$,
    $(\nu_{1}^{s},\ldots,\nu_{q(s)}^{s})\overset{d}{=}(\nu_{\pi(1)}^{s},\ldots,\nu_{\pi(q(s))}^{s});$
  • Generation-wise independence: the vectors $\nu^{s}$ are independent across $s$.
When $q(s)\equiv N$ is constant and $(\nu^{s})_{s\in\mathbb{N}}$ are i.i.d., the model reduces to the classical Cannings model. Important examples of the inhomogeneous case include the inhomogeneous Wright–Fisher model and the Galton–Watson model conditioned on generation-size profile.
The genealogical structure induced by an inhomogeneous Cannings model can be naturally represented as a random rooted ordered tree, which we call a Cannings tree. A formal definition is given in Section 2.1.
Brownian motion conditioned on local time profile
Let $W$ be a reflected Brownian motion on $\mathbb{R}^{+}=[0,\infty)$, and denote by $L_{t}(x)$ its local time at level $x$ up to time $t$. Define the inverse local time at the origin by
$\tau_{u}:=\inf\{t\geq 0:L_{t}(0)>u\}$
A function $\ell:\mathbb{R}^{+}\to\mathbb{R}^{+}$ is called a continuous profile if it satisfies:
  • $\ell(x)=0$ for all $x\geq h(\ell):=\inf\{y>0:\ell(y)=0\}$;
  • $\ell$ is continuous on $0,h(\ell))$ and admits a finite left limit at $h(\ell)$ if $h(\ell)<\infty$.
(Note that $\ell$ may be discontinuous at $h(\ell)$.)
Given a continuous profile $\ell$, we define $W^{\ell}$ as the process $W$ up to time $\tau_{\ell(0)}$, conditioned on the local time profile satisfying
$L_{\tau_{\ell(0)}}(x)=\ell(x)\quad\text{for all }x\geq 0.$
This conditioned process was constructed explicitly by Warren–Yor [warren98] and characterized via continuum random trees by Aldous [1998BROWNIAN]. In Section 2.3, we recall these constructions and introduce a canonical version of $W^{\ell}$. Throughout the paper, we always work with that canonical version.
Convergence assumptions and results
We now consider a sequence of inhomogeneous Cannings models indexed by $n$, with generation-size profiles $q_{n}=(q_{n}(s))_{s\in\mathbb{N}_{+}}$ and offspring vectors $(\nu^{n,s})_{s\in\mathbb{N}_{+}}$. For $n,s\in\mathbb{N}_{+}$, set
$\sigma_{n}(s)^{2}:=\mathbf{1}\{s<h(q_{n})\}\,\mathrm{Var}(\nu_{1}^{n,s}).$
The assumptions for our main result are as follows:
  1. (H1)
    There exist continuous profiles $\ell$ and $\sigma$ with $h(\ell)=h(\sigma)<\infty$ such that
    1. (i)
      As $n\to\infty$,
      $\frac{h(q_{n})}{n}\to h(\ell),\qquad\frac{q_{n}(\lfloor n\,\cdot\rfloor)}{n}\to\ell\ \text{uniformly on compact subsets of }(0,h(\ell)),$
      and
[…]
    2. (ii)
      $\lim\limits_{\delta\rightarrow 0}\sup\limits_{n\in\mathbb{N}_{+}}\dfrac{1}{n^{2}}\sum\limits_{s\in\mathbb{N}_{+}\setminus[\delta n,h(q_{n})-\delta n]}q_{n}(s)=0$;
    3. (iii)
      $\lim\limits_{s\rightarrow 0+}\dfrac{\ell(s)}{\sigma(s)^{2}}$ exists.
  2. (H2)
    One of the following conditions holds:
    1. (i)
      For every $\delta>0$, there exists $\varepsilon>0$ such that uniformly for $\delta n\leq s\leq h(q_{n})-\delta n$,
      $\mathbb{E}\big[(\nu_{1}^{n,s})^{3}\big]=o\!\left(\frac{n}{(\log n)^{2+\varepsilon}}\right);$
    2. (ii)
      For every $\delta>0$, the family $\big\{\nu^{n,s}_{1}:n\geq 1,\ \delta n\leq s\leq h(q_{n})-\delta n\big\}$ is uniformly $L^{2}$-integrable, i.e.
      $\lim_{K\to\infty}\ \sup_{n\geq 1}\ \sup_{\delta n\leq s\leq h(q_{n})-\delta n}\mathbb{E}\big[(\nu_{1}^{n,s})^{2}\,\mathbf{1}\{\nu^{n,s}_{1}>K\}\big]=0.$ (1)
Remark. Condition H1(ii) in (H1) ensures that the contribution to the rescaled population size from the very early and very late generations—near the beginning and near extinction—is negligible. Combined with H1(i), this yields the convergence of the rescaled total population size:
$\frac{1}{n^{2}}\sum_{s\in\mathbb{N}_{+}}q_{n}(s)\;\longrightarrow\;\int_{0}^{\infty}\ell(s)\,\mathrm{d}s.$ (2)
Condition H1(iii) in (H1) ensures that the function
$\ell^{\sigma}(s):=\mathbf{1}\{s<h(\ell)\}\,\frac{4\ell(s)}{\sigma(s)^{2}}$
extends continuously to $s=0$, and hence defines a continuous profile. As will be seen in Theorem 1.1, the limiting object involves Brownian motion conditioned on the local time profile $\ell^{\sigma}$. Hence condition H1(iii) rules out the pathological case we would otherwise have to handle—conditioning on a profile whose value at the origin fails to converge.
Let $\mathscr{T}^{n}$ denote the Cannings tree associated with the model indexed by $n$, and let $C_{\cdot}(\mathscr{T}^{n})$ and $H_{\cdot}(\mathscr{T}^{n})$ be its contour and height functions, respectively (see Section 2.2.1 for precise definitions). Our main result is as follows.
Theorem 1.1.
Assume (H1)–(H2). Define $\alpha(t):=\frac{1}{4}\int_{0}^{t}\sigma(W^{\ell^{\sigma}}_{s})^{2}\,\mathrm{d}s$. Let $\alpha^{-1}(t)$ be the inverse of $\alpha(t)$ [^4] We will verify in Lemma 2.16 that $\alpha$ is a.s. strictly increasing, so $\alpha^{-1}(t)$ is well-defined.. Then
$\Big(n^{-1}C_{\lfloor 2n^{2}t_{1}\rfloor}(\mathscr{T}^{n}),n^{-1}H_{\lfloor n^{2}t_{2}\rfloor}(\mathscr{T}^{n})\Big)_{0\leq t_{1},t_{2}<\infty}\overset{d}{\longrightarrow}\Big(W_{\alpha^{-1}(t_{1})}^{\ell^{\sigma}},W^{\ell^{\sigma}}_{\alpha^{-1}(t_{2})}\Big)_{0\leq t_{1},t_{2}<\infty}$ (3)
in the product space $\big(D[0,\infty)\big)^{2}$ equipped with the product Skorokhod topology (where the functions on both sides are extended by 0 outside their original domains).
It is a standard fact that convergence of contour functions implies convergence of the associated trees in the Gromov-Hausdorff topology (see [le2006random, Lemma 2.4]). Applying this, we obtain the following corollary.
Corollary 1.2.
Assume (H1)–(H2). Let $\widetilde{\mathscr{T}}^{n}$ be the geometric tree obtained from $\mathscr{T}^{n}$ by assigning length $1/n$ to each edge. Let $\mathcal{T}^{\ell,\sigma}$ denote the continuum random tree encoded by the time-changed conditioned Brownian motion $W^{\ell^{\sigma}}_{\alpha^{-1}(t)}$. Then
$\widetilde{\mathscr{T}}^{n}\overset{d}{\longrightarrow}\mathcal{T}^{\ell,\sigma}\quad\text{under the Gromov-Hausdorff topology.}$
(See Section 2.2 below for precise definitions of geometric trees and tree encodings.)
Remark 1.3.
Condition (H2) is a sufficient but not optimal assumption for the convergence result (3). A sharper condition consists of (H2’) (stated in Section 3.3) together with
$\lim_{n\rightarrow\infty}\ \sup_{\delta n\leq s\leq h(q_{n})-\delta n}\frac{\mathbb{E}\big[(\nu_{1}^{n,s})^{3}\big]}{n}=0.$ (H2”)
The corresponding convergence result under (H2’) and (H2”) is given in Theorem 5.6. We show in Proposition 5.8 that the condition (H2) implies both (H2’) and (H2”).
By Aldous’ theorem [1993Aldous, Theorem 20], the convergence of the contour functions is equivalent to the combination of tightness and finite-dimensional convergence of the associated tree sequence. The condition (H2”) is in fact equivalent to the finite-dimensional convergence (see Remark 5.3); while condition (H2’) is closely related to the tightness. In Section B, we present a counterexample in which the offspring distribution slightly violates (H2’), leading to a failure of tightness. This indicates a form of sharpness of condition (H2’).
We also remark that the counterexample in Section B shows that the condition (H2”), although sufficient for finite-dimensional convergence, is not enough to guarantee tightness.
Remark 1.4.
The convergence results can also be extended to other population models, such as the continuous-time Moran model and the diploid Cannings model (as defined in [mohle2003coalescent]), using a very similar approach.
1.2 Related works
Coalescent theory
A major focus in the study of Cannings models lies in coalescent theory, which describes the genealogies of finite samples via limiting processes such as Kingman’s coalescent [kingman1982genealogy]. As explained in Remark 5.3, convergence to Kingman’s coalescent is in fact equivalent to the finite-dimensional convergence of the associated tree sequence.
In [mohle2000total, Theorem 4], Möhle gave the necessary and sufficient condition for convergence to Kingman’s coalescent in the classical Cannings model. Condition (H2”) serves as the analogue of this criterion in the inhomogeneous Cannings setting. Since it is equivalent to the finite-dimensional convergence, it constitutes a necessary condition of the convergence result (3), as discussed in Remark 1.3.
Encoding tree of conditioned Brownian motion
It is classical that every continuous bridge function naturally encodes a continuum random tree (see Section 2.2.2 for details). For the conditioned Brownian motion $W^{\ell}$, Aldous [1998BROWNIAN] derived the finite-dimensional distributions of the corresponding encoding tree. Subsequently, Warren [warren1999result] provided two alternative, direct proofs of this result. This distribution plays a key role in identifying the limiting process in our setting.
Processes conditioned on local time profile
The conditioned Brownian motion $W^{\ell}$ is a self-interacting diffusion constructed by Warren and Yor [warren98], and independently by Aldous [1998BROWNIAN]. Detailed constructions will be presented in Section 2.3. Subsequently, Warren [warren05] further established a close connection between this conditioning problem and a stochastic flow of the Bass-Burdzy type.
More recently, Lupu, Sabot, and Tarrès [lupu2021inverting], as well as Aïdékon, Hu, and Shi [aidekon2023stochastic], have further developed this framework by analyzing perturbed reflecting Brownian motions and Brownian loop soups conditioned on local time profiles [^5] Specifically, Lupu, Sabot, and Tarrès [lupu2021inverting] studied the Gaussian free field conditioned on local time profile, which is equivalent—via Dynkin’s isomorphism—to a Brownian loop soup of intensity $1/2$, or equivalently, a perturbed reflecting Brownian motion with parameter $2$, conditioned on local time profile. Aïdékon, Hu, and Shi [aidekon2023stochastic] later generalized this result to arbitrary loop soup intensities (or equivalently, arbitrary parameters for perturbed reflecting Brownian motion).. A natural extension of our setting is to incorporate immigration into the inhomogeneous Cannings model. In that case, the rescaled contour and height processes will converge to the (time-changed) perturbed reflecting Brownian motion conditioned on its local time profile.
On the discrete side, the conditional laws of Markov chains and Markovian loop soups given their local time profiles have been studied in [2016Inverting, lupu2019inverting, li2024inverting].
1.3 Proof outline
Our approach to proving Theorem 1.1 builds on the framework developed by Aldous in [1993Aldous, Theorem 20], where he showed that for a sequence of random rooted ordered trees, convergence of the contour process is equivalent to the combination of finite-dimensional convergence and tightness of the trees (see Theorem 2.6 for a modified version of this result adapted to our setting). In our case, finite-dimensional convergence is obtained via its equivalence with convergence to Kingman’s coalescent (see Proposition 5.1). The main technical challenge lies in proving tightness. Once tightness is established, the convergence of the contour process follows, and the joint convergence with the height process is then deduced using a deterministic relation between the two processes from [marckert2003depth] (see Section 5.2).
We now outline the key ideas used in the proof of tightness. Roughly speaking, tightness of a sequence of trees means that for any $\varepsilon,\delta>0$, there exists a positive integer $k$ such that for every tree in the sequence, with probability at least $1-\varepsilon$, a uniformly chosen set of $k$ vertices forms a $\delta$-net of the tree—that is, every vertex lies within distance $\delta$ of at least one of the selected vertices. In the context of the (unrescaled) Cannings trees $\mathscr{T}^{n}$, this corresponds to a $\lfloor\delta n\rfloor$-net formed by $k$ uniformly chosen vertices.
To this end, we analyze the ancestral structure of the tree from top to bottom. We partition the height of the tree into intervals of length $\lfloor\delta n\rfloor$. For each such interval $\big[s\lfloor\delta n\rfloor,(s+1)\lfloor\delta n\rfloor\big)$, we consider the ancestors at height $s\lfloor\delta n\rfloor$ of all vertices located at height $(s+1)\lfloor\delta n\rfloor$. These ancestral vertices are referred to as the $\delta$-coalescent vertices at the $s$-th height interval; see Definition 4.2 for the precise definition.
The union of all $\delta$-coalescent vertices forms a $2\lfloor\delta n\rfloor$-net of the tree. Tightness then follows by showing that, with high probability, these vertices lie within distance $\lfloor\delta n\rfloor$ of the $k$ uniformly chosen vertices. This relies on two key facts, which hold with high probability:
  1. 1.
    The total number of $\delta$-coalescent vertices is bounded by a universal constant (Proposition 4.3);
  2. 2.
    Each $\delta$-coalescent vertex has offspring of size $\Theta(n^{2})$ within distance $\lfloor\delta n\rfloor$ (Proposition 4.5).
We highlight that the first fact relies on a sequential version of the coming-down-from-infinity property, established in Theorem 3.26. To this end, we introduce a coalescent process, a Markov process that tracks the number of ancestral lineages backward from height $(s+1)\lfloor\delta n\rfloor$. In particular, at height $s\lfloor\delta n\rfloor$ this count is exactly the number of $\delta$-coalescent vertices at the $s$-th height interval. Our assumption (H2’) is used in the proof to control the merge rates of these lineages. Heuristically, under (H2’) the process contracts at a uniform multiplicative rate on mesoscopic windows: on average, over each window of length $\frac{n}{(\log n)^{1+\varepsilon}},$ the lineage count shrinks by a factor $(\log n)^{-c}$. So over a time horizon $\delta n$, the cumulative shrinkage is
$(\log n)^{-c\delta(\log n)^{1+\varepsilon}}=\exp\!\big(-c\delta(\log n)^{1+\varepsilon}\log\log n\big)\ll 1/n.$
This results in a rapid decrease in the number of lineages—from $\Theta(n)$ down to a constant—within $\delta n$ steps. By contrast, in the counterexample of Section B (where (H2’) fails), a contraction occurs only every $\frac{n}{(\log n)^{1-\varepsilon}}$ steps; over the same horizon $\delta n$ the cumulative shrinkage is
$(\log n)^{-c^{\prime}\delta(\log n)^{1-\varepsilon}}=\exp\!\big(-c^{\prime}\delta(\log n)^{1-\varepsilon}\log\log n\big)\gg 1/n,$
which is insufficient to reduce the lineage count from order $\Theta(n)$ to $O(1)$. This leads to non-convergence of the rescaled contour functions in this example (see Proposition B.1).
For the second fact, we analyze the ancestral structure of all vertices at height $(s+1)\lfloor\delta n\rfloor$. We show that, with high probability, there exists a stretch of height—around $(s+\tfrac{1}{2})\lfloor\delta n\rfloor$ and of length $\Theta(n)$—during which the number of their ancestors remains constant and is bounded by a constant independent of $n$. This stability implies that a large proportion of other vertices at these heights are likely to coalesce onto these few persistent ancestral lines within this time window. These vertices already account for $\Theta(n^{2})$ descendants of the $\delta$-coalescent vertices at the $s$-th height interval. See Sections 4.34.4 for details.
1.4 Organization of the paper
The remainder of the paper is organized as follows. In Section 2, we present several important preliminaries, including the formal definition of Cannings trees, Aldous’ theorem on the convergence of contour functions, and the construction of the conditioned Brownian motion $W^{\ell}$. Section 3 introduces the coalescent process and derives some of its basic properties, including the coming-down-from-infinity property. Section 4 is devoted to proving the tightness of the rescaled trees. Finally, in Section 5, we complete the proof of the main theorem by establishing the finite-dimensional convergence and the joint convergence of the contour and height functions. Appendix A proves convergence to Kingman’s coalescent under (H2”), and Appendix B provides a counterexample demonstrating the near-sharpness of (H2’).
As a brief guide to the assumptions, our proof of Theorem 1.1 proceeds in two steps. First, under (H1), (H2’) and (H2”), we prove the convergence stated in Theorem 1.1. Second, we show that (H2) implies both (H2’) and (H2”). Accordingly, most results in Sections 35 are stated under (H2’) and (H2”), rather than (H2).
2 Preliminaries
2.1 Cannings tree
Let
$\mathcal{U}:=\bigcup_{n\geq 0}(\mathbb{N}_{+})^{n},$
where by convention $(\mathbb{N}_{+})^{0}:=\{\emptyset\}$. Every rooted ordered tree can be canonically represented as a prefix-stable finite subset of $\mathcal{U}$. See, for example, [le2006random, Section 1.1] for details.
Using this formalism, we now give a precise definition of the Cannings tree associated with an inhomogeneous Cannings model. For a sequence $u=(u_{1},\ldots,u_{m})\in\mathcal{U}$, we write $|u|=m$ for its length. For $u=(u_{1},\ldots,u_{m})$ and $j\in\mathbb{N}_{+}$, we denote by $uj$ the concatenated sequence $(u_{1},\ldots,u_{m},j)$. For $u,v\in\mathcal{U}$, we write $u\prec v$ if $u$ precedes $v$ in the lexicographical order on $\mathcal{U}$.
Definition 2.1 (Cannings tree).
Given a population profile $q=(q(s))_{s\in\mathbb{N}_{+}}$ and offspring vectors $\nu^{s}=(\nu_{1}^{s},\ldots,\nu_{q(s)}^{s})$, we define the associated Cannings tree $\mathscr{T}$ as a subset of $\mathcal{U}$ constructed recursively as follows:
  • $\emptyset\in\mathscr{T}$;
  • Suppose all elements $u\in\mathscr{T}$ with $|u|=s$ have been defined. Label them in lexicographic order as $u^{1}\prec u^{2}\prec\cdots\prec u^{q(s)}$. Then the elements $u\in\mathscr{T}$ with $|u|=s+1$ are defined as
    $\left\{u^{i}j:1\leq i\leq q(s),\,1\leq j\leq\nu^{s}_{i}\right\},$
    where we set $q(0):=1$ and $\nu^{0}_{1}:=q(1)$.
The resulting set $\mathscr{T}$ is a prefix-stable subset of $\mathcal{U}$ representing a random rooted ordered tree.
Intuitively, this construction starts by introducing an artificial ancestor, the root node $\emptyset$, which serves as the common parent of all individuals in generation $1$. Each vertex at height $s$ (i.e., with label length $s$) represents an individual in generation $s$, and the corresponding entry $\nu^{s}_{i}$ in the offspring vector determines the number of children that the $i$-th individual (in lexicographic order) in generation $s$ has. These children are then represented by appending $1,2,\ldots,\nu^{s}_{i}$ to the parent’s representation.
In this way, the tree structure encodes the entire genealogy of the population.
2.2 Foundations of tree encoding and convergence
This section introduces the framework for encoding trees as height and contour functions, and then provides necessary and sufficient conditions for the convergence of contour functions.
2.2.1 Geometric trees
In this section, we extend the notion of rooted ordered trees by assigning a positive length to each edge, thereby introducing the concept of a rooted ordered geometric tree. This enriched structure is better suited for studying convergence, where edge lengths need to be rescaled.
A rooted ordered geometric tree is defined as a pair $(T,w)$, where:
  • $T$ is a rooted ordered tree with vertex set $V(T)$ and edge set $E(T)$;
  • $w:E(T)\to(0,\infty)$ is a function that assigns a positive length $w(e)$ to each edge $e\in E(T)$.
We denote the root of $T$ by $\rho$. The function $w$ naturally induces a metric on $V(T)$. For any two vertices $v,v^{\prime}\in V(T)$, the distance is given by
$d(v,v^{\prime}):=\sum_{e\in[\![v,v^{\prime}]\!]}w(e),$
where $v,v^{\prime}]\!]$ denotes the unique path connecting $v$ and $v^{\prime}$ in $T$. In particular,
$\left\lvert v\right\rvert:=d(\rho,v)$
represents the height (or depth) of $v$ measured in terms of edge lengths.
Height function
The height function $H(T)$ of a rooted ordered tree $(T,w)$ records the height of each vertex according to the lexicographic order on $T$. Specifically, if
$V(T)=\{v_{0},v_{1},\dots,v_{k-1}\}\quad\text{with}\quad\rho=v_{0}\prec v_{1}\prec\cdots\prec v_{k-1},$
then the height function is defined by
$H_{n}(T):=\left\lvert v_{n}\right\rvert,\quad 0\leq n<k.$
Contour function
The contour function $C(T)$ encodes the tree via a depth-first traversal that respects the order on $T$. Imagine a particle starting at the root $\rho$ and exploring $T$ “from left to right” according to this order. The particle takes one time unit to traverse each edge, and since each edge is crossed twice (once forward and once backward), the total number of steps in the traversal is 2(|V(T)|1)2(|V(T)|-1), where $|V(T)|$ is the number of vertices in $V(T)$. For each integer $n\in\{0,1,\dots,2(|V(T)|-1)\}$, let $f(n)$ denote the vertex visited at step $n$. Then the contour function is defined by
$C_{n}(T):=\left\lvert f(n)\right\rvert,\quad 0\leq n\leq 2(|V(T)|-1).$
Remark 2.2.
A rooted ordered tree can be viewed as a special case of a rooted ordered metric tree in which all edge lengths are equal to one (i.e., $w(e)\equiv 1$ for all $e\in E(T)$). In this case, the classical definitions of height and contour functions, typically defined by the number of edges, are recovered.
2.2.2 Encoding compact real trees by bridge functions
Following Le Gall’s survey [jean2005random], compact real trees are defined as compact metric spaces satisfying the standard tree properties (see [jean2005random, Section 2.1] for details). A canonical way to construct such trees, which simultaneously endows them with a root and an ordered structure, is through encoding functions drawn from the space
$C_{\rm bridge}:=\{g\in C[0,1]:g(0)=g(1)=0,\ g(x)\geq 0\text{ for }x\in[0,1]\}.$
Specifically, given $g\in C_{\rm bridge}$, we define a pseudo-metric:
$d_{g}(s,t):=g(s)+g(t)-2\inf_{r\in[s\wedge t,s\vee t]}g(r),\quad\text{for }s,t\geq 0.$
This induces an equivalence relation:
$s\sim t\quad\text{if and only if}\quad d_{g}(s,t)=0.$
The quotient space $\mathcal{T}_{g}:=[0,1]/\sim$, equipped with the inherited metric $d_{g}$, is a compact real tree by [jean2005random, Theorem 2.2]. Let $p_{g}:[0,1]\to\mathcal{T}_{g}$ be the canonical projection. We regard $\mathcal{T}_{g}$ as rooted at $p_{g}(0)$ and equip it with an order by declaring $s_{1}\prec_{g}s_{2}$ if and only if
$p_{g}^{-1}(s_{1})<p_{g}^{-1}(s_{2}),\qquad p_{g}^{-1}(s):=\min\{t\in[0,1]:p_{g}(t)=s\}.$
Intuitively, this construction can be understood by visualizing the graph of $g$ and identifying points that are connected through horizontal compression. This process “flattens” the graph into a tree structure.
Remark 2.3.
Le Gall’s encoding of compact real trees uses continuous functions with compact support [jean2005random]. Our encoding through $C_{\rm bridge}$ functions is essentially equivalent because, in Le Gall’s setting, the function $g(x)$ and its rescaled version $g(cx)$ encode the same real tree. Thus, without loss of generality, we can always assume $g\in C_{\rm bridge}$.
We now consider a random rooted ordered real tree $(\mathcal{T}_{g},d_{g},\preceq_{g})$ encoded by a random function $g\in C_{\rm bridge}$. Its finite-dimensional distributions are captured by the uniform $k$-point subtrees defined below.
For any $t_{1},\ldots,t_{k}\in[0,1]$, let $\mathcal{T}_{g}(t_{1},\ldots,t_{k})$ be the subtree of $\mathcal{T}_{g}$ generated by the points $\{p_{g}(t_{1}),\ldots,p_{g}(t_{k})\}$:
$\mathcal{T}_{g}(t_{1},\ldots,t_{k})=\bigcup_{i=1}^{k}[\![\rho,\,p_{g}(t_{i})]\!],$
where $x,y]\!]\subset\mathcal{T}_{g}$ denotes the unique geodesic path between $x,y\in\mathcal{T}_{g}$.
This subtree is naturally viewed as a rooted ordered geometric tree: its vertex set consists of the root $\rho$, the sampled points $p_{g}(t_{1}),\dots,p_{g}(t_{k})$, and all branch points along their geodesics; edges are the geodesic segments connecting consecutive vertices, endowed with their metric lengths; and the left–right order is inherited from $\preceq_{g}$.
Definition 2.4 (Uniform $k$-point subtree induced by $g$).
Let $k\geq 1$, and let $U_{1},\dots,U_{k}$ be i.i.d. $\mathrm{Uniform}[0,1]$, independent of $g$. The uniform $k$-point subtree of $\mathcal{T}_{g}$ is
$\mathcal{T}_{g,k}:=\mathcal{T}_{g}(U_{1},\ldots,U_{k}).$
Remark 2.5.
Suppose $g,g^{\prime}\in C_{\rm bridge}$ differ only by a strictly increasing time change; that is, there exists a strictly increasing homeomorphism $\phi:[0,1]\to[0,1]$ such that
$g(t)=g^{\prime}\big(\phi_{t}\big),\qquad t\in[0,1].$
Then the map
$\Psi:\mathcal{T}_{g}\longrightarrow\mathcal{T}_{g^{\prime}},\qquad\Psi\big(p_{g}(t)\big):=p_{g^{\prime}}\big(\phi_{t}\big),$
defines an isomorphism between the rooted ordered real trees $\mathcal{T}_{g}$ and $\mathcal{T}_{g^{\prime}}$. In particular, we have for any $t_{1},\ldots,t_{k}\in[0,1]$, the subtrees
$\mathcal{T}_{g}(t_{1},\ldots,t_{k})\quad\text{and}\quad\mathcal{T}_{g^{\prime}}(\phi_{t_{1}},\ldots,\phi_{t_{k}})$
are isomorphic.
2.2.3 Convergence of contour functions
In this part, we present a fundamental criterion, originally due to Aldous, which characterizes the weak convergence of contour functions in terms of tightness and finite-dimensional convergence of the tree sequence.
[…]
where $b^{i}_{(0)}=b^{i}_{(k)}$ is the root of $T^{i}$, and for $1\leq j\leq k-1$, $b^{i}_{(j)}$ denotes the branch point of $V^{i}_{(j)}$ and $V^{i}_{(j+1)}$ in $T^{i}$.
Theorem 2.6.
Let $(\mathcal{T},d)=(\mathcal{T}_{g},d_{g})$ be a random rooted ordered real tree encoded by a random $g\in C_{\rm bridge}$. Consider a sequence $(\mathscr{T}^{n}:n\geq 1)$ of random ordered discrete trees with $a_{n}:=|V(\mathscr{T}^{n})|\to\infty$, and suppose that the maximal edge length of $\mathscr{T}^{n}$ converges to 0 in distribution. Then the following are equivalent:
  1. (i)
    The contour functions
    $C_{\lfloor 2(a_{n}-1)\,\cdot\,\rfloor}(\mathscr{T}^{n})\ \overset{d}{\longrightarrow}\ g\quad\text{in $D[0,1]$.}$
  2. (ii)
    For each $1\leq k\leq a_{n}$, let $(V_{n,1},\dots,V_{n,a_{n}})$ be a uniform random ordering of the vertices of $\mathscr{T}^{n}$, and let $\mathscr{T}^{n}_{k}$ be the rooted ordered subtree spanned by the root $\rho$ and $\{V_{n,1},\dots,V_{n,k}\}$. Then it holds that:
    • Tightness: define
      $\Delta(n,k):=\max_{v\in V(\mathscr{T}^{n})}\min_{u\in V(\mathscr{T}^{n}_{k})}d(v,u).$ (5)
      Then, for any $\delta>0$,
      $\lim_{k\to\infty}\limsup_{n\to\infty}\mathbb{P}(\Delta(n,k)>\delta)=0.$
    • Finite-dimensional convergence: for every $k\in\mathbb{N}_{+}$,
      $\big(\mathscr{T}^{n}_{k},\{V_{n,1},\dots,V_{n,k}\}\big)\overset{d}{\longrightarrow}\big(\mathcal{T}_{g,k},\{p_{g}(U_{1}),\dots,p_{g}(U_{k})\}\big)$ (6)
      under the distance (4), where $\{V_{n,1},\dots,V_{n,k}\}$ and $\{p_{g}(U_{1}),\dots,p_{g}(U_{k})\}$ are considered as multisets.
Remark 2.7.
Aldous originally proved this equivalence in [1993Aldous, Theorem 20] for the case where the limiting tree is a compact binary real tree encoded by an excursion. In that setting, his notion of finite-dimensional distributions differs from (6). Nevertheless, the first step of his proof shows that his definition is equivalent to requiring, for every $k\geq 1$,
$\begin{aligned} &\big(C_{U_{n,(1)}}(\mathscr{T}^{n}),\inf_{U_{(1)}<t<U_{(2)}}C_{t}(\mathscr{T}^{n}),C_{U_{n,(2)}}(\mathscr{T}^{n}),\inf_{U_{(2)}<t<U_{(3)}}C_{t}(\mathscr{T}^{n}),\ldots,C_{U_{n,(k)}}(\mathscr{T}^{n})\big)\\ &\overset{d}{\longrightarrow}\big(g(U_{(1)}),\inf_{U_{(1)}<t<U_{(2)}}g(t),g(U_{(2)}),\inf_{U_{(2)}<t<U_{(3)}}g(t),\ldots,g(U_{(k)})\big),\end{aligned}$ (7)
where $(U_{n,1},\ldots,U_{n,k})$ is a random vector sampled uniformly without replacement from $\{0,\ldots,2(a_{n}-1)\}$, with $U_{n,(i)}$ denoting its order statistics, and $U_{(1)},\ldots,U_{(k)}$ are the order statistics of $k$ i.i.d. uniform $(0,1)$ random variables. By [1993Aldous, Lemma 12], (7) is readily seen to be equivalent to (6). This is the only step in his argument that relies on the assumption of a binary limiting tree; the remainder of the proof proceeds without invoking it. Thus, Theorem 2.6 is already implicit in Aldous’ proof.
The formulation in (6) is more suitable for our setting, where the root of the limiting tree is not necessarily binary.
2.3 Conditioned Brownian motion and its subtree distribution
Recall the notion of a continuous profile introduced in Section 1. In this section, we first construct a canonical version of $W^{\ell}$ based on Aldous’ construction [1998BROWNIAN]. We then study the laws of the uniform $k$-point subtrees induced by $W^{\ell}$ and by the time-changed process $W^{\ell^{\sigma}}_{\alpha^{-1}(t)}$. These results will be used in the proof of finite-dimensional convergence in Section 5.1.
2.3.1 Canonical version of $W^{\ell}$
Recall that $L_{t}(x)$ denotes the local time of $W$ at level $x$ up to time $t$, and that $\tau_{u}$ represents the inverse local time at the origin. For any non-negative measurable function $f$ on $\mathbb{R}^{+}$, we write
$I(f):=\int_{0}^{\infty}f(x)\,\mathrm{d}x.$
In [1998BROWNIAN, Section 4], Aldous derived the law of the reflected Brownian bridge on time interval $[0,1]$ conditioned on having local time profile $\ell$, where $\ell$ may lie in a broad class of functions. This class includes, in particular, all continuous profiles. In our notation, this is exactly the law of $W^{\ell}$ for continuous profiles $\ell$ satisfying $h(\ell)<\infty$ and $I(\ell)=1$.
We now explain how this conditional law is described. Let $\mathcal{T}^{\ell}$ denote the rooted ordered real tree encoded by $W^{\ell}$, and let $\mathcal{T}^{\ell}_{k}$ denote its uniform $k$-point subtree. Aldous [1998BROWNIAN] and, subsequently, Warren [warren1999result] determined the law of $\mathcal{T}^{\ell}_{k}$, which will be described in detail later in Section 2.3.2. Since the family $\{\mathcal{T}^{\ell}_{k}\}_{k\geq 1}$ is consistent in $k$, the correspondence theorem of Aldous [1993Aldous, Theorem 15] [^6] A direct extension of the correspondence theorem from excursion functions to bridge functions applies here. yields a version of $W^{\ell}$. Moreover, this version of the conditional law is continuous in $\ell$ in a suitable sense.
Next, we extend Aldous’ version of $W^{\ell}$ to all continuous profiles $\ell$ with $h(\ell)<\infty$ (i.e., without the restriction $I(\ell)=1$). This follows from the scaling invariance of reflected Brownian motion and its local times:
Lemma 2.8.
For any $c,u>0$,
$\big(W_{t},\ L_{\tau_{u}}(x):0\leq t\leq\tau_{u},\ x\geq 0\big)\ \overset{d}{=}\ \big(c\,W_{t/c^{2}},\ c\,L_{\tau_{u/c}}(x/c):0\leq t\leq c^{2}\tau_{u/c},\ x\geq 0\big).$
Definition 2.9 (Canonical version of $W^{\ell}$).
Let $\ell$ be a continuous profile with $h(\ell)<\infty$. Set
$\tilde{\ell}(s):=I(\ell)^{-1/2}\,\ell\big(I(\ell)s\big),\qquad s\geq 0,$
so that $I(\tilde{\ell})=1$. Let $W^{\tilde{\ell}}$ denote the version of the conditioned reflected Brownian motion associated with the normalized profile $\tilde{\ell}$ constructed above. We then define
$W^{\ell}_{t}:=I(\ell)^{1/2}\,W^{\tilde{\ell}}_{\,t/I(\ell)},\qquad 0\leq t\leq I(\ell).$
By Lemma 2.8, this defines a version of conditional law, which will serve as our canonical version of $W^{\ell}$ throughout the paper.
Many properties for the case $I(\ell)=1$ (cf. [1998BROWNIAN, Construction 1]) extend to this general setting, including continuity in $\ell$ and the following:
Proposition 2.10.
  1. (i)
    For any continuous profile $\ell$ with $h(\ell)<\infty$, the local time profile of $W^{\ell}$ is $\ell$; that is, for every non-negative measurable $f$ on $\mathbb{R}_{+}$,
    $\int_{0}^{\infty}\ell(s)\,f(s)\,\mathrm{d}s\ =\ \int_{0}^{I(\ell)}f\big(W^{\ell}_{t}\big)\,\mathrm{d}t\quad\text{a.s.}.$ (8)
    In particular, the total time duration of $W^{\ell}$ is $I(\ell)$.
  2. (ii)
    The process $W^{\ell}$ is a bridge from 0 to 0 on the time interval $0,I(\ell)]$.
We now extend the associated tree notation to this setting.
Definition 2.11 (Real tree encoded by $W^{\ell}$).
Let $\ell$ be a continuous profile with $h(\ell)<\infty$. Since $W^{\ell}_{I(\ell)t}\in C_{\mathrm{bridge}}$ by Proposition 2.10 (ii), we define $\mathcal{T}^{\ell}$ to be the rooted ordered real tree encoded by $W^{\ell}_{I(\ell)t}$. We denote by $\mathcal{T}^{\ell}_{k}$ its uniform $k$-point subtree.
Remark 2.12 (Warren–Yor’s construction).
In [warren98], Warren and Yor provided an explicit construction of $W^{\ell}$ for continuous profiles $\ell$ satisfying $\ell(0)>0$ and $h(\ell)<\infty$. Roughly speaking, they introduced a process called the Brownian burglar, which is obtained through a suitable time-space transformation of a Brownian path. A version of $W^{\ell}$ can then be constructed from the Brownian burglar by applying another time-space transformation.
In particular, this version of $W^{\ell}$ is continuous in $\ell$ in the following sense (cf. [li2024inverting, Proposition 5.5]): if $(\ell^{n})$ and $\ell$ are continuous profiles with $h(\ell_{n})\rightarrow h(\ell)$ and $\ell_{n}\rightarrow\ell$ uniformly on compact subsets of $(0,h(\ell))$, then
$W^{\ell^{n}}\xrightarrow{d}W^{\ell}\quad\text{in }C[0,\infty),$
where each process is extended by 0 outside its original domain. In particular, this continuity allows the definition of $W^{\ell}$ to be extended to continuous profiles with $\ell(0)=0$ by approximation.
By comparing the continuity of the canonical version $W^{\ell}$ with respect to $\ell$, one can readily deduce that the two versions coincide for all continuous profiles $\ell$ with $h(\ell)<\infty$. Hence, the Warren–Yor’s approach gives an explicit realization of $W^{\ell}$.
2.3.2 Distribution of $\mathcal{T}^{\ell}_{k}$
In this part, we describe the distribution of the uniform $k$-point subtree $\mathcal{T}^{\ell}_{k}$. We recall Aldous’ coalescent construction and then reformulate it in terms of a piecewise Kingman representation. The latter formulation will play a key role in Section 5.1. In particular, it makes explicit that the finite-dimensional convergence of the tree sequence is equivalent to the convergence of the associated coalescent dynamics to Kingman’s coalescent.
Aldous’ coalescent construction
In [1998BROWNIAN], the uniform $k$-point subtree $\mathcal{T}^{\ell}_{k}$ is described via a coalescent process of $k$ particles. Fix a continuous profile $\ell$ with $h(\ell)<\infty$ and $I(\ell)=1$. Think of “time” $t$ running downward from $\infty$ to 0. Each of $k$ particles is born at independent random times with probability density $\ell(x)\,\mathrm{d}x$. As time decreases, clusters of particles merge according to the following rule: in the interval $t,t-\mathrm{d}t]$, each pair of existing clusters coalesces into one cluster with probability $\frac{4}{\ell(t)}\,\mathrm{d}t$. At time $t=0$, all remaining clusters merge into a single cluster.
The coalescent dynamics described above determine a rooted unordered real tree with $k$ leaves. The $k$ birth times specify the heights of the leaves, and each merge event at height $t>0$ creates a branching point at height $t$. The root is located at height 0.
To obtain an ordered tree, we assign a left-right ordering at each branching point. Whenever two clusters merge at height $t>0$, one is designated as the left subtree and the other as the right subtree, with the choice made uniformly at random. At time $t=0$, when the final merge occurs, a uniformly random ordering is again assigned. This procedure yields a rooted ordered real tree. Aldous [1998BROWNIAN] showed that the law of this tree coincides with the law of the uniform $k$-point subtree $\mathcal{T}^{\ell}_{k}$.
For a general continuous profile $\ell$ with $h(\ell)<\infty$, one easily checks that the law of $\mathcal{T}^{\ell}_{k}$ is given by the same coalescent construction, except that the birth density $\ell(x)\,\mathrm{d}x$ is replaced by its normalized form $\frac{\ell(x)}{I(\ell)}\,\mathrm{d}x$.
[…]
Figure 1: An illustration of the coalescent construction of the uniform $6$-point subtree. The vertical axis represents time decreasing from top to bottom. Each leaf $i=1,\ldots,6$ is placed at height $x_{i}$ equal to its birth time, and each $b_{i,j}$ denotes the branching point at which the ancestral clusters of $i$ and $j$ merge. The final merge occurs at height 0, which corresponds to the root.
Piecewise Kingman representation
The Aldous’ coalescent construction suggests a useful way to understand the tree $\mathcal{T}^{\ell}_{k}$ through the evolution of the number of ancestral lineages across heights. For $s\geq 0$, let $Y^{(k)}_{s}$ denote the number of points in $\mathcal{T}^{\ell}_{k}$ at height $s$. Then $Y^{(k)}_{s}$ can be naturally linked to a piecewise Kingman’s coalescent (here and below, “Kingman’s coalescent” always refers to its block-counting process), defined as follows.
Definition 2.13 (Piecewise Kingman’s coalescent).
For real numbers $x_{1}>x_{2}>\cdots>x_{k}>0$, the piecewise Kingman’s coalescent $\big(Z^{(k)}_{s}=Z^{(k)}_{s}(x_{1},\ldots,x_{k}):s\geq 0\big)$ with rate $\frac{4}{\ell(s)}$ is defined as:
  • For $s>x_{1}$, we set $Z^{(k)}_{s}=0$.
  • Recursively, for each $i=1,\ldots,k-1$, conditionally on $\big(Z^{(k)}_{s}:s>x_{i}\big)$, the process $Z^{(k)}_{s}$ on the interval $s\in(x_{i+1},x_{i}]$ evolves as a Kingman’s coalescent started with $\lim_{s\downarrow x_{i}}Z^{(k)}_{s}+1$ lineages, where each pair of lineages coalesces at rate $\frac{4}{\ell(s)}$. Here the height parameter $s$ is interpreted as backward time, so the coalescent runs as $s$ decreases.
Proposition 2.14.
For any $k\geq 1$ and real numbers $x_{1}>x_{2}>\cdots>x_{k}>0$, conditionally on the leaf heights being $x_{1},\ldots,x_{k}$, the process $Y^{(k)}_{s}$ evolves as the piecewise Kingman’s coalescent $\big(Z^{(k)}_{s}(x_{1},\ldots,x_{k}):s\geq 0\big)$.
We mention that the additional term +1+1 at height $x_{i+1}$ in the definition of $Z^{(k)}$ corresponds, in the process $Y^{(k)}$, to the introduction of a new leaf at its birth height $x_{i+1}$.
This proposition gives rise to the following equivalent construction of $\mathcal{T}^{\ell}_{k}$ for continuous profile $\ell$ with $h(\ell)<\infty$:
  1. 1.
    Sample $k$ heights independently with density $\frac{\ell(x)}{I(\ell)}\,\mathrm{d}x$, and reorder them so that $x_{1}>\cdots>x_{k}$.
  2. 2.
    Sample a piecewise Kingman’s coalescent process $\big(Z^{(k)}_{s}(x_{1},\ldots,x_{k}):s\geq 0\big)$.
  3. 3.
    Conditional on $\{Z^{(k)}_{s}\}_{s\geq 0}$, sample uniformly from the set of rooted ordered real trees whose lineage counts at each height agree with $Z^{(k)}$.
Proposition 2.15.
The tree obtained from this construction has the same distribution as $\mathcal{T}^{\ell}_{k}$.
2.4 Distribution of uniform $k$-point subtrees induced by $W^{\ell^{\sigma}}_{\alpha^{-1}(t)}$
We now turn to the time-changed process $W^{\ell^{\sigma}}_{\alpha^{-1}(t)}$. As a first step, we verify that $\alpha$ is strictly increasing, so that its inverse $\alpha^{-1}$ is well-defined. We then show that the total time duration of $W^{\ell^{\sigma}}_{\alpha^{-1}(t)}$ is equal to $I(\ell)$, which implies that the rescaled process $W^{\ell^{\sigma}}_{\alpha^{-1}(I(\ell)t)}\in C_{\mathrm{bridge}}$. After that, we describe the law of the uniform $k$-point subtrees induced by this process.
Lemma 2.16.
For any continuous profiles $\ell$ and $\sigma$ satisfying (H1), the function
$\alpha(t)=\frac{1}{4}\int_{0}^{t}\sigma\big(W^{\ell^{\sigma}}_{s}\big)^{2}\,\mathrm{d}s$
is a.s. strictly increasing.
Proof 2.17.
By Proposition 2.10, the total time duration of $W^{\ell^{\sigma}}$ equals $I(\ell^{\sigma})$. Using (8), we have a.s.
$\mathrm{Leb}\{s\in[0,I(\ell^{\sigma})]:W^{\ell^{\sigma}}_{s}\notin(0,h(\ell))\}=0.$
Since $\sigma>0$ on $(0,h(\ell))$, it follows that $\sigma(W^{\ell^{\sigma}}_{s})^{2}>0$ for a.e. $s$, hence $\alpha$ is a.s. strictly increasing.
Proposition 2.18.
For any continuous profile $\ell$ with $h(\ell)<\infty$, The total time duration of $W^{\ell^{\sigma}}_{\alpha^{-1}(t)}$ is $I(\ell)$. Consequently, $W^{\ell^{\sigma}}_{\alpha^{-1}(I(\ell)t)}\in C_{\mathrm{bridge}}$.
Proof 2.19.
The total time duration of the time-changed process $W^{\ell^{\sigma}}_{\alpha^{-1}(t)}$ is
[…]
The fact that $W^{\ell^{\sigma}}_{\alpha^{-1}(I(\ell)t)}\in C_{\mathrm{bridge}}$ then follows from Proposition 2.10 (ii).
Let $\mathcal{T}^{\ell,\sigma}$ denote the rooted ordered real tree encoded by $W^{\ell^{\sigma}}_{\alpha^{-1}(I(\ell)t)}$, and let $\mathcal{T}^{\ell,\sigma}_{k}$ be its uniform $k$-point subtree. We now describe the distribution of $\mathcal{T}^{\ell,\sigma}_{k}$.
Theorem 2.20.
For any $k\in\mathbb{N}_{+}$, the tree $\mathcal{T}^{\ell,\sigma}_{k}$ admits the following construction:
  1. 1.
    Sample $k$ heights independently with density $\frac{\ell(x)}{I(\ell)}\,\mathrm{d}x$, and reorder them so that $x_{1}>\cdots>x_{k}$.
  2. 2.
    Sample a piecewise Kingman coalescent $\big(Z^{(k)}_{s}(x_{1},\ldots,x_{k}):s\geq 0\big)$ in which each pair of lineages coalesces at rate $\frac{4}{\ell^{\sigma}(s)}$.
  3. 3.
    Conditional on $\{Z^{(k)}_{s}\}_{s\geq 0}$, sample uniformly from the set of rooted ordered real trees whose lineage counts at each height agree with $Z^{(k)}$.
The proof proceeds by relating $\mathcal{T}^{\ell,\sigma}_{k}$ to $\mathcal{T}^{\ell^{\sigma}}_{k}$, the uniform $k$-point subtree induced by $W^{\ell^{\sigma}}_{I(\ell^{\sigma})t}$. Note that $W^{\ell^{\sigma}}_{\alpha^{-1}(I(\ell)t)}$ differs from $W^{\ell^{\sigma}}_{I(\ell^{\sigma})t}$ only by a strictly increasing time change:
$W^{\ell^{\sigma}}_{\alpha^{-1}(I(\ell)t)}=W^{\ell^{\sigma}}_{I(\ell^{\sigma})\,\phi_{t}},\qquad\phi_{t}:=\frac{\alpha^{-1}(I(\ell)t)}{I(\ell^{\sigma})}.$
Thus by Remark 2.5, for any $t_{1},\ldots,t_{k}\in(0,1)$,
$\mathcal{T}^{\ell,\sigma}(t_{1},\ldots,t_{k})\;\cong\;\mathcal{T}^{\ell^{\sigma}}(\phi_{t_{1}},\ldots,\phi_{t_{k}}).$ (9)
Proposition 2.21.
For any $k\in\mathbb{N}_{+}$, the laws of $\mathcal{T}^{\ell,\sigma}_{k}$ and $\mathcal{T}^{\ell^{\sigma}}_{k}$ are mutually absolutely continuous. More precisely, for any test function $f$,
$\mathbb{E}\big[f(\mathcal{T}^{\ell,\sigma}_{k})\big]=\left(\frac{I(\ell^{\sigma})}{I(\ell)}\right)^{k}\mathbb{E}\!\big[f(\mathcal{T}^{\ell^{\sigma}}_{k})\,\sigma(\mathcal{T}^{\ell^{\sigma}}_{k})^{2}\big],$ (10)
where
$\sigma(\mathcal{T}^{\ell^{\sigma}}_{k})=\prod_{j=1}^{k}\sigma(h_{j}),\qquad\text{$h_{1},\ldots,h_{k}$ denote the leaf heights of $\mathcal{T}^{\ell^{\sigma}}_{k}$.}$
Proof 2.22.
For any test function $f$,
[…]
Since
$\phi^{-1}(s)=\frac{\alpha(I(\ell^{\sigma})s)}{I(\ell)},\qquad\mathrm{d}\phi^{-1}(s)=\frac{I(\ell^{\sigma})}{I(\ell)}\,\sigma(W^{\ell^{\sigma}}_{I(\ell^{\sigma})s})^{2}\,\mathrm{d}s,$
substituting into (11) yields
$\mathbb{E}\big[f(\mathcal{T}^{\ell,\sigma}_{k})\big]$ $=\left(\frac{I(\ell^{\sigma})}{I(\ell)}\right)^{k}\int_{0}^{1}\cdots\int_{0}^{1}\mathbb{E}\big[f(\mathcal{T}^{\ell^{\sigma}}(s_{1},\ldots,s_{k}))\big]\prod_{j=1}^{k}\sigma(W^{\ell^{\sigma}}_{I(\ell^{\sigma})s_{j}})^{2}\,\mathrm{d}s_{1}\cdots\,\mathrm{d}s_{k}$
$=\left(\frac{I(\ell^{\sigma})}{I(\ell)}\right)^{k}\mathbb{E}\big[f(\mathcal{T}^{\ell^{\sigma}}_{k})\sigma(\mathcal{T}^{\ell^{\sigma}}_{k})^{2}\big].$
This completes the proof.
Proof 2.23 (Proof of Theorem 2.20).
Let $\mathring{\mathcal{T}}^{\ell,\sigma}_{k}$ denote the random rooted ordered tree defined by the three-step procedure in the statement of the theorem. We assume that $\mathcal{T}^{\ell^{\sigma}}$ is realized via the piecewise Kingman representation in Proposition 2.14.
For any $x_{1}>\cdots>x_{k}>0$, the conditional laws of $\mathring{\mathcal{T}}^{\ell,\sigma}_{k}$ and $\mathcal{T}^{\ell^{\sigma}}_{k}$ given that the leaf heights equal $(x_{1},\ldots,x_{k})$ coincide. Thus, the only difference between their unconditional distributions comes from the respective sampling densities of the leaf heights. Comparing these densities yields, for any test function $f$,
$\mathbb{E}\big[f(\mathring{\mathcal{T}}^{\ell,\sigma}_{k})\big]=\left(\frac{I(\ell^{\sigma})}{I(\ell)}\right)^{k}\,\mathbb{E}\!\big[f(\mathcal{T}^{\ell^{\sigma}}_{k})\,\sigma(\mathcal{T}^{\ell^{\sigma}}_{k})^{2}\big].$
Together with (10), this implies $\mathring{\mathcal{T}}^{\ell,\sigma}_{k}\overset{d}{=}\mathcal{T}^{\ell,\sigma}_{k}$, completing the proof.
3 Coalescent process
In this section, we first derive preparatory moment and probability estimates for elementary coalescence events (Section 3.1); these bounds will be used repeatedly in Sections 35. We then introduce the coalescent process arising in Cannings trees, show that it forms a Markov chain, and prove a coming-down-from-infinity property that serves as a central ingredient for the tightness of $\widetilde{\mathscr{T}}^{n}$.
We begin with some notation. Denote by $x_{1}^{n,s},\dots,x_{q_{n}(s)}^{n,s}$ the vertices at generation $s$ of $\mathscr{T}^{n}$, listed according to the lexicographic order. Let $(u_{1}^{n,s},\dots,u_{q_{n}(s)}^{n,s})$ be a uniform random permutation of $(x_{1}^{n,s},\dots,x_{q_{n}(s)}^{n,s})$, independent across different generations and independent of $\mathscr{T}^{n}$.
For any vertex $x$, we write $\xi(x)$ for the number of children of $x$ and $\mathfrak{p}(x)$ for the parent of $x$. In particular, we have $\nu_{i}^{n,s}=\xi(x_{i}^{n,s})$.
For notational convenience, from this point onward we write $\mathbb{P}_{n}$ and $\mathbb{E}_{n}$ for the probability and expectation under the model indexed by $n$, respectively, and we omit the superscript $n$ in $\nu^{n,s}$, $x_{i}^{n,s}$, and $u_{i}^{n,s}$ whenever it is clear from context.
3.1 Coalescent probability
In this subsection, we introduce coalescent events in Cannings trees and express their probabilities in terms of the moments of the offspring counts $\nu$. We then estimate these moments, which in turn yields quantitative estimates for the coalescent probabilities.
For $n\geq 1$ and $1\leq s<h(q_{n})$, define
$M^{s}=M^{n,s}:=\Big\{f:A\to\{1,\ldots,q_{n}(s)\}\,:\,A\subset\{1,\ldots,q_{n}(s+1)\}\Big\},$ (12)
the collection of all mappings from a subset of $\{1,\ldots,q_{n}(s+1)\}$ to $\{1,\ldots,q_{n}(s)\}$.
Each $f\in M^{s}$ naturally encodes a coalescent event:
$\text{Coal}^{s}(f)=\text{Coal}^{n,s}(f):=\Big\{\forall\,j\in\mathrm{Dom}(f),\ \mathfrak{p}(u^{n,s+1}_{j})=x^{n,s}_{f(j)}\Big\}.$ (13)
Suppose the range of $f$ is $\{j_{1}<j_{2}<\cdots<j_{m}\}\subset\{1,\ldots,q_{n}(s+1)\}$. We then define
$N_{f}:=\big(\#f^{-1}(j_{1}),\ldots,\#f^{-1}(j_{m})\big),$
which records the multiplicities of coalescence in the event $\text{Coal}^{s}(f)$.
3.1.1 Linking coalescent probability to offspring moments
Proposition 3.1.
For $n\geq 1$, $1\leq s<h(q_{n})$, and $f\in M^{s}$ with $N_{f}=(N_{1},\ldots,N_{m})$ and $\mathrm{Dom}(f):=\sum_{i=1}^{m}N_{i}$, we have
${\mathbb{P}}_{n}\big(\text{Coal}^{s}(f)\big)=\frac{1}{\big(q_{n}(s+1)\big)_{\#\mathrm{Dom}(f)}}\,\mathbb{E}_{n}\!\Bigg[\prod_{i=1}^{m}\big(\nu^{s}_{j_{i}}\big)_{N_{i}}\Bigg],$ (14)
where $(l)_{k}:=l(l-1)\cdots(l-k+1)$ denotes the falling factorial.
Proof 3.2.
[…]
$\frac{\prod_{i=1}^{m}(v^{s}_{j_{i}})_{N_{i}}}{(q_{n}(s+1))_{\#\mathrm{Dom}(f)}}.$
Taking expectation over $(\nu^{s}_{j_{i}})$ yields
${\mathbb{P}}_{n}\big(\text{Coal}^{s}(f)\big)=\frac{1}{(q_{n}(s+1))_{\#\mathrm{Dom}(f)}}\,{\mathbb{E}}_{n}\Bigg[\prod_{i=1}^{m}\big(\nu^{s}_{j_{i}}\big)_{N_{i}}\Bigg],$
as claimed.
We now record several direct consequences of Proposition 3.1, which will be used in the sequel.
Corollary 3.3.
For $n\geq 1$, $1\leq s<h(q_{n})$, $1\leq l\leq m\leq q_{n}(s+1)$, it holds:
  1. (1)
    ${\mathbb{P}}_{n}\!\Big(\{\mathfrak{p}(u_{1}^{s+1})=\cdots=\mathfrak{p}(u_{l}^{s+1})\}\cap\{\mathfrak{p}(u_{l}^{s+1}),\dots,\mathfrak{p}(u_{m}^{s+1})~\text{are distinct}\}\Big)$
    $=$ $\frac{(q_{n}(s))_{m-l+1}}{(q_{n}(s+1))_{m}}\,{\mathbb{E}}_{n}\!\Big[(\nu_{1}^{s})_{l}\prod_{i=2}^{m-l+1}\nu_{i}^{s}\Big],$
  2. (2)
    ${\mathbb{P}}_{n}\!\Big(\mathfrak{p}(u_{1}^{s+1})=\mathfrak{p}(u_{2}^{s+1})\neq\mathfrak{p}(u_{3}^{s+1})=\mathfrak{p}(u_{4}^{s+1})\Big)=\frac{(q_{n}(s))_{2}}{(q_{n}(s+1))_{4}}\,{\mathbb{E}}_{n}\!\Big[(\nu_{1}^{s})_{2}(\nu_{2}^{s})_{2}\Big].$
For later use in the proof of the main theorem, when applying part (1) of Corollary 3.3, we will mainly need the cases $l=1$ or $2$. Following the same argument as in the proof of Proposition 3.1, Corollary 3.3 admits the following refinement.
Corollary 3.4.
For $n,K\geq 1$, $1\leq s<h(q_{n})$ and $1\leq l\leq q_{n}(s+1)$, it holds:
  • (1)
    ${\mathbb{P}}_{n}\!\Big(\mathfrak{p}(u_{1}^{s+1})=\cdots=\mathfrak{p}(u_{l}^{s+1}),\;\xi(\mathfrak{p}(u_{1}^{s+1}))\leq K\Big)=\frac{q_{n}(s)}{(q_{n}(s+1))_{l}}\,{\mathbb{E}}_{n}\!\Big[(\nu_{1}^{s})_{l}\mathbf{1}\{\nu_{1}^{s}\leq K\}\Big],$
  • (2)
    ${\mathbb{P}}_{n}\!\Big(\mathfrak{p}(u_{1}^{s+1})=\mathfrak{p}(u_{2}^{s+1})\neq\mathfrak{p}(u_{3}^{s+1})=\mathfrak{p}(u_{4}^{s+1}),\,\xi(\mathfrak{p}(u_{1}^{s+1}))\leq K,\;\xi(\mathfrak{p}(u_{3}^{s+1}))\leq K\Big)$
    $=$ $\frac{(q_{n}(s))_{2}}{(q_{n}(s+1))_{4}}\,{\mathbb{E}}_{n}\!\Big[(\nu_{1}^{s})_{2}(\nu_{2}^{s})_{2}\mathbf{1}\{\nu_{1}^{s},\,\nu_{2}^{s}\leq K\}\Big].$
3.1.2 Estimates for offspring moments
In this part, we provide estimates for the offspring moments appearing in Corollary 3.3. We start with the $(2,2)$-type moment appearing in Corollary 3.3 (2). Recall the conditions (H1) and (H2”) from Section 1.1.
Lemma 3.5.
For any $n$ sufficiently large, $\delta>0$ small and $\delta n\leq s\leq h(q_{n})-\delta n$,
${\mathbb{E}}_{n}[(\nu^{s}_{1})^{2}(\nu^{s}_{2})^{2}]\leq\frac{q_{n}(s+1)}{q_{n}(s)-1}{\mathbb{E}}_{n}[(\nu^{s}_{1})^{3}]$ (15)
In particular, under conditions (H1) and (H2”), for any $\delta>0$,
$\lim_{n\to\infty}\sup_{\delta n\leq s\leq h(q_{n})-\delta n}\frac{\mathbb{E}_{n}[(\nu_{1}^{s})^{2}(\nu_{2}^{s})^{2}]}{n}=0.$ (16)
Proof 3.6.
We observe that $(\nu^{s}_{1})^{2}(\nu^{s}_{2})^{2}\leq\nu^{s}_{1}(\nu^{s}_{2})^{3}+(\nu^{s}_{1})^{3}\nu^{s}_{2}$ and hence ${\mathbb{E}}_{n}[(\nu^{s}_{1})^{2}(\nu^{s}_{2})^{2}]\leq{\mathbb{E}}_{n}[(\nu^{s}_{1})^{3}\nu^{s}_{2}]$ by the exchangeability of $\nu$. In the same reason, we derive that
${\mathbb{E}}_{n}[(\nu^{s}_{1})^{3}\nu^{s}_{2}]=\frac{1}{q_{n}(s)-1}{\mathbb{E}}_{n}\Big[(\nu^{s}_{1})^{3}\sum_{j=2}^{q_{n}(s)}\nu^{s}_{j}\Big]$
$=\,$ $\frac{q_{n}(s+1)}{q_{n}(s)-1}{\mathbb{E}}_{n}[(\nu^{s}_{1})^{3}]-\frac{1}{q_{n}(s)-1}{\mathbb{E}}_{n}[(\nu^{s}_{1})^{4}]\leq\frac{q_{n}(s+1)}{q_{n}(s)-1}{\mathbb{E}}_{n}[(\nu^{s}_{1})^{3}].$
Therefore (15) follows. Under assumption (H1), since $\ell$ is continuous, it is easy to obtain that for all $\delta>0$, $\sup_{\delta n\leq s\leq h(q_{n})-\delta n}\frac{q_{n}(s+1)}{q_{n}(s)-1}\to 1$ as $n\to\infty$. (16) is then a consequence of (15) and (H2”).
Next, we consider the moments in Corollary 3.3 (1) with $l=1,2$.
Lemma 3.7.
Assume (H1) and (H2”). For any $\delta>0$ and $m\geq 1$, uniformly for $\delta n\leq s\leq h(q_{n})-\delta n$, we have:
  • (1)
    $\mathbb{E}_{n}\Big[\prod_{i=1}^{m}\nu_{i}^{s}\Big]=1+o_{n}(1),$ (17)
    and more precisely,
    $\mathbb{E}_{n}\Big[\prod_{i=1}^{m}\nu_{i}^{s}\Big]=\frac{(q_{n}(s+1))^{m}}{(q_{n}(s))_{m}}-\dfrac{\binom{m}{2}\big(1+\sigma_{n}^{2}(s)\big)}{q_{n}(s)}\big(1+o_{n}(1)\big);$ (18)
  • (2)
    $\mathbb{E}_{n}\Big[(\nu_{1}^{s})^{2}\prod_{i=2}^{m}\nu_{i}^{s}\Big]=1+\sigma_{n}(s)^{2}+o_{n}(1).$ (19)
Proof 3.8.
Set
$a_{m}$ $:=\mathbb{E}_{n}\Big[\prod_{i=1}^{m}\nu_{i}^{s}\Big],\quad 1\leq m\leq q_{n}(s),$
$b_{1}$ $:=\mathbb{E}_{n}[(\nu_{1}^{s})^{2}],\quad b_{m}:=\mathbb{E}_{n}\Big[(\nu_{1}^{s})^{2}\prod_{i=2}^{m}\nu_{i}^{s}\Big],\quad 2\leq m\leq q_{n}(s),$
$c_{1}$ $:=\mathbb{E}_{n}[(\nu_{1}^{s})^{3}],\quad c_{m}:=\mathbb{E}_{n}\left[(\nu_{1}^{s})^{3}\prod_{i=2}^{m}\nu_{i}^{s}\right],\quad 2\leq m\leq q_{n}(s),$
$d_{2}$ $:=\mathbb{E}_{n}[(\nu_{1}^{s})^{2}(\nu_{2}^{s})^{2}],\quad d_{m}:=\mathbb{E}_{n}\left[(\nu_{1}^{s})^{2}(\nu_{2}^{s})^{2}\prod_{i=3}^{m}\nu_{i}^{s}\right],\quad 3\leq m\leq q_{n}(s).$
By conditioning on $\nu_{1}^{s},\dots,\nu_{m-1}^{s}$, we obtain by induction that
$a_{m}$ $=\mathbb{E}_{n}\Bigg[\frac{q_{n}(s+1)-\sum_{i=1}^{m-1}\nu_{i}^{s}}{q_{n}(s)-(m-1)}\prod_{i=1}^{m-1}\nu_{i}^{s}\Bigg]=\frac{q_{n}(s+1)}{q_{n}(s)-(m-1)}a_{m-1}-\frac{m-1}{q_{n}(s)-(m-1)}b_{m-1}$
$=\cdots=\frac{(q_{n}(s+1))^{m}}{(q_{n}(s))_{m}}-\sum_{i=1}^{m-1}\frac{(q_{n}(s+1))^{m-i-1}}{(q_{n}(s)-i)_{m-i}}\cdot ib_{i},\quad m\geq 1.$ (20)
where in the last step we use the fact ${\mathbb{E}}_{n}[\nu^{s}_{1}]=\frac{q_{n}(s+1)}{q_{n}(s)}$. Similarly, for $b_{m}$,
$b_{m}=\frac{(q_{n}(s+1))^{m-1}}{(q_{n}(s)-1)_{m-1}}b_{1}-\sum_{i=1}^{m-1}\frac{(q_{n}(s+1))^{m-i-1}}{(q_{n}(s)-i)_{m-i}}(c_{i}+(i-1)d_{i}),$ (21)
and in particular,
$b_{m}\leq\frac{(q_{n}(s+1))^{m-1}}{(q_{n}(s)-1)_{m-1}}b_{1}.$ (22)
Combining (3.8) and (22) gives that
$\left|a_{m}-\frac{(q_{n}(s+1))^{m}}{(q_{n}(s))_{m}}\right|$ $\leq\sum_{i=1}^{m-1}\frac{(q_{n}(s+1))^{m-i-1}}{(q_{n}(s)-i)_{m-i}}\cdot\frac{(q_{n}(s+1))^{i-1}}{(q_{n}(s)-1)_{i-1}}\cdot ib_{1}$
$=\frac{m(m-1)}{2}\frac{q_{n}(s+1)^{m-2}}{(q_{n}(s)-1)_{m-1}}b_{1}.$
By (H1), we have
$\frac{(q_{n}(s+1))^{m}}{(q_{n}(s))_{m}}=1+o_{n}(1),\qquad\frac{(q_{n}(s+1))^{m-2}}{(q_{n}(s)-1)_{m-1}}=O_{n}\!\left(\tfrac{1}{n}\right),$
where both $o_{n}(1)$ and $O_{n}(\tfrac{1}{n})$ are uniform for $\delta n\leq s\leq h(q_{n})-\delta n$. Hence, (17) follows.
The bound (19) follows similarly. Indeed, in the same way as (22),
$c_{m}\leq\frac{(q_{n}(s+1))^{m-1}}{(q_{n}(s)-1)_{m-1}}c_{1},\quad d_{m}\leq\frac{(q_{n}(s+1))^{m-2}}{(q_{n}(s)-1)_{m-2}}d_{2}.$
Approximating $b_{1}$ by (H1), $c_{1}$ by (H2”) and $d_{2}$ by Lemma 3.5, we deduce (19) from (21). Finally, by (H1), uniformly over $i=1,\cdots,m-1$,
$\frac{(q_{n}(s+1))^{m-i-1}}{(q_{n}(s)-i)_{m-i}}=\frac{1}{q_{n}(s)}\big(1+o_{n}(1)\big),$
and thus (18) follows directly from (19) together with (3.8).
Corollary 3.9.
Assume (H1) and (H2”). For any $\delta>0$ and $m\geq 1$, uniformly for $\delta n\leq s\leq h(q_{n})-\delta n$, we have:
  • (1)
    ${\mathbb{P}}_{n}\Big(\mathfrak{p}(u_{1}^{s+1}),\ldots,\mathfrak{p}(u_{m}^{s+1})~\text{are distinct}\Big)=1+o_{n}(1),$
    and more precisely,
    $\begin{aligned} {\mathbb{P}}_{n}\Big(\mathfrak{p}(u_{1}^{s+1}),\ldots,\mathfrak{p}(u_{m}^{s+1})~\text{are distinct}\Big)=1-\dfrac{\binom{m}{2}\sigma_{n}(s)^{2}}{q_{n}(s)}\big(1+o_{n}(1)\big);\end{aligned}$ (23)
  • (2)
    $\begin{aligned} &{\mathbb{P}}_{n}\Big(\{\mathfrak{p}(u_{1}^{s+1})=\mathfrak{p}(u_{2}^{s+1})\}\cap\{\mathfrak{p}(u_{2}^{s+1}),\ldots,\mathfrak{p}(u_{m}^{s+1})~\text{are distinct}\}\Big)\\ =\,&\frac{1}{q_{n}(s)}\big(\sigma_{n}(s)^{2}+o_{n}(1)\big).\end{aligned}$ (24)
Proof 3.10.
Recall that by (H1), $\frac{q_{n}(s+1)}{q_{n}(s)}=1+o(1)$ uniformly for $\delta n\leq s\leq h(q_{n})-\delta n$. Hence by Corollary 3.3 (with $l=1$) and (18), we deduce that
${\mathbb{P}}_{n}\!\Big(\mathfrak{p}(u_{1}^{s+1}),\dots,\mathfrak{p}(u_{m}^{s+1})~\text{are distinct}\Big)=\frac{(q_{n}(s))_{m}}{(q_{n}(s+1))_{m}}\,{\mathbb{E}}_{n}\!\Big[\prod_{i=1}^{m}\nu_{i}^{s}\Big]$
$=$ $\frac{(q_{n}(s+1))^{m}}{(q_{n}(s+1))_{m}}-\dfrac{\binom{m}{2}\big(1+\sigma_{n}^{2}(s)\big)}{q_{n}(s)}\big(1+o_{n}(1)\big).$
Expanding the first term on the last line gives
$\frac{(q_{n}(s+1))^{m}}{(q_{n}(s+1))_{m}}=\prod_{i=0}^{m-1}\left(1-\frac{i}{q_{n}(s+1)}\right)^{-1}=1+\binom{m}{2}\frac{1}{q_{n}(s+1)}+o_{n}\left(\frac{1}{q_{n}(s+1)}\right),$
which yields (23).
Moreover,
${\mathbb{E}}_{n}\!\Big[(\nu_{1}^{s})_{2}\prod_{i=2}^{m-1}\nu_{i}^{s}\Big]={\mathbb{E}}_{n}\!\Big[(\nu_{1}^{s})^{2}\prod_{i=2}^{m-1}\nu_{i}^{s}\Big]-{\mathbb{E}}_{n}\!\Big[\prod_{i=1}^{m-1}\nu_{i}^{s}\Big]=\sigma_{n}(s)^{2}+o_{n}(1).$
Therefore (24) follows from Corollary 3.3 (with $l=2$) and (19).
3.2 Coalescent process
In this subsection, we define the coalescent process and derive estimates for its transition probabilities.
We begin with the relevant notation. For any vertex $v\in\mathscr{T}^{n}$ and $0\leq i\leq|v|$, let $\mathrm{Anc}(v,i)$ denote the ancestor of $v$ at height $i$. Define the set of descendants of $v$ by
$\mathrm{Desc}(v)=\mathrm{Desc}(v,n):=\big\{\,u\in V(\mathscr{T}^{n}):|u|\geq|v|\text{ and }\mathrm{Anc}(u,|v|)=v\,\big\}.$
Definition 3.11 (Coalescent process).
Fix a height $0\leq h^{*}\leq h(q_{n})$ and $1\leq k\leq q_{n}(h^{*})$. For $0\leq j\leq h^{*}$, set
$\mathcal{X}^{(k)}_{j}=\mathcal{X}^{(k)}_{j}(n,h^{*})$ $=\big\{\text{Anc}(u^{h^{*}}_{i},j):1\leq i\leq k\big\}$ (25)
$=\big\{v\in V(\mathscr{T}^{n}):|v|=j,\ \operatorname{Desc}(v)\cap\{u^{h^{*}}_{i}:1\leq i\leq k\}\neq\emptyset\big\}.$
The coalescent process (based at height $h^{*}$) is then defined as
$X^{(k)}_{j}=X^{(k)}_{j}(n,h^{*}):=\#\mathcal{X}^{(k)}_{j},\qquad 0\leq j\leq h^{*}.$
In the special case $k=q_{n}(h^{*})$, we write $\mathcal{X}_{j}$ and $X_{j}$ in place of $\mathcal{X}^{(k)}_{j}$ and $X^{(k)}_{j}$, respectively.
In other words, $X^{(k)}_{j}$ represents the number of distinct ancestral lineages at height $j$ traced from the $k$ uniformly sampled vertices at height $h^{*}$. At $j=h^{*}$, we have $X^{(k)}_{0}=k$, since all sampled vertices are distinct. As $j$ decreases, the lineages are traced backward through the tree and may coalesce whenever two or more sampled vertices share a common ancestor. Thus, the process $\big(X^{(k)}_{j}:0\leq j\leq h^{*}\big)$ is non-increasing as $j$ decreases, and each downward jump corresponds to a coalescence event between ancestral lineages.
For each $v\in\mathscr{T}^{n}$ with $|v|\leq h^{*}$, we write
$\mathscr{T}^{n}(v):=\mathscr{T}^{n}(v,h^{*})$ (26)
for the subtree of $\mathscr{T}^{n}$ rooted at $v$, truncated at height $h^{*}$; that is, the subtree consisting of vertices
$\{u\in\operatorname{Desc}(v):|u|\leq h^{*}\}.$
We then define, for $0\leq j\leq h^{*}$ and $1\leq k\leq q_{n}(h^{*})$,
$\mathcal{G}^{(k)}_{j}=\mathcal{G}^{(k)}_{j}(h^{*}):=\sigma\!\big(\{\mathscr{T}^{n}(v):|v|=j\}\big)\vee\sigma(\{u^{h^{*}}_{i}:1\leq i\leq k\}).$ (27)
We emphasize that both collections $\{\mathscr{T}^{n}(v):|v|=j\}$ and $\{u^{h^{*}}_{i}:1\leq i\leq k\}$ are treated as unordered sets; in particular, $\mathcal{G}^{(k)}_{j}$ does not record any ordering among vertices at height $j$. When $k=q_{n}(h^{*})$, we write simply $\mathcal{G}_{j}=\mathcal{G}^{(k)}_{j}$, in which case
$\mathcal{G}_{j}=\sigma\!\big(\{\mathscr{T}^{n}(v):|v|=j\}\big).$
We now present some basic properties of the coalescent process $(X^{(k)}_{j})$.
Proposition 3.12.
For any $1\leq j\leq h^{*}$ and $1\leq k\leq q_{n}(h^{*})$, conditionally on $\mathcal{G}^{(k)}_{j}$, if $X^{(k)}_{j}=m$, then the set $\mathcal{X}_{j}$ is distributed as $\{u_{1}^{j},\dots,u_{m}^{j}\}$.
Proof 3.13.
By the exchangeability of the offspring vector $\nu^{j}=\big(\nu_{1}^{j},\ldots,\nu_{q_{n}(j)}^{j}\big)$, the vector of subtrees
$(\mathscr{T}^{n}(v):|v|=j)$
is also exchangeable, a fact that can be readily verified. The conclusion then follows immediately.
Theorem 3.14.
For any $1\leq k\leq q_{n}(h^{*})$, the process $(X^{(k)}_{h^{*}-j}:0\leq j\leq h^{*})$ is a Markov chain with respect to $(\mathcal{G}^{(k)}_{h^{*}-j}:0\leq j\leq h^{*})$. Moreover, for any $1\leq j\leq h^{*}$, its transition probabilities satisfy
${\mathbb{P}}_{n}\big(X^{(k)}_{j-1}=m_{2}\,\big|\,X^{(k)}_{j}=m_{1}\big)={\mathbb{P}}_{n}\!\left(\#\{\mathfrak{p}(u^{j}_{i}):1\leq i\leq m_{1}\}=m_{2}\right).$ (28)
We give the proof only for the special case $k=q_{n}(h^{*})$. The argument for general $k$ is entirely analogous and is therefore omitted. In the proof, we will use the following elementary property of conditional expectation.
Lemma 3.15.
Let $X,Y,Z$ be random elements on the same probability space, with $Y$ independent of $(X,Z)$. For any test function $f(x,y)$ on the value space of $(X,Y)$, define
$g(x):=\mathbb{E}[f(x,Y)].$
Then
$\mathbb{E}\big(f(X,Y)\mid Z\big)=\mathbb{E}\big(g(X)\mid Z\big).$ (29)
In particular, if there exists a random element $W$, independent of $Y$ and distributed as $(X\mid Z)$, then
$\mathbb{E}\big(f(X,Y)\mid Z\big)=\mathbb{E}\big(f(W,Y)\big).$ (30)
Proof 3.16.
By the standard property of conditional expectation for independent random elements (see, e.g., [durrett2019probability, Example 4.1.7]),
$\mathbb{E}\big(f(X,Y)\mid X,Z\big)=g(X).$
Applying the tower property of conditional expectation yields (29).
For (30), note that
$\mathbb{E}\big(g(X)\mid Z\big)=\mathbb{E}[g(W)]=\mathbb{E}[f(W,Y)],$
since $W$ and $Y$ are independent.
Lemma 3.17.
Let $1\leq m_{2}\leq m_{1}\leq q_{n}(j)$ and $k_{1},\ldots,k_{m_{2}}\in[1,q_{n}(j-1)]$ be distinct. Then on $\{X_{j}=m_{1}\}$,
$\mathbb{P}_{n}\big(\mathcal{X}_{j}=\{x^{j-1}_{k_{1}},\ldots,x^{j-1}_{k_{m_{2}}}\}\,\big|\,\mathcal{G}_{j}\big)=\binom{q_{n}(j-1)}{m_{2}}^{-1}{\mathbb{P}}_{n}\big(\#\{\mathfrak{p}(u^{j-1}_{i}):1\leq i\leq m_{1}\}=m_{2}\big).$ (31)
Proof 3.18.
We shall apply Lemma 3.15 to
$\begin{aligned} &\mathbb{P}_{n}\big(\mathcal{X}_{j}=\{x^{j-1}_{k_{1}},\ldots,x^{j-1}_{k_{m_{2}}}\}\,\big|\,\mathcal{G}_{j}\big)\\ =\;&{\mathbb{P}}_{n}\Big(\big\{\mathfrak{p}(x):x\in\mathcal{X}_{j}\big\}=\{x^{j-1}_{k_{1}},\ldots,x^{j-1}_{k_{m_{2}}}\}\,\big|\,\big(\mathscr{T}^{n}(v):|v|=j\big)\Big).\end{aligned}$ (32)
Observe the following:
  • (a)
    Both $\mathcal{X}_{j}$ and $\big(\mathscr{T}^{n}(v):|v|=j\big)$ are measurable with respect to
    $\sigma\big(\nu^{j},\nu^{j+1}\ldots,\nu^{h^{*}}\big),$
    and are independent of $\nu^{j-1}$ due to generation-wise independence of $\nu^{s}$.
  • (b)
    The event $\big\{\mathfrak{p}(x):x\in\mathcal{X}_{j}\big\}=\{x^{j-1}_{k_{1}},\ldots,x^{j-1}_{k_{m_{2}}}\}$ is determined by $\mathcal{X}_{j}$ and $\nu^{j-1}$, i.e.,
    $\Big\{\big\{\mathfrak{p}(x):x\in\mathcal{X}_{j}\big\}=\{x^{j-1}_{k_{1}},\ldots,x^{j-1}_{k_{m_{2}}}\}\Big\}\in\sigma(\mathcal{X}_{j},\nu^{j-1}).$
  • (c)
    By Proposition 3.12, on $\{X_{j}=m_{1}\}$,
    $\big(\mathcal{X}_{j}\mid\big(\mathscr{T}^{n}(v):|v|=j\big)\big)\overset{d}{=}\{u^{j}_{1},\ldots,u^{j}_{m_{1}}\}.$
Applying Lemma 3.15 with
$X=\mathcal{X}_{j},\quad Y=\nu^{j-1},\quad Z=\big(\mathscr{T}^{n}(v):|v|=j\big),\quad W=\{u^{j}_{1},\ldots,u^{j}_{m_{1}}\},$
and letting $f(X,Y)$ be the function determined by $\mathcal{X}_{j-1}$ and $\nu^{d_{j}}$ corresponding to the event
$\Big\{\big\{\mathfrak{p}(x):x\in\mathcal{X}_{j}\big\}=\{x^{j-1}_{k_{1}},\ldots,x^{j-1}_{k_{m_{2}}}\}\Big\},$
we obtain that on $\{X_{j}=m_{1}\}$,
${\mathbb{P}}_{n}\Big(\big\{\mathfrak{p}(x):x\in\mathcal{X}_{j}\big\}=\{x^{j-1}_{k_{1}},\ldots,x^{j-1}_{k_{m_{2}}}\}\,\big|\,\big(\mathscr{T}^{n}(v):|v|=j\big)\Big)$ (33)
$=$ ${\mathbb{P}}_{n}\big(\{\mathfrak{p}(u^{j}_{i}):1\leq i\leq m_{1}\}=\{x^{j-1}_{k_{1}},\ldots,x^{j-1}_{k_{m_{2}}}\}\big)$
$=$ $\binom{q_{n}(j-1)}{m_{2}}^{-1}{\mathbb{P}}_{n}\big(\#\{\mathfrak{p}(u^{j}_{i}):1\leq i\leq m_{1}\}=m_{2}\big).$
Substituting back into (32) gives the desired conclusion.
Proof 3.19 (Proof of Theorem 3.14).
It follows from (31) that on $\{X_{j}=m_{1}\}$,
${\mathbb{P}}_{n}\big(X_{j-1}=m_{2}\,\big|\,\mathcal{G}_{j}\big)$ $=\sum_{1\leq k_{1}<\cdots<k_{m_{1}}\leq q_{n}(j-1)}\mathbb{P}_{n}\big(\mathcal{X}_{j}=\{x^{j-1}_{k_{1}},\ldots,x^{j-1}_{k_{m_{1}}}\}\,\big|\,\mathcal{G}_{j}\big)$
$={\mathbb{P}}_{n}\!\left(\#\{\mathfrak{p}(u^{j}_{i}):1\leq i\leq m_{1}\}=m_{2}\right).$
This confirms the Markov property and the transition formula (28).
We close this section by a simple estimate.
Lemma 3.20.
Assume (H1). Then for any $\delta>0$, uniformly for $1\leq k\leq q_{n}(h^{*})$, $j\in[\delta n,\,h(q_{n})-\delta n]$ and $1\leq m\leq k$,
${\mathbb{P}}_{n}\big(X^{(k)}_{j-1}\leq m-1\,\big|\,X^{(k)}_{j}=m\big)\lesssim\frac{m^{2}}{n}.$ (34)
Proof 3.21.
By Corollary 3.3 and Theorem 3.14,
${\mathbb{P}}_{n}\big(X^{(k)}_{j-1}\leq m-1\,\big|\,X^{(k)}_{j}=m\big)={\mathbb{P}}_{n}\Big(\#\{\mathfrak{p}(u^{j}_{i}):1\leq i\leq m\}\leq m-1\Big)$
$\leq\binom{m}{2}\,{\mathbb{P}}_{n}\big(\mathfrak{p}(u^{j}_{1})=\mathfrak{p}(u^{j}_{2})\big)=\binom{m}{2}\,\frac{q_{n}(j-1)}{(q_{n}(j))_{2}}\,{\mathbb{E}}_{n}\big[(\nu_{1}^{j})_{2}\big].$
By (H1), this is $O(m^{2}/n)$, yielding (34).
3.3 Coming-down-from-infinity property
In this subsection, we establish the coming-down-from-infinity property for the coalescent process $X$ under the condition (H2’) introduced below.
3.3.1 Condition (H2’)
We now introduce the condition required to establish the coming-down-from-infinity property. Define for $n\geq 1$, $0\leq s\leq h(q_{n})$ and $1\leq m\leq q_{n}(s)$,
$p_{n}(m,s):={\mathbb{P}}_{n}\!\left(\#\{\mathfrak{p}(u^{s}_{1}),\ldots,\mathfrak{p}(u^{s}_{m})\}\leq m-1\right).$
By Theorem 3.14, this quantity also equals ${\mathbb{P}}_{n}(X_{s-1}\leq m-1\,|\,X_{s}=m)$ for any coalescent process $X$ based at a height strictly greater than $s$.
The condition is formulated as follows:
  1. (H2’)
    For every $\delta>0$, there exists $\varepsilon>0$ such that uniformly for all $n\geq 1$ and $\delta n\leq s\leq h(q_{n})-\delta n$,
    $p_{n}\!\left(\lfloor(\log n)^{1+\varepsilon}\rfloor,s\right)\gtrsim\frac{(\log n)^{2(1+\varepsilon)}}{n}.$ (35)
Remark 3.22.
Condition (H2’) provides a lower bound of the form
$p_{n}(m,s)\gtrsim\frac{m}{n^{2}}$ (36)
when $m=\lfloor(\log n)^{1+\varepsilon}\rfloor$. As will be shown in Lemma 3.23, in order for (H2’) to hold, it suffices that for each $n\geq 1$, there exists some $m^{\prime}\geq(\log n)^{1+\varepsilon}$ for which the bound (36) holds; the bound then automatically propagates to all smaller $m\leq m^{\prime}$.
The following lemma formalizes this monotonicity property and will be used to control the coalescence rate in the proof of the coming-down-from-infinity property.
Lemma 3.23.
For any $n\geq 1$, $1\leq s\leq h(q_{n})$, and $1\leq m^{\prime}\leq m\leq q_{n}(s)$,
$p_{n}(m^{\prime},s)\;\geq\;\frac{(m^{\prime})_{2}}{(m)_{2}}\,p_{n}(m,s).$
Proof 3.24.
Recall the notion of coalescent event from (13). Using this, we have the disjoint decomposition:
$\big\{\#\{\mathfrak{p}(u^{s}_{1}),\ldots,\mathfrak{p}(u^{s}_{m})\}\leq m-1\big\}=\bigcup_{\begin{subarray}{c}f\in M^{s}:\text{Dom}(f)=\{1,\ldots,m\}\\ \text{Range}(f)\leq m-1\end{subarray}}\text{Coal}^{s}(f).$
On any event $\mathrm{Coal}^{s}(f)$ appearing above, there is at least one collision of parents: there exist indices $i<j\leq m$ such that $\mathfrak{p}(u^{s}_{i})=\mathfrak{p}(u^{s}_{j})$. Note that conditionally on $\mathrm{Coal}^{s}(f)$, the $m^{\prime}$ individuals $\{u^{s}_{1},\ldots,u^{s}_{m^{\prime}}\}$ are obtained by choosing an $m^{\prime}$-subset uniformly from $\{u^{s}_{1},\ldots,u^{s}_{m}\}$. Thus the conditional probability that the colliding pair $\{i,j\}$ is entirely contained in $\{1,\ldots,m^{\prime}\}$ is $\frac{(m^{\prime})_{2}}{(m)_{2}}$. Consequently,
$\mathbb{P}_{n}\big(\#\{\mathfrak{p}(u^{s}_{1}),\ldots,\mathfrak{p}(u^{s}_{m})\}\leq m-1\,|\,\text{Coal}^{s}(f)\big)\geq\frac{(m^{\prime})_{2}}{(m)_{2}}.$
This leads to the conclusion.
Corollary 3.25.
Assume (H2’). For every $\delta>0$, there exists $\varepsilon>0$ such that uniformly for all $n\geq 1$, $\delta n\leq s\leq h(q_{n})-\delta n$ and $2\leq m\leq(\log n)^{1+\varepsilon}$,
$p_{n}\!\left(m,s\right)\gtrsim\frac{m^{2}}{n}.$ (37)
3.3.2 Coming-down-from-infinity property
Fix two sequences $h^{*}(n)$ and $h_{*}(n)$, $n\geq 1$, such that
$0<h_{*}(n)<h^{*}(n)<h(q_{n}),\qquad n\geq 1,$
and assume there exist constants $0<H_{*}<H^{*}<h(\ell)$ such that
$\frac{h_{*}(n)}{n}\to H_{*}\quad\text{and}\quad\frac{h^{*}(n)}{n}\to H^{*}.$ (38)
In particular, (38) implies that there exists $\delta>0$ such that, for all $n\geq 1$,
$\delta n\leq h_{*}(n)<h^{*}(n)\leq h(q_{n})-\delta n\quad\text{and}\quad h^{*}(n)-h_{*}(n)\geq\delta n.$ (39)
We consider the coalescent process $X_{j}=X_{j}(n,h^{*}(n))$ based at height $h^{*}(n)$ and focus on its value at height $h_{*}(n)$:
$N_{n}:=X_{h_{*}(n)}(n,h^{*}(n)).$
For brevity, we write $h^{*}=h^{*}(n)$ and $h_{*}=h_{*}(n)$ when $n$ is clear from the context.
Theorem 3.26 (Coming-down-from-infinity property).
Assume (H1) and (H2’). Then the sequence $(r_{n})$ is tight, i.e.,
$\lim_{M\to\infty}\,\limsup_{n\to\infty}\,{\mathbb{P}}_{n}(N_{n}>M)=0.$
Remark 3.27.
This can be regarded as a sequential version of the coming-down-from-infinity property, since the bound $M$ must be chosen independently of $n$.
Proof 3.28.
For each $M\geq 1$, define the hitting time
$T_{M}:=T_{M}(n):=\sup\{0\leq j\leq h^{*}(n):X_{j}\leq M\}.$
Clearly,
$\{N_{n}>M\}=\{T_{M}<h_{*}\}\subset\{T_{M}\leq h^{*}-\delta n\}.$
Thus, it suffices to prove
$\lim_{M\to\infty}\limsup_{n\to\infty}{\mathbb{P}}_{n}(T_{M}\leq h^{*}-\delta n)=0,$ (40)
by (39). Hence it is enough to show that for any $\eta>0$, there exists $M\geq 1$ such that, for all sufficiently large $n$,
${\mathbb{P}}_{n}\Big(T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\leq h^{*}-\tfrac{2}{3}\delta n\Big)<\tfrac{\eta}{2},$ (41)
${\mathbb{P}}_{n}\Big(T_{M}\leq T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}-\tfrac{1}{3}\delta n\,\Big|\,T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}>h^{*}-\tfrac{2}{3}\delta n\Big)<\tfrac{\eta}{2}.$ (42)
Proof 3.29 (Proof of (42)).
Define
$\zeta_{m}:=\begin{cases}\inf\{i\geq 1:X_{T_{m}-i}<m\},&\text{if }X_{T_{m}}=m,\\ 0,&\text{otherwise}.\end{cases}$
Then, for any $m_{1}<m_{2}$,
$T_{m_{2}}-T_{m_{1}}=\sum_{m=m_{1}+1}^{m_{2}}\zeta_{m}.$ (43)
By the strong Markov property, conditionally on $T_{m}=j$ and $X_{T_{m}}=m$,
$\zeta_{m}\overset{d}{=}\inf\{1\leq i\leq j:Y_{i}=1\},$ (44)
where $Y_{i}$ are independent Bernoulli random variables with success probability
${\mathbb{P}}(Y_{i}=1)={\mathbb{P}}_{n}(X_{j-i}\leq m-1\mid X_{j-i+1}=m).$
By (37), there exist $C^{*}>0$ such that for $n\geq 1$, $h_{*}\leq i\leq h^{*}$ and $2\leq m\leq\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor$,
${\mathbb{P}}(Y_{i}=1)\geq C^{*}\frac{m^{2}}{n}.$ (45)
We next record two auxiliary lemmas to be used in the proof.
Lemma 3.30.
For any $2\leq m<\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor$, conditionally on $T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}>h^{*}-\frac{2}{3}\delta n$, we have the stochastic bound
$\zeta_{m}\wedge\left(\frac{1}{3}\delta n-\sum_{j=m+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{j}\right)^{+}\;\leq_{\rm sto}\;{\rm Geometric}(C^{*}m^{2}/n).$ (46)
Lemma 3.31.
If $T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}-T_{M}=\sum_{m=M+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{m}\geq\tfrac{1}{3}\delta n$, then
$\sum_{m=M+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{m}\wedge\left(\frac{1}{3}\delta n-\sum_{j=m+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{j}\right)^{+}\geq\tfrac{1}{3}\delta n.$ (47)
Based on these lemmas, let $\gamma_{m}$ be independent ${\rm Geometric}(C^{*}m^{2}/n)$ variables. Then
[…]
Since
${\mathbb{P}}\Big(\sum_{m=M+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\gamma_{m}\geq\tfrac{1}{3}\delta n\Big)\;\leq\;\frac{3}{\delta n}{\mathbb{E}}\Big[\sum_{m=M+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\gamma_{m}\Big]\;\leq\;\frac{c}{M},$
for some constant $c>0$, substituting into (48) gives (42) once $M$ is chosen sufficiently large.
Proof 3.32 (Proof of Lemma 3.30).
It suffices to show that conditionally on $\big(X_{j}:T_{m}\leq j\leq h^{*}\big)$ with $T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}>h^{*}-\frac{2}{3}\delta n$, $X_{T_{m}}=m$ and $\frac{1}{3}\delta n-\sum_{j=m+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{j}>0$,
$\zeta_{m}\wedge\left(\frac{1}{3}\delta n-\sum_{j=m+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{j}\right)\;\leq_{\rm sto}\;{\rm Geometric}(C^{*}m^{2}/n).$ (49)
Note that under the conditioning, we have the following bound for the height $T_{m}$:
$T_{m}=T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}-\sum_{i=j+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{j}$ $>h^{*}-\delta n+\left(\frac{1}{3}\delta n-\sum_{j=m+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{j}\right)$
[…]
Then (49) follows from (44) and (45).
Proof 3.33 (Proof of Lemma 3.31).
Assume $\sum_{m=M+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{m}\geq\tfrac{1}{3}\delta n.$ Define
$m^{*}:=\sup\Big\{\,m\geq M+1:\ \sum_{j=m}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{j}\geq\tfrac{1}{3}\delta n\,\Big\}.$
Then by definition,
$\sum_{j=m^{*}}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{j}\geq\tfrac{1}{3}\delta n\quad\text{and}\quad\sum_{j=m}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{j}<\tfrac{1}{3}\delta n\ \text{for all }m>m^{*}.$
Therefore
$\sum_{m=m^{*}}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{m}\wedge\left(\frac{1}{3}\delta n-\sum_{j=m+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{j}\right)^{+}$ $=\sum_{m=m^{*}}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{m}+\left(\frac{1}{3}\delta n-\sum_{j=m^{*}+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{j}\right)^{+}$
$\geq\frac{1}{3}\delta n.$
This proves (47).
We now turn to the proof of (41). The main technical estimate we require is the following proposition.
Proposition 3.34.
Assume (H1) and (H2’). Then there exist constants $c>0$ such that for any $n\geq 1$, $h_{*}+\frac{n}{(\log n)^{1+\frac{\varepsilon}{4}}}\leq s\leq h^{*}$ and $1\leq k\leq q_{n}(j)$ and , we have
${\mathbb{P}}_{n}\left(X_{s-{\frac{n}{(\log n)^{1+\frac{\varepsilon}{4}}}}}\leq\dfrac{k}{(\log n)^{\frac{\varepsilon}{6}}}+(\log n)^{1+\frac{\varepsilon}{2}}\,\Big|\,X_{s}=k\right)\geq 1-e^{-c(\log n)^{\frac{\varepsilon}{4}}}.$ (50)
Proof 3.35 (Proof of (41) assuming Proposition 3.34).
Set
$l_{i}=h^{*}-\dfrac{n\cdot i}{(\log n)^{1+\frac{\varepsilon}{4}}}$
for $i=0,\cdots,\lfloor\frac{2}{3}\delta(\log n)^{1+\frac{\varepsilon}{4}}\rfloor$. We use the following elementary fact:
Lemma 3.36.
For sufficiently large $n$, we have if
$X_{l_{i+1}}\leq\dfrac{X_{l_{i}}}{(\log n)^{\frac{\varepsilon}{6}}}+(\log n)^{1+\frac{\varepsilon}{2}},$
for all $i=0,\ldots,\lfloor\frac{2}{3}\delta(\log n)^{1+\frac{\varepsilon}{4}}\rfloor$, then $X_{\lfloor\frac{2}{3}\delta n\rfloor}<2(\log n)^{1+\frac{\varepsilon}{2}}$.
Applying this, we obtain for $n$ large enough,
[…]
which goes to 0 as $n\rightarrow\infty$. This proves (41).
The remaining part is devoted to the proof of Proposition 3.34. Let $r=r(n):=\lfloor(\log n)^{1+\varepsilon}\rfloor.$ Fixing $h_{*}+\frac{n}{(\log n)^{1+\frac{\varepsilon}{4}}}\leq s\leq h^{*}$, consider the coalescent process
$X^{(r)}_{j}=X^{(r)}_{j}(n,s),$
and define the hitting time
$T^{(r)}_{m}:=\sup\{0\leq j\leq s:X^{(r)}_{j}\leq m\}.$
We begin with a coming-down-from-infinity type estimate.
Lemma 3.37.
Assume (H1) and (H2’). Then there exists constant $c>0$ such that for all $n\geq 1$ and $h_{*}+\frac{n}{(\log n)^{1+\frac{\varepsilon}{4}}}\leq s\leq h^{*}$,
${\mathbb{P}}_{n}\!\left(T^{(r)}_{\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}>s-\frac{n}{\lfloor(\log n)^{1+\frac{\varepsilon}{4}}\rfloor}\right)\;\geq\;1-e^{-c(\log n)^{\frac{\varepsilon}{4}}}.$
Proof 3.38.
By the same argument as in the proof of (42), we have with $\gamma_{m}$ independent $\sim{\rm Geometric}(C^{*}m^{2}/n)$,
${\mathbb{P}}_{n}\!\left(T^{(r)}_{\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\geq\frac{n}{\lfloor(\log n)^{1+\frac{\varepsilon}{4}}\rfloor}\right)\;\leq\;{\mathbb{P}}\!\left(\sum_{m=\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor+1}^{r}\gamma_{m}\;\geq\;\frac{n}{\lfloor(\log n)^{1+\frac{\varepsilon}{4}}\rfloor}\right).$ (51)
Since
${\mathbb{E}}\!\left[\sum_{m=\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor+1}^{r}\gamma_{m}\right]=\sum_{m=\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor+1}^{r}\frac{n}{C^{*}m^{2}}<\frac{n}{C^{*}\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor},$
we obtain
RHS of (51) $\leq{\mathbb{P}}\!\left(\exists\,m=\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor+1,\ldots,r,\,\ \gamma_{m}>C^{*}(\log n)^{\frac{\varepsilon}{4}}\,{\mathbb{E}}[\gamma_{m}]\right)$
$\leq\sum_{m=\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor+1}^{r}{\mathbb{P}}\!\left(\gamma_{m}>C^{*}(\log n)^{\frac{\varepsilon}{4}}\,{\mathbb{E}}[\gamma_{m}]\right)$
$\leq r\,e^{-C^{*}(\log n)^{\frac{\varepsilon}{4}}}\;\leq\;e^{-c(\log n)^{\frac{\varepsilon}{4}}}$
for some constant $c>0$, completing the proof.
Proof 3.39 (Proof of Proposition 3.34).
For $1\leq k\leq(\log n)^{1+\varepsilon}$, the bound (50) follows from Lemma 3.37. We therefore focus on the case $k>(\log n)^{1+\varepsilon}$.
Fix $h_{*}+\frac{n}{(\log n)^{1+\frac{\varepsilon}{4}}}\leq s\leq h^{*}$, denote
$j^{*}=j^{*}(s,n):=s-\frac{n}{(\log n)^{1+\frac{\varepsilon}{4}}}.$
By Theorem 3.14, it suffices to prove that
${\mathbb{P}}_{n}\left(X^{(k)}_{s-{\frac{n}{(\log n)^{1+\frac{\varepsilon}{4}}}}}\leq\dfrac{k}{(\log n)^{\frac{\varepsilon}{6}}}+(\log n)^{1+\frac{\varepsilon}{2}}\right)\geq 1-e^{-c(\log n)^{\frac{\varepsilon}{4}}}.$ (52)
[…]
where $\pi^{n}$ is a uniform permutation in the symmetric group $S_{q_{n}(s)}$. Let $\pi^{n}_{(1)}<\cdots<\pi^{n}_{(r)}$ denote the order statistics of $\{\pi^{n}_{1},\ldots,\pi^{n}_{r}\}$.
Observe that for any $1\leq m\leq r$, the event $\{X^{(r)}_{j^{*}}=m\}$ means that there exist indices
$0=i_{0}<i_{1}<\cdots<i_{m}=r$
and $m$ distinct vertices $v_{1},\ldots,v_{m}\in\mathcal{X}^{(r)}_{j^{*}}$ such that for each $q=1,\ldots,m$ and every $t\in\{i_{q-1}+1,\ldots,i_{q}\}$,
$\text{Anc}(x^{s}_{\pi^{n}_{(t)}},\,j^{*})=v_{q}.$
That is, the $r$ sampled vertices induce $m$ distinct ancestral lineages at height $j^{*}$, partitioning the index set $\{\pi^{n}_{(1)},\ldots,\pi^{n}_{(r)}\}$ into $m$ consecutive blocks
$I_{q}:=\{\pi^{n}_{(i_{q-1}+1)},\ldots,\pi^{n}_{(i_{q})}\},\qquad q=1,\ldots,m,$
such that all samples in $I_{q}$ share the same ancestor $v_{q}$ at height $j^{*}$.
On the event $\{X^{(r)}_{j^{*}}=m\}$, this gives a natural decomposition of all vertices at height $s$:
$B_{1}$ $:=\{x^{s}_{i}:\text{Anc}(x^{s}_{i},j^{*})\in\{v_{1},\ldots,v_{m}\}\}=\{x^{s}_{i}:i\in\cup_{q=1}^{m}I_{q}\},$
$B_{2}$ $:=\{x^{s}_{1},\ldots,x^{s}_{q_{n}(s)}\}\setminus B_{1}.$
We now return to the coalescent process $X^{(k)}$. Recall that
$X^{(k)}_{j^{*}}=\#\{\text{Anc}(u^{s}_{i},j^{*}):1\leq i\leq k\}.$
By splitting the $k$ sampled vertices according to whether $u^{s}_{i}\in B_{1}$ or $u^{s}_{i}\in B_{2}$, we obtain
$X^{(k)}_{j^{*}}\leq X^{(r)}_{j^{*}}+\#\big(B_{2}\cap\{u^{s}_{i}:1\leq i\leq k\}\big).$ (53)
It follows from Lemma 3.37 that
$\mathbb{P}_{n}\left(X^{(r)}_{j^{*}}\leq\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor\right)\geq{\mathbb{P}}_{n}\!\left(T^{(r)}_{\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}>s-\frac{n}{\lfloor(\log n)^{1+\frac{\varepsilon}{4}}\rfloor}\right)\geq 1-e^{-c(\log n)^{\frac{\varepsilon}{4}}}.$ (54)
The next estimate, whose proof will be given later, is the key to completing the argument.
Lemma 3.40.
There exists constant $c^{\prime}>0$ such that for any $n\geq 1$, $h_{*}+\frac{n}{(\log n)^{1+\frac{\varepsilon}{4}}}\leq s\leq h^{*}$ and $(\log n)^{1+\varepsilon}\leq k\leq q_{n}(s)$,
$\mathbb{P}_{n}\left[\#\big(B_{2}\cap\{u^{s}_{i}:1\leq i\leq k\}\big)>\frac{k}{(\log n)^{\varepsilon/6}}\right]\leq e^{-\frac{c^{\prime}k}{(\log n)^{\varepsilon/6}}}.$ (55)
Substituting (54) and (55) to (53) yields that
$\quad\,{\mathbb{P}}_{n}\left(X^{(k)}_{j^{*}}>\dfrac{k}{(\log n)^{\frac{\varepsilon}{6}}}+(\log n)^{1+\frac{\varepsilon}{2}}\right)$
$\leq\mathbb{P}_{n}\left(X^{(r)}_{j^{*}}>\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor\right)+\mathbb{P}_{n}\left[\#\big(B_{2}\cap\{u^{s}_{i}:1\leq i\leq k\}\big)>\frac{k}{(\log n)^{\varepsilon/6}}\right]$
[…]
for some constant $c^{\prime\prime}>0$. This completes the proof.
Proof 3.41 (Proof of Lemma 3.40).
By (54), it suffices to show that for any $1\leq m\leq\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor$,
${\mathbb{P}}_{n}\!\left(\#\big(B_{2}\cap\{u^{s}_{i}:1\leq i\leq k\}\big)>\frac{k}{(\log n)^{\varepsilon/6}}\,\Big|\,X^{(r)}_{j^{*}}=m\right)\leq e^{-\frac{c^{\prime}k}{(\log n)^{\varepsilon/6}}}.$ (56)
We use the following elementary property of uniform permutations.
Lemma 3.42.
Let $\pi^{n}_{(0)}:=0$ and $\pi^{n}_{(r+1)}:=q_{n}(s)$. Then there exists a constant $c^{\prime}>0$ such that for all $n\geq 1$ and $h_{*}+\frac{n}{(\log n)^{1+\frac{\varepsilon}{4}}}\leq s\leq h^{*}$,
${\mathbb{P}}\!\left(\max_{0\leq i\leq r}(\pi^{n}_{(i+1)}-\pi^{n}_{(i)})\leq\frac{n}{(\log n)^{1+\frac{3\varepsilon}{4}}}\right)\geq 1-e^{-c^{\prime}(\log n)^{\frac{\varepsilon}{4}}}.$
On the event $\{X^{(r)}_{j^{*}}=m\}$,
[…]
and hence for any $1\leq m\leq\lfloor(\log n)^{1+\frac{\varepsilon}{2}}\rfloor$,
${\mathbb{P}}_{n}\!\left(\#B_{2}\leq\frac{n}{(\log n)^{\frac{\varepsilon}{4}}}\,\Big|\,X^{(r)}_{j^{*}}=m\right)\geq 1-e^{-c^{\prime}(\log n)^{\frac{\varepsilon}{4}}}.$ (57)
Finally, note that $(u^{s}_{1},\ldots,u^{s}_{r})\subset B_{1}$, while the remaining $(u^{s}_{r+1},\ldots,u^{s}_{k})$ are chosen uniformly among the other $q_{n}(s)-r$ vertices at height $s$. Conditionally on $B_{2}$, the random variable $(B_{2}\cap\{u^{s}_{i}:1\leq i\leq k\})$ follows a hypergeometric distribution with population size $q_{n}(s)$, number of successes $B_{2}$ and sample size $k$. Combining the tail estimate of the hypergeometric distribution with (57) yields (56).
4 Tightness
Recall that $\widetilde{\mathscr{T}}^{n}$ is the geometric tree obtained from $\mathscr{T}^{n}$ by assigning length $1/n$ to each edge. In this section, we prove the tightness (Theorem 4.1) of the sequence $(\widetilde{\mathscr{T}}^{n})_{n\in\mathbb{N}}$, which is the main technical step in establishing Theorem 1.1.
Recall the definition of $\Delta(n,k)$ from (5). Throughout this section, we use this notation specifically for the sequence $(\mathscr{T}^{n})$, i.e.,
$\Delta(n,k):=\max_{v\in V(\mathscr{T}^{n})}\min_{u\in V(\mathscr{T}^{n}_{k})}d(v,u),\quad\text{for }1\leq k\leq\#V(\mathscr{T}^{n}),$
where $\mathscr{T}^{n}_{k}=\mathscr{T}^{n}(V_{n,1},\dots,V_{n,k})$ is the subtree of $\mathscr{T}^{n}$ spanned by $(V_{n,1},\dots,V_{n,k})$ and root $\rho$, with $(V_{n,1},\dots,V_{n,\#V(\mathscr{T}^{n})})$ being a uniform ordering of $V(\mathscr{T}^{n})$.
Theorem 4.1.
Assume (H1),(H2’) and (H2”). Then the sequence $\widetilde{\mathscr{T}}^{n}$ is tight. Equivalently, for any $\delta>0$,
$\lim_{k\rightarrow\infty}\limsup_{n\rightarrow\infty}{\mathbb{P}}_{n}\left(\Delta(n,k)>\delta n\right)=0.$
4.1 Strategy of the proof
We begin by introducing the notion of $\delta$-coalescent vertices, which plays a central role in the argument.
Definition 4.2 ($\delta$-coalescent vertices).
For $n\geq 1$ and $\delta>0$, define
$H=H(n,\delta):=\left\lfloor\frac{h(q_{n})}{\lfloor\delta n\rfloor}\right\rfloor.$ (58)
The interval $\big[s\lfloor\delta n\rfloor,(s+1)\lfloor\delta n\rfloor\big)$ is called the $s$-th height interval. We focus on the coalescent process starting from height $(s+1)\lfloor\delta n\rfloor$ for $s=1,\dots,H-2$. Recall that $\mathcal{X}_{s\lfloor\delta n\rfloor}\big(n,(s+1)\lfloor\delta n\rfloor\big)$ consists of all vertices at height $s\lfloor\delta n\rfloor$ that have at least one descendant at height $(s+1)\lfloor\delta n\rfloor$. Those vertices are referred to as the $\delta$-coalescent vertices at the $s$-th height interval. The collection of all $\delta$-coalescent vertices over these height intervals is denoted by
$\bar{\mathcal{X}}(n,\delta):=\bigcup_{s=1}^{H-2}\mathcal{X}_{s\lfloor\delta n\rfloor}\big(n,(s+1)\lfloor\delta n\rfloor\big).$
These $\delta$-coalescent vertices serve as structural anchors within the tree, ensuring that every vertex lies within a distance of at most $3\lfloor\delta n\rfloor$ from one of these vertices. (In general, the distance bound is $\lfloor\delta n\rfloor$, but it increases to $2\lfloor\delta n\rfloor$ near the root since $s=0$ is not considered, and up to $3\lfloor\delta n\rfloor$ near height $h(q_{n})$ as s=K1s=K-1 is also excluded.)
To establish the tightness of the sequence $\widetilde{\mathscr{T}}^{n}$, we need to show that the maximum distance $\Delta(n,k)$ divided by $n$ between any vertex $v$ in $\mathscr{T}^{n}$ and its closest vertex in the subtree $\mathscr{T}^{n}_{k}$ vanishes as $k\to\infty$. This is achieved by proving two key results:
  • (1)
    $\bar{\mathcal{X}}(n,\delta)$ is tight (see Proposition 4.3);
  • (2)
    each $\delta$-coalescent vertex has at least $\Theta(n^{2})$ descendants within distance $\lfloor\delta n\rfloor$ (see Proposition 4.5).
These two results imply that, as $k\to\infty$, with high probability, every $\delta$-coalescent vertex will have a descendant among the $k$ selected vertices $\{V_{n,1},\ldots,V_{n,k}\}$ within its $\lfloor\delta n\rfloor$-neighborhood. Consequently, $\frac{1}{n}\Delta(n,k)$ remains small as $n\to\infty$ and $k\to\infty$.
We now formulate the two results as propositions.
Proposition 4.3.
Assume (H1) and (H2’). Then for any $\delta>0$, the number of $\delta$-coalescent vertices, $\bar{\mathcal{X}}(n,\delta)$, is tight, i.e.,
$\lim_{M\rightarrow\infty}\limsup_{n\rightarrow\infty}{\mathbb{P}}_{n}\big(\#\bar{\mathcal{X}}(n,\delta)>M\big)=0.$
Proof 4.4.
By (H1), the quantity $H=H(n,\delta)$ converges as $n\rightarrow\infty$. Therefore, it suffices to prove that for any $\delta>0$ and $1\leq s\leq H-2$,
$\lim_{M\to\infty}\limsup_{n\to\infty}{\mathbb{P}}_{n}\Big(X_{s\lfloor\delta n\rfloor}\big(n,(s+1)\lfloor\delta n\rfloor\big)>M\Big)=0.$
This follows directly from Theorem 3.26, with $h^{*}=(s+1)\lfloor\delta n\rfloor$ and $h_{*}=s\lfloor\delta n\rfloor$.
Recall that for $v\in\mathscr{T}^{n}$, we write $\text{Desc}(v)$ for the set of its descendants. For any $\gamma>0$, define $B(n,\gamma,\delta)$ as the event that every $\delta$-coalescent vertex has at least $\gamma n^{2}$ descendants within graph distance $\delta n$ in $\mathscr{T}^{n}$. That is,
$B(n,\gamma,\delta):=\Big\{\forall\,v\in\bar{\mathcal{X}}(n,\delta):\ \#\{u\in\text{Desc}(v):d(u,v)<\delta n\}\geq\gamma n^{2}\Big\}.$ (59)
Proposition 4.5.
Assume (H1) and (H2’). Then
$\lim_{\gamma\to 0+}\,\liminf_{n\to\infty}{\mathbb{P}}_{n}\big(B(n,\gamma,\delta)\big)=1.$
We defer the proof of Proposition 4.5.
4.2 Proof of Theorem 4.1 assuming Proposition 4.5
Define the event
$A(n,k,\delta):=\big\{\forall\,v\in\bar{\mathcal{X}}(n,\delta),\ \exists\,1\leq i\leq k\text{ such that }V_{n,i}\in\text{Desc}(v)\text{ and }d(v,V_{n,i})\leq\delta n\big\}.$
By (2), there exists constant $C>0$ such that
$V(\mathscr{T}^{n})\leq Cn^{2}.$
It follows that for any $M\geq 1$ and $\gamma>0$,
${\mathbb{P}}_{n}\Big(A(n,k,\delta)\,\big|\,\{\#\bar{\mathcal{X}}(n,\delta)\leq M\}\cap B(n,\gamma,\delta)\Big)\geq 1-M\big(1-\gamma/C\big)^{k}.$
Applying Propositions 4.3 and 4.5, we conclude that
$\lim_{k\rightarrow\infty}\liminf_{n\rightarrow\infty}{\mathbb{P}}_{n}\big(A(n,k,\delta)\big)=1.$
The proof is then completed by the following key observation:
$A(n,k,\delta)\subset\big\{\Delta(n,k)\leq 4\delta n\big\}.$ (60)
Indeed, since the root $v_{0}$ is a vertex in $\mathscr{T}^{n}_{k}$, for any vertex $v$ with $\left\lvert v\right\rvert\in\big[0,4\delta n\big]$, we have
$\min_{u\in V(\mathscr{T}^{n}_{k})}d(v,u)\leq d(v,v_{0})\leq 4\delta n.$
For all vertex $v$ with $\left\lvert v\right\rvert\in\big(s\delta n,(s+1)\delta n\big]$ for some $s=4,\dots,H$, let $v^{*}$ be its ancestor at height $(s-2)\lfloor\delta n\rfloor$. By definition, $v^{*}\in\bar{\mathcal{X}}(n,\delta)$. Moreover, on the event $A(n,k,\delta)$, there exists $1\leq i_{0}\leq k$ such that $V_{n,i_{0}}\in\text{Desc}(v^{*})$ and $d(v^{*},V_{n,i_{0}})<\delta n$. Consequently,
$\min_{u\in V(\mathscr{T}^{n}_{k})}d(v,u)\leq d(v,V_{n,i_{0}})\leq d(v,v^{*})+d(v^{*},V_{n,i_{0}})\leq 4\delta n,\quad\text{on }A(n,k,\delta).$
This verifies (60) and thus concludes the proof.
4.3 Proof of Proposition 4.5
Recall the definition of $B(n,\eta,\delta)$ from (59). For $s=1,\ldots,H-2$, define the event
$B(n,s,\gamma,\delta):=\Big\{\forall\,v\in X_{s\lfloor\delta n\rfloor}\big(n,(s+1)\lfloor\delta n\rfloor\big),\ \#\{u\in\text{Desc}(v):d(u,v)<\delta n\}\geq\gamma n^{2}\Big\}.$ (61)
This represents the “restriction” of $B(n,\gamma,\delta)$ to the $s$-th height interval. It suffices to fix $\delta>0$ and $1\leq s\leq H-2$, and show that for any $\eta>0$, there exists $\gamma=\gamma(\eta)>0$ such that for all $n\geq 1$,
${\mathbb{P}}_{n}\big(B(n,s,\gamma,\delta)\big)\geq 1-\eta.$ (62)
Our strategy is as follows. Recall the base height $h^{*}$ in the definitions of $\mathcal{X}_{j}=\mathcal{X}_{j}(n,h^{*})$ and $X_{j}=X_{j}(n,h^{*})$. In this subsection we fix
$h^{*}:=(s+1)\lfloor\delta n\rfloor$
and consider the coalescent process $(X_{j})_{0\leq j\leq\lfloor\delta n\rfloor}$ from this base height. We first analyze the behavior of the coalescent process $(X_{j})$ and identify a typical event that occurs with high probability and is more tractable for subsequent estimates. We then condition on this event and establish the desired bound for the probability of $B(n,s,\gamma,\delta)$.
A typical event
As preparation, we identify a typical event in the coalescent process, which will serve as the basis for our later estimates. By Theorem 3.26,
$\lim_{M\to\infty}\liminf_{n\to\infty}{\mathbb{P}}_{n}\Big(X_{h^{*}-\lfloor\frac{1}{2}\delta n\rfloor}\leq M\Big)=1.$
Hence, for any $\eta>0$, there exists $M_{0}=M_{0}(\eta)\geq 1$ such that
${\mathbb{P}}_{n}\!\Big(X_{h^{*}-\lfloor\frac{1}{2}\delta n\rfloor}\leq M_{0}\Big)\geq 1-\eta/4,\quad\text{for all }n\geq 1.$ (63)
Next, consider the holding time
$\bar{T}=\bar{T}(n,\delta):=\inf\Big\{j\geq 1:X_{h^{*}-\lfloor\frac{1}{2}\delta n\rfloor-j}<X_{h^{*}-\lfloor\frac{1}{2}\delta n\rfloor}\Big\}.$
Using the same argument as in (46) together with (34), we obtain:
Proposition 4.6.
There exists a constant $C=C(\delta,\varepsilon,M_{0})>0$ such that for any $1\leq m\leq M_{0}$, conditionally on $X_{h^{*}-\lfloor\frac{1}{2}\delta n\rfloor}=m$, we have
$\bar{T}\;\geq_{\rm sto}\;\text{Geometric}(C/n)\wedge\tfrac{1}{4}\delta n.$
Consequently, there exists $c_{0}=c_{0}(C,\delta)\in(0,\delta/2)$ such that for any $1\leq m\leq M_{0}$,
${\mathbb{P}}_{n}\!\big(\bar{T}>\lfloor c_{0}n\rfloor\,\big|\,X_{h^{*}-\lfloor\frac{1}{2}\delta n\rfloor}=m\big)\geq 1-\eta/4,\quad\text{for all }n\geq 1.$ (64)
Set
$\bar{h}:=(s+1)\lfloor\delta n\rfloor-\left\lfloor\frac{1}{2}\delta n\right\rfloor-\lfloor c_{0}n\rfloor.$
Combining (63) and (64), we arrive at the following high-probability event:
$\begin{aligned} &E(n,\delta,\eta,m):=\Big\{\bar{T}>\lfloor c_{0}n\rfloor,X_{h^{*}-\lfloor\frac{1}{2}\delta n\rfloor}=m\Big\}=\Big\{X_{k}=m,\ \forall\,k\in[\,\bar{h},\bar{h}+\lfloor c_{0}n\rfloor]\Big\},\\ &{\mathbb{P}}_{n}\bigg(\bigcup_{m=1}^{M_{0}}E(n,\delta,\eta,m)\bigg)\geq 1-\eta/2.\end{aligned}$ (65)
We shall take the event $E(n,\delta,\eta,m)$ as the typical event on which our subsequent analysis will be conditioned. Define for $1\leq m\leq M_{0}$ the conditional probability measure
${\mathbb{Q}}^{m}_{n}(\cdot):={\mathbb{P}}_{n}\big(\,\cdot\,\big|\,E(n,\delta,\eta,m)\big).$
Structure of $\mathscr{T}^{n}$ on $E(n,\delta,\eta,m)$
The purpose of this subsection is to set up the structural ingredients — coalescent subtrees, trunk paths and vertices — which will be crucial for the forthcoming estimates. Recall the subtree notation $\mathscr{T}^{n}(v)=\mathscr{T}^{n}(v,h^{*})$ from (26), which denotes the subtree of $\mathscr{T}^{n}$ rooted at $v$ and truncated at height $h^{*}$.
Definition 4.7 (Coalescent subtrees).
For $0\leq j\leq h^{*}$, define the collection of the coalescent subtrees at height $j$ by
$\mathbf{T}_{j}=\mathbf{T}_{j}(n,h^{*}):=\big\{\mathscr{T}^{n}(v):v\in\mathcal{X}_{j}\big\}.$
Each element of $\mathbf{T}_{j}$ is a subtree rooted at a vertex at height $j$ that admits at least one descendant at height $h^{*}$. Note that $|\mathbf{T}_{j}|=|\mathcal{X}_{j}|=X_{j}$.
We then write $\mathsf{T}_{1},\ldots,\mathsf{T}_{|\mathbf{T}_{j}|}$ for a uniform random ordering of the elements of $\mathbf{T}_{j}$. Concretely, conditionally on $\mathscr{T}^{n}$, we first list the elements of $\mathbf{T}_{j}$ according to the lexicographical order of their roots, denoted by $\mathbf{T}_{j}^{(1)},\ldots,\mathbf{T}_{j}^{(|\mathbf{T}_{j}|)}$. Then we choose a permutation $\pi$ uniformly at random from the symmetric group on $|\mathbf{T}_{j}|$ elements, and define
$(\mathsf{T}_{1},\ldots,\mathsf{T}_{|\mathbf{T}_{j}|}):=(\mathbf{T}_{j}^{(\pi(1))},\ldots,\mathbf{T}_{j}^{(\pi(|\mathbf{T}_{j}|))}).$
Finally, we introduce the labeling map
$\Lambda_{j}:\mathbf{T}_{j}\longrightarrow\{1,\ldots,|\mathbf{T}_{j}|\},\qquad\Lambda_{j}(\mathsf{T}_{i}):=i,$ (66)
which assigns to each coalescent subtree its index in the random ordering. The random orderings (and hence the labeling maps $\Lambda_{j}$) are taken independently across different heights $j$.
Definition 4.8 (Trunk paths and vertices).
Under ${\mathbb{Q}}^{m}_{n}$, we have
$X_{j}=m,\quad\text{for all }\bar{h}\leq j\leq\bar{h}+\lfloor c_{0}n\rfloor.$
In other words, for each $j$ between $\bar{h}$ and $\bar{h}+\lfloor c_{0}n\rfloor$, there are exactly $m$ vertices in $\mathcal{X}_{j}$. These vertices belong to $m$ distinct ancestral lineages that remain separate and do not coalesce.
Let $\mathsf{T}_{1},\ldots,\mathsf{T}_{m}$ be the $m$ coalescent subtrees at height $\bar{h}$. For $\bar{h}\leq j\leq\bar{h}+\lfloor c_{0}n\rfloor$ and $1\leq i\leq m$, denote by
$z^{j}_{i}\in V(\mathsf{T}_{i})$
the unique vertex of $\mathsf{T}_{i}$ at height $j$. The sequence
$Z_{i}:=\big(z^{\bar{h}+\lfloor c_{0}n\rfloor}_{i},z^{\bar{h}+\lfloor c_{0}n\rfloor-1}_{i},\ldots,z^{\bar{h}}_{i}\big)$
is called the trunk path of $\mathsf{T}_{i}$, and its elements are referred to as the trunk vertices of $\mathsf{T}_{i}$.
The remaining $q_{n}(j)-m$ non-trunk vertices at height $j$ are denoted by $y^{j}_{1},\ldots,y^{j}_{q_{n}(j)-m}$, where the subscripts are assigned arbitrary.
Merging vertices and key theorem
While the $m$ trunk paths remain disjoint, the lineages of the $y^{j}_{l}$ may merge into some $Z_{i}$, implying $y^{j}_{l}\in{\rm Desc}(z^{\bar{h}}_{i})$. To quantify this, define
$\mathcal{N}_{i}:=\sum_{j=\bar{h}+1}^{\bar{h}+\lfloor c_{0}n\rfloor}\sum_{\ell=1}^{q_{n}(j)-m}\mathbf{1}\big\{y^{j}_{l}\in{\rm Desc}(z^{\bar{h}}_{i})\big\},$
the total number of vertices whose lineages merge into $Z_{i}$. The following theorem provides a lower bound on $\mathcal{N}_{i}$.
Theorem 4.9.
Assume (H1) and (H2’). For fixed $\delta,s,\eta,M_{0},c_{0}$, there exists $\gamma>0$ such that for all $1\leq m\leq M_{0}$ and sufficiently large $n$,
${\mathbb{Q}}^{m}_{n}\big(\mathcal{N}_{i}\geq\gamma n^{2},\ \forall 1\leq i\leq m\big)\geq 1-\eta/2.$ (67)
Proof 4.10 (Proof of (62) assuming Theorem 4.9).
Each $\delta$-coalescent vertex at height $s\lfloor\delta n\rfloor$ has at least one descendant in $\{z^{0}_{1},\dots,z^{0}_{m}\}$. Hence, for all large $n$ and $1\leq m\leq M_{0}$,
${\mathbb{Q}}^{m}_{n}\big(B(n,s,\gamma,\delta)\big)\geq{\mathbb{Q}}^{m}_{n}\big(\mathcal{N}_{i}\geq\gamma n^{2},\ \forall 1\leq i\leq m\big)\geq 1-\eta/2.$
Combining with (65) completes the proof of (62).
4.4 Proof of Theorem 4.9
The remainder of this section is devoted to the proof of Theorem 4.9. Throughout, we fix $\delta,s,M_{0},c_{0}$, and regard them as constants (i.e. omit the dependence of other constants on $\delta,s,M_{0},c_{0}$).
4.4.1 Outline of the proof
We begin with a brief outline of the proof strategy.
The first step is to partition the interval $\big[\bar{h},\bar{h}+\lfloor c_{0}n\rfloor\big]$ into $K_{\theta}$ sub-intervals $(h_{r},h_{r-1}]$, where $\bar{h}+\lfloor c_{0}n\rfloor=h_{0}>h_{1}>\cdots>h_{K_{\theta}}$ and each sub-interval has length $\lfloor\theta n\rfloor$ (see (68) below for precise definitions). For each $r=0,\ldots,K_{\theta}-1$, let $\mathcal{N}_{i,r}$ be the number of $y^{j}_{l}$ in the height interval $(h_{r},h_{r-1}]$ that merge into $Z_{i}$ before height $h_{r}$.
We then establish the independence of the family $(\mathcal{N}_{i,r})$ under $\mathbb{Q}^{m}_{n}$ (Proposition 4.11) and obtain a uniform lower bound for their tail probabilities (Proposition 4.12). Since $\mathcal{N}_{i}$ exceeds the sum of $\mathcal{N}_{i,r}$, the event that $\mathcal{N}_{i}>\gamma n^{2}$ is implied by the event that some $\mathcal{N}_{i,r}>\gamma n^{2}$. The independence and uniform lower bound together yield the desired probability estimate in (67).
To prove the independence and the uniform lower bound, we introduce the notion of uniform selection pairs (Definition 4.14). Roughly speaking, a random vector $\mathcal{V}$ consisting of $k$ vertices of $\mathscr{T}^{n}$ at some height $h$, together with an event $A$ depending only on the genealogy above that height, forms a uniform selection pair if, conditionally on $A$, $\mathcal{V}$ is distributed as a uniformly chosen $k$-tuple of distinct vertices at height $h$. Such structures naturally appear in the events we consider.
For these events, we prove a general lemma (Lemma 4.15) showing that their conditional probabilities can be decomposed into contributions from the genealogies above and below a given height. This decomposition directly yields the desired independence property and simplifies the probability estimates required for proving the uniform lower bound.
Finally, the uniform lower bound is obtained via the second moment method, by relating the merging probabilities to the coalescent probabilities in Corollary 3.9 (see Proposition 4.19).
4.4.2 Key propositions
Let $\theta\in(0,c_{0})$. Define
$K_{\theta}:=\Big\lfloor\frac{\lfloor c_{0}n\rfloor}{\lfloor\theta n\rfloor}\Big\rfloor,\qquad h_{r}:=\bar{h}+\lfloor c_{0}n\rfloor-r\lfloor\theta n\rfloor,\quad 0\leq r\leq K_{\theta},$ (68)
and for $1\leq i\leq m$ and $1\leq r\leq K_{\theta}$, set
$\mathcal{N}_{i,r}:=\sum_{j=h_{r}+1}^{h_{r-1}}\sum_{l=1}^{q_{n}(j)-m}\mathbf{1}\big\{y^{j}_{l}\in{\rm Desc}(z_{i}^{h_{r}})\big\}.$
Clearly,
$\mathcal{N}_{i}\;\geq\;\sum_{r=1}^{K_{\theta}}\mathcal{N}_{i,r},$
with equality when $\lfloor c_{0}n\rfloor/\lfloor\theta n\rfloor\in{\mathbb{N}}$.
For later use, we introduce the quantity $\widetilde{\mathcal{N}}_{i,r}$, which serve as an analog of $\mathcal{N}_{i,r}$ for uniformly chosen vertices $(u^{h_{r-1}}_{1},\ldots,u^{h_{r-1}}_{m})$ and will naturally appear in subsequent arguments.
First, for any vertex $v$ and $0\leq j\leq|v|$, denote by $\text{Anc}(v,j)$ the ancestor of $v$ at height $j$. For $0\leq r\leq K_{\theta}$ and $1\leq i\leq m$, denote the ancestral lineage of $u^{h_{r-1}}_{i}$ up to height $h_{r}$ by
$U_{i,r}:=U_{i}\big([h_{r-1},h_{r}]\big):=\Big(u^{h_{r-1}}_{i},\text{Anc}(u^{h_{r-1}}_{i},h_{r-1}-1),\ldots,\text{Anc}(u^{h_{r-1}}_{i},h_{r})\Big).$
For each $j$ with $h_{r}<j\leq h_{r-1}$, denote the remaining $q_{n}(j)-m$ vertices at height $j$ by $v^{j}_{1},\ldots,v^{j}_{q_{n}(j)-m}$, where the subscripts are assigned, conditionally on $\mathscr{T}^{n}$, uniformly at random and independently across heights. (Unlike $(y^{j}_{l})$ before, we specify the distribution of subscripts to simplify later analysis.)
Then, for $1\leq i\leq m$ and $1\leq r\leq K_{\theta}$, set
$\widetilde{\mathcal{N}}_{i,r}:=\sum_{j=h_{r}+1}^{h_{r-1}}\sum_{l=1}^{q_{n}(j)-m}\mathbf{1}\Big\{v^{j}_{l}\in\text{Desc}\big(\text{Anc}(u^{h_{r-1}}_{i},h_{r})\big)\Big\}.$
That is, $\widetilde{\mathcal{N}}_{i,r}$ counts the number of vertices $v^{j}_{l}$ at heights $(h_{r},h_{r-1}]$ that are not part of any ancestral lineage $(U_{k,r}:k=1,\ldots,m)$ but eventually merge into $U_{i,r}$ at some height greater than or equal to $h_{r}$.
For $0\leq r\leq K_{\theta}$, we define the event
$\widetilde{B}_{r}:=\big\{U_{1,r},\ldots,U_{m,r}\text{ are disjoint}\big\},$
Proposition 4.11.
The random variables $(\mathcal{N}_{i,r}:1\leq r\leq K_{\theta})$ are independent under ${\mathbb{Q}}_{n}^{m}$. Moreover, for any $1\leq r\leq K_{\theta}$ and test function $f$ on $\mathbb{N}$,
$\mathbb{Q}^{m}_{n}\big(f(\mathcal{N}_{i,r})\big)=\mathbb{P}_{n}\left(f(\widetilde{\mathcal{N}}_{i,r})\,|\,\widetilde{B}_{r}\right)$ (69)
Proposition 4.12.
Assume (H1) and (H2’). Then there exists constant $C_{0}\in(0,1)$ such that for any $\theta\in(0,c_{0})$ sufficiently small, there exists $\gamma=\gamma(\theta,C_{0})>0$ satisfying
${\mathbb{Q}}_{n}^{m}(\mathcal{N}_{i,r}\geq\gamma n^{2})\;\geq\;C_{0},$
for all sufficiently large $n$, $1\leq i\leq m\leq M_{0}$ and $0\leq r\leq K_{\theta}-1$.
The proofs of Propositions 4.11 and 4.12 are postponed, where we will see what additional condition $\theta$ should satisfy. We first explain how these results imply Theorem 4.9.
Proof 4.13 (Proof of Theorem 4.9 assuming Propositions 4.11 and 4.12).
Observe that if $\mathcal{N}_{i,r}\geq\gamma n^{2}$ for some $1\leq r\leq K_{\theta}$, then $\mathcal{N}_{i}\geq\gamma n^{2}$. Hence, by Propositions 4.11 and 4.12, for any $\theta>0$, there exists $\gamma>0$ such that for sufficiently large $n$, $1\leq i\leq m\leq M_{0}$ and $0\leq r\leq K_{\theta}-1$,
${\mathbb{Q}}_{n}^{m}(\mathcal{N}_{i}\geq\gamma n^{2})\;\geq\;1-{\mathbb{Q}}_{n}^{m}\!\big(\mathcal{N}_{i,r}<\gamma n^{2},\ \forall\,1\leq r\leq K_{\theta}\big)\;\geq\;1-(1-C_{0})^{K_{\theta}}.$
Since $K_{\theta}\sim c_{0}/\theta$ as $n\to\infty$, for any $\eta>0$ one can choose $\theta=\theta(\eta)>0$ sufficiently small such that, for all $n\geq 1$,
$(1-C_{0})^{K_{\theta}}\leq\frac{\eta}{2M_{0}}.$
Thus, there exists $\gamma>0$ such that
[…]
Consequently,
${\mathbb{Q}}_{n}^{m}\big(\mathcal{N}_{i}\geq\gamma n^{2},\,\forall\,1\leq i\leq m\big)\;\geq\;1-\sum_{j=1}^{m}{\mathbb{Q}}_{n}^{m}\big(\mathcal{N}_{i}<\gamma n^{2}\big)\;\geq\;1-\frac{\eta}{2},$
which completes the proof.
4.4.3 Proof of Propositions 4.11 and 4.12
We begin with a preliminary lemma, which will be used repeatedly in the sequel. Recall the definitions of $\mathcal{G}_{j}$ in (27) and labeling map $\Lambda_{j}$ in (66).
Definition 4.14 (Uniform selection pair).
Fix a base height $h^{*}$. For $0\leq h\leq h^{*}$ and $1\leq k\leq q_{n}(h)$, let $(A,\mathcal{V})$ be a pair consisting of
  • an event $A$ measurable with respect to $\mathcal{G}_{h}\vee\sigma(\Lambda_{h})$, and
  • a random vector $\mathcal{V}$ of $k$ vertices taking values in $\{x^{h}_{1},\ldots,x^{h}_{q_{n}(h)}\}$.
We say that $(A,\mathcal{V})$ is a $(h,k)$-uniform selection pair if, under $\mathbb{P}_{n}$,
$(\mathcal{V}\mid A)\,\overset{d}{=}\,\{u^{h}_{1},\ldots,u^{h}_{k}\},$
that is, conditionally on $A$, the vector $\mathcal{V}$ has the same distribution as a uniformly chosen $k$-tuple of distinct vertices at height $h$.
Recall the notion of offspring vector $\nu^{s}$.
Lemma 4.15.
Let $0\leq h\leq h^{*}$ and $1\leq k\leq q_{n}(h)$, and let $(A,\mathcal{V})$ be a $(h,k)$-uniform selection pair. Then for any test function $f$ defined on the value space of $\big((u^{h}_{i})_{i=1,\ldots,k},(\nu^{i})_{1\leq i\leq h-1}\big)$, we have
$\mathbb{P}_{n}\Big[f\big(\mathcal{V},(\nu^{i})_{1\leq i\leq h-1}\big)\,\big|\,A\Big]=\mathbb{P}_{n}\Big[f\big((u^{h}_{i})_{1\leq i\leq k},(\nu^{i})_{1\leq i\leq h-1}\big)\Big].$
The proof follows the same reasoning as that in the proof of Theorem 3.14, and is therefore omitted.
Proof 4.16 (Proof of Proposition 4.11).
Recall the definition of $E(n,\delta,\eta,m)$ in (65). For each $0\leq r\leq K_{\theta}$, we split the event $E(n,\delta,\eta,m)$ according to the behavior of the process above and below $h_{r}$:
$E(n,\delta,\eta,m)=A_{r}\cap B_{r},$
where
$A_{r}:=\big\{X_{\bar{h}+\lfloor c_{0}n\rfloor}=X_{h_{r}}=m\big\},\qquad B_{r}:=\big\{X_{\bar{h}}=X_{h_{r}-1}=m\big\}.$
Then for any $1\leq r\leq K_{\theta}$ and subsets $D_{1},\ldots,D_{r}$ of $\mathbb{N}$, we have
$\mathbb{Q}^{m}_{n}\Bigg[\bigcap_{l=1}^{r}\{\mathcal{N}_{i,l}\in D_{l}\}\Bigg]=\frac{\mathbb{P}_{n}\Big[A_{r}\cap B_{r}\cap\bigcap_{l=1}^{r}\{\mathcal{N}_{i,l}\in D_{l}\}\Big]}{\mathbb{P}_{n}(A_{r}\cap B_{r})}.$ (70)
Lemma 4.17.
For any $0\leq r\leq K_{\theta}$, both $\big(A_{r},\mathcal{X}_{h_{r}}\big)$ and $\big(A_{r}\cap\bigcap_{l=1}^{r}\{\mathcal{N}_{i,l}\in D_{l}\},\mathcal{X}_{h_{r}}\big)$ are $(h_{r},m)$-uniform selection pairs.
Proof 4.18.
Clearly, both $A_{r}$ and $\bigcap_{l=1}^{r}\{\mathcal{N}_{i,l}\in D_{l}\}$ are measurable with respect to $\mathcal{G}_{h_{r}}\vee\sigma(\Lambda_{h_{r}})$. Furthermore, as in the proof of Proposition 3.12, the uniformity condition follows directly from the exchangeability of $\big(\mathscr{T}^{n}(v):|v|=h_{r}\big)$. This concludes the lemma.
Returning to (70), the next step is to apply Lemma 4.15 to the numerator on the RHS. Recall the definitions of $\widetilde{B}_{r}$ and $\widetilde{\mathcal{N}}_{i,r}$ introduced before Proposition 4.11.
By Lemmas 4.15 and 4.17, for $1\leq r\leq K_{\theta}$,
$\mathbb{P}_{n}\Big[A_{r}\cap B_{r}\cap\bigcap_{l=1}^{r}\{\mathcal{N}_{i,l}\in D_{l}\}\Big]$ $=\mathbb{P}_{n}\Big[A_{r-1}\cap\bigcap_{l=1}^{r-1}\{\mathcal{N}_{i,l}\in D_{l}\}\Big]\,\mathbb{P}_{n}\big(\widetilde{B}_{r}\cap\{\widetilde{\mathcal{N}}_{i,r}\in D_{r}\}\big),$
$\mathbb{P}_{n}\Big[A_{r}\cap B_{r}\cap\bigcap_{l=1}^{r-1}\{\mathcal{N}_{i,l}\in D_{l}\}\Big]$ $=\mathbb{P}_{n}\Big[A_{r-1}\cap\bigcap_{l=1}^{r-1}\{\mathcal{N}_{i,l}\in D_{l}\}\Big]\,\mathbb{P}_{n}(\widetilde{B}_{r}).$
Consequently,
$\mathbb{P}_{n}\Big[A_{r}\cap B_{r}\cap\bigcap_{l=1}^{r}\{\mathcal{N}_{i,l}\in D_{l}\}\Big]=\mathbb{P}_{n}\Big[A_{r}\cap B_{r}\cap\bigcap_{l=1}^{r-1}\{\mathcal{N}_{i,l}\in D_{l}\}\Big]\,\mathbb{P}_{n}\big(\widetilde{\mathcal{N}}_{i,r}\in D_{r}\,\big|\,\widetilde{B}_{r}\big).$
Substituting into (70), we obtain
$\mathbb{Q}^{m}_{n}\Bigg[\bigcap_{l=1}^{r}\{\mathcal{N}_{i,l}\in D_{l}\}\Bigg]=\mathbb{Q}^{m}_{n}\Bigg[\bigcap_{l=1}^{r-1}\{\mathcal{N}_{i,l}\in D_{l}\}\Bigg]\,\mathbb{P}_{n}\big(\widetilde{\mathcal{N}}_{i,r}\in D_{r}\,\big|\,\widetilde{B}_{r}\big).$
Iterating this identity over $1\leq r\leq K_{\theta}$ gives
$\mathbb{Q}^{m}_{n}\Bigg[\bigcap_{l=1}^{K_{\theta}}\{\mathcal{N}_{i,l}\in D_{l}\}\Bigg]=\prod_{r=1}^{K_{\theta}}\mathbb{P}_{n}\big(\widetilde{\mathcal{N}}_{i,r}\in D_{r}\,\big|\,\widetilde{B}_{r}\big),$
which proves the desired results.
Next, we turn to the proof of Proposition 4.12. By (69), it suffices to show that
$\mathbb{P}_{n}\big(\widetilde{\mathcal{N}}_{i,r}\geq\gamma n^{2}\,\big|\,\widetilde{B}_{r}\big)\geq C_{0}.$
This will be established using the second moment method. To this end, we shall estimate, for each $v^{j}_{l}$, the probability
$\mathbb{P}_{n}\Big[v^{j}_{l}\in\text{Desc}\big(\text{Anc}(u^{h_{r-1}}_{i},h_{r})\big)\,\Big|\,\widetilde{B}_{r}\Big],$
and for distinct $v^{j}_{l},v^{j^{\prime}}_{l^{\prime}}$,
$\mathbb{P}_{n}\Big[v^{j}_{l},v^{j^{\prime}}_{l^{\prime}}\in\text{Desc}\big(\text{Anc}(u^{h_{r-1}}_{i},h_{r})\big)\,\Big|\,\widetilde{B}_{r}\Big].$
We begin with some notation. For any $1\leq i\leq m\leq M_{0}$, $1\leq r\leq K_{\theta}$, and vertex $v^{j}_{l}$ with $h_{r}<j\leq h_{r-1}$ and $1\leq l\leq q_{n}(j)-m$, define
$\tau_{i}(v^{j}_{l}):=\inf\Big\{1\leq k\leq j-h_{r}:\text{Anc}(v^{j}_{l},j-k)\in U_{i,r}\Big\},$
with the convention $\inf\emptyset=\infty$. In words, $\tau_{i}(v^{j}_{l})$ is the “first time” $k$ at which the ancestral lineage of $v^{j}_{l}$ merges into the trunk path $U_{i,r}$. Clearly,
$\big\{v^{j}_{l}\in\text{Desc}\big(\text{Anc}(u^{h_{r-1}}_{i},h_{r})\big)\big\}=\big\{\tau_{i}(v^{j}_{l})<\infty\big\}.$
Proposition 4.19.
Assume (H1) and (H2’). Then:
  • (1)
    Uniformly for all $1\leq i\leq m\leq M_{0}$, $1\leq r\leq K_{\theta}$, and vertices $v^{j}_{l}$ with $h_{r}<j\leq h_{r-1}$ and $1\leq l\leq q_{n}(j)-m$,
    ${\mathbb{P}}_{n}\big(\tau_{i}(v^{j}_{l})<\infty\,\big|\,\widetilde{B}_{r}\big)=\frac{1}{m}\left\{1-\prod_{h=0}^{j-h_{r}-1}\Bigg[1-\frac{m\sigma^{2}_{n}(j-h)}{q_{n}(j-h)}\big(1+o_{n}(1)\big)\Bigg]\right\}.$ (71)
  • (2)
    Uniformly for all $1\leq i\leq m\leq M_{0}$ and vertices $v^{j}_{l},\,v^{j^{\prime}}_{l^{\prime}}$ with $h_{r}<j\leq j^{\prime}\leq h_{r-1}$, $1\leq l\leq q_{n}(j)-m$, and $1\leq l^{\prime}\leq q_{n}(j^{\prime})-m$,
    ${\mathbb{P}}_{n}\big(\tau_{i}(v^{j}_{l})<\infty\,\big|\,\tau_{i}(v^{j^{\prime}}_{l^{\prime}})<\infty,\,\widetilde{B}_{r}\big)$ (72)
    $\leq$ $\frac{1}{m}\left\{1-\big(1+o_{n}(1)\big)\prod_{h=0}^{j-h_{r}-1}\Bigg[1-\frac{(m+1)\sigma^{2}_{n}(j-h)}{q_{n}(j-h)}\big(1+o_{n}(1)\big)\Bigg]\right\}.$
Proof 4.20 (Proof of Proposition 4.12 assuming Proposition 4.19).
Throughout the proof, we use $c_{1},c_{2},\dots$ to denote positive constants, independent of $\theta$ (though they may depend on other fixed parameters $\delta,s,M_{0},c_{0}$).
We estimate the first and second moments of $\widetilde{\mathcal{N}}_{i,r}$. For the first moment,
$\mathbb{E}_{n}\big(\widetilde{\mathcal{N}}_{i,r}\,\big|\,\widetilde{B}_{r}\big)$ $=\sum_{j=h_{r}+1}^{h_{r-1}}\sum_{l=1}^{q_{n}(j)-m}{\mathbb{P}}_{n}\big(\tau_{i}(v^{j}_{l})<\infty\,\big|\,\widetilde{B}_{r}\big).$ (73)
We notice that $j-h_{r}\leq\theta n$ for $h_{r}<j\leq h_{r-1}$. By (71), (H1) and (H2’), we deduce that for all $n$ sufficiently large,
${\mathbb{P}}_{n}\big(\tau_{i}(v^{j}_{l})<\infty\,\big|\,\widetilde{B}_{r}\big)$ $\geq\frac{1}{m}\left\{1-\bigg[1-\frac{c_{1}m}{n}\bigg]^{j-h_{r}}\right\}\geq\frac{c_{1}(j-h_{r})}{2n},$
provided that $\theta\leq\frac{1}{2c_{1}M_{0}}\log 2$ (in which case $\frac{c_{1}m}{n}\leq\frac{\log 2}{2\theta n}\leq 1-(\frac{1}{2})^{\frac{1}{\theta n}}$ for all $n$ sufficiently large); the last equality then follows since $(1-x)^{k}\leq 1-\frac{k}{2}x$ for all $0\leq x\leq 1-(\frac{1}{2})^{\frac{1}{\theta n}}$ and $1\leq k\leq\theta n$. Substituting this into (73) gives
$\mathbb{E}_{n}\big(\widetilde{\mathcal{N}}_{i,r}\,\big|\,\widetilde{B}_{r}\big)\geq\sum_{j=h_{r}+1}^{h_{r-1}}c_{2}n\cdot\frac{c_{1}(j-h_{r})}{2n}=\frac{c_{1}c_{2}}{4}\lfloor\theta n\rfloor(\lfloor\theta n\rfloor-1)\geq c_{3}\theta^{2}n^{2}.$ (74)
We next derive an upper bound for $\mathbb{E}_{n}\big((\widetilde{\mathcal{N}}_{i,r})^{2}\,\big|\,\widetilde{B}_{r}\big)$. Note that
$\begin{aligned} \mathbb{E}_{n}\big((\widetilde{\mathcal{N}}_{i,r})^{2}\,\big|\,\widetilde{B}_{r}\big)&=\mathbb{E}_{n}\big(\widetilde{\mathcal{N}}_{i,r}\,\big|\,\widetilde{B}_{r}\big)+\sum_{(j,l)\neq(j^{\prime},l^{\prime})}\mathbb{P}_{n}\big(\tau_{i}(y^{j}_{l}),\,\tau_{i}(y^{j^{\prime}}_{l^{\prime}})<\infty\,\big|\,\widetilde{B}_{r}\big)\\ &\leq\mathbb{E}_{n}\big(\widetilde{\mathcal{N}}_{i,r}\,\big|\,\widetilde{B}_{r}\big)+2\sum_{(j,l)\neq(j^{\prime},l^{\prime}):j\leq j^{\prime}}\mathbb{P}_{n}\big(\tau_{i}(y^{j}_{l}),\,\tau_{i}(y^{j^{\prime}}_{l^{\prime}})<\infty\,\big|\,\widetilde{B}_{r}\big),\end{aligned}$ (75)
where the sum on the first line runs over all $h_{r}<j,j^{\prime}\leq h_{r-1}$, $1\leq l\leq q_{n}(j)-m$, $1\leq l^{\prime}\leq q_{n}(j^{\prime})-m$ with $(j,l)\neq(j^{\prime},l^{\prime})$, while the sum on the last line further restricts to $j\leq j^{\prime}$.
By (79), (H1) and (H2’), for $(j,l)\neq(j^{\prime},l^{\prime})$ with $j\leq j^{\prime}$,
${\mathbb{P}}_{n}\big(\tau_{i}(v^{j}_{l})<\infty\,\big|\,\tau_{i}(v^{j^{\prime}}_{l^{\prime}})<\infty,\,\widetilde{B}_{r}\big)$ $\leq\frac{1}{m}\left\{1-\big(1+o_{n}(1)\big)\Bigg[1-\frac{c_{4}m}{n}\Bigg]^{\theta n}\right\}\leq c_{5}\theta,$
for all $1\leq m\leq M_{0}$ and sufficiently large $n$. Consequently,
$\sum_{(j,l)\neq(j^{\prime},l^{\prime}):j\leq j^{\prime}}\mathbb{P}_{n}\big(\tau_{i}(y^{j}_{l}),\,\tau_{i}(y^{j^{\prime}}_{l^{\prime}})<\infty\,\big|\,\widetilde{B}_{r}\big)$
$=\sum_{(j,l)\neq(j^{\prime},l^{\prime}):j\leq j^{\prime}}\mathbb{P}_{n}\big(\tau_{i}(y^{j}_{l})<\infty\,\big|\,\tau_{i}(y^{j^{\prime}}_{l^{\prime}})<\infty,\,\widetilde{B}_{r}\big)\,\mathbb{P}_{n}\big(\tau_{i}(y^{j^{\prime}}_{l^{\prime}})<\infty\,\big|\,\widetilde{B}_{r}\big)$
$\leq c_{5}\theta\cdot\sum_{(j,l)\neq(j^{\prime},l^{\prime}):j\leq j^{\prime}}\mathbb{P}_{n}\big(\tau_{i}(y^{j^{\prime}}_{l^{\prime}})<\infty\,\big|\,\widetilde{B}_{r}\big)$
$\leq c_{6}\theta^{2}n^{2}\cdot\mathbb{E}_{n}\big(\widetilde{\mathcal{N}}_{i,r}\,\big|\,\widetilde{B}_{r}\big).$
Substituting this estimate into (75) yields
[…]
$\,\ \leq c_{7}\Big[\mathbb{E}_{n}\big(\widetilde{\mathcal{N}}_{i,r}\,\big|\,\widetilde{B}_{r}\big)\Big]^{2},$
for all large $n$. Finally, taking $\gamma=\tfrac{1}{2}c_{3}\theta^{2}$, we obtain
$\mathbb{P}_{n}(\mathcal{N}_{i,r}\geq\gamma n^{2}\,\big|\,\widetilde{B}_{r})\;\geq\;\mathbb{P}_{n}\!\left(\mathcal{N}_{i,r}\geq\tfrac{1}{2}\mathbb{E}_{n}(\mathcal{N}_{i,r})\,\Big|\,\widetilde{B}_{r}\right)\;\geq\;\frac{\Big[\mathbb{E}_{n}\big(\widetilde{\mathcal{N}}_{i,r}\,\big|\,\widetilde{B}_{r}\big)\Big]^{2}}{4\,\mathbb{E}_{n}\big((\widetilde{\mathcal{N}}_{i,r})^{2}\,\big|\,\widetilde{B}_{r}\big)}\;\geq\;\frac{1}{4c_{7}}.$
This completes the proof.
Finally, we proceed to the proof of Proposition 4.19.
Proof 4.21 (Proof of Proposition 4.19 (1)).
For any vertex $v^{j}_{l}$, set
$\tau(v^{j}_{l}):=\inf\{\tau_{i}(v^{j}_{l}):1\leq i\leq m\}=\inf\Big\{1\leq k\leq j-h_{r}:\text{Anc}(v^{j}_{l},j-k)\in\bigcup_{i=1}^{m}U_{i,r}\Big\}.$
We first record the following estimate, whose proof will be given afterwards.
Lemma 4.22.
Uniformly for all $1\leq m\leq M_{0}$, $1\leq r\leq K_{\theta}$, and vertices $v^{j}_{l}$ satisfying $h_{r}<j\leq h_{r-1}$ and $1\leq l\leq q_{n}(j)-m$, we have
$\mathbb{P}_{n}\big(\tau(v^{j}_{l})=\infty\,\big|\,\widetilde{B}_{r}\big)=\prod_{h=0}^{j-h_{r}-1}\Bigg[1-\frac{m\sigma^{2}_{n}(j-h)}{q_{n}(j-h)}\big(1+o_{n}(1)\big)\Bigg].$ (76)
Assuming Lemma 4.22, the desired estimate (71) follows directly from the exchangeability of the vector $\big(\mathscr{T}^{n}(v):|v|=h_{r}\big)$.
Proof 4.23 (Proof of Lemma 4.22).
We first note that
$\{\tau(v^{j}_{l})=\infty\}=\{\tau(v^{j}_{l})>j-h_{r}\}.$
Thus it suffices to prove that, uniformly for the quantities appearing in the lemma and $0\leq h\leq j-h_{r}-1$,
$\mathbb{P}_{n}\big(\tau(v^{j}_{l})>h+1\,\big|\,\tau(v^{j}_{l})>h,\,\widetilde{B}_{r}\big)=1-\frac{m\sigma^{2}_{n}(j-h)}{q_{n}(j-h)}\big(1+o_{n}(1)\big).$ (77)
The argument parallels that of Proposition 4.11. For $h_{r}\leq j_{2}<j_{1}\leq h_{r-1}$, set
$\widetilde{B}_{r}^{[j_{1},j_{2}]}$ $:=\Big\{\#\big\{\text{Anc}(u^{h_{r-1}}_{i},j_{2}):1\leq i\leq m\big\}=\#\big\{\text{Anc}(u^{h_{r-1}}_{i},j_{1}):1\leq i\leq m\big\}=m\Big\}$
={# $=\Big\{\#\text{\,ancestors of }\{u^{h_{r-1}}_{i}:1\leq i\leq m\}\text{ remains $m$ between heights }[j_{2},j_{1}]\Big\}.$
Then we have
$\widetilde{B}_{r}=\widetilde{B}^{[j-h,h_{r-1}]}_{r}\cap\widetilde{B}^{[h_{r},j-h]}_{r}$.
By Proposition 3.12 and the uniform ordering of vertices $v^{j}_{l}$, the vector
$\Big(\text{Anc}\big(u^{h_{r-1}}_{1},j-h\big),\ldots,\text{Anc}\big(u^{h_{r-1}}_{m},j-h\big),\text{Anc}\big(v^{j}_{l},j-h\big)\Big),$
together with the event $\{\tau(v^{j}_{l})>h\}\cap\widetilde{B}^{[j-h,h_{r-1}]}_{r}$, forms a (jh,m+1)(j-h,m+1)-uniform selection pair, with $h_{r-1}$ serving as the base height $h^{*}$ in Definition 4.14. Hence, by Lemma 4.17,
$\begin{aligned} \mathbb{P}_{n}\big(\tau(v^{j}_{l})>h+1,\,\widetilde{B}_{r}\big)&=\mathbb{P}_{n}\big(\tau(v^{j}_{l})>h,\,\widetilde{B}^{[j-h,h_{r-1}]}_{r}\big)\\ &\quad\cdot\mathbb{P}_{n}\Big[\mathfrak{p}(u^{j-h}_{1}),\ldots,\mathfrak{p}(u^{j-h}_{m+1})\text{ are distinct,}\\ &\quad\quad\quad\quad\#\big\{\text{Anc}(u^{j-h}_{i},h_{r}):1\leq i\leq m\big\}=m\Big].\end{aligned}$ (78)
For brevity, denote
$p_{n}:=\mathbb{P}_{n}\Big[\mathfrak{p}(u^{j-h}_{1}),\ldots,\mathfrak{p}(u^{j-h}_{m+1})\text{ are distinct, }\#\big\{\text{Anc}(u^{j-h}_{j},h_{r}):1\leq i\leq m\big\}=m\Big].$
Note that
$\Big(\big\{\mathfrak{p}(u^{j-h}_{j}):1\leq i\leq m\big\},\ \{\mathfrak{p}(u^{j-h}_{1}),\ldots,\mathfrak{p}(u^{j-h}_{m+1})\text{ are distinct}\}\Big)$
is a $(1,m)$-uniform selection pair with jhj-h as the base height. Applying Lemma 4.17 again yields
$p_{n}=\,$ $\mathbb{P}_{n}\big(\mathfrak{p}(u^{j-h}_{1}),\ldots,\mathfrak{p}(u^{j-h}_{m+1})\text{ are distinct}\big)\cdot\mathbb{P}_{n}\big(\#\big\{\text{Anc}(u^{j-h-1}_{j},h_{r}):1\leq i\leq m\big\}=m\big).$
By (24) in Corollary 3.9,
$\quad\,\mathbb{P}_{n}\Big(\mathfrak{p}(u^{j-h}_{1}),\ldots,\mathfrak{p}(u^{j-h}_{m+1})\text{ are distinct}\,\Big|\,\mathfrak{p}(u^{j-h}_{1}),\ldots,\mathfrak{p}(u^{j-h}_{m})\text{ are distinct}\Big)$
$=1-\dfrac{m\sigma_{n}^{2}(j-h)}{q_{n}(j-h)}\big(1+o_{n}(1)\big).$
Substituting into (78), we obtain
[…]
$\quad\cdot\mathbb{P}_{n}\big(\mathfrak{p}(u^{j-h}_{1}),\ldots,\mathfrak{p}(u^{j-h}_{m+1})\text{ are distinct}\big)$
$\quad\cdot\mathbb{P}_{n}\big(\#\big\{\text{Anc}(u^{j-h-1}_{j},h_{r}):1\leq i\leq m\big\}=m\big).$
Applying Lemma 4.17 once more, we deduce
[…]
$\quad\cdot\mathbb{P}_{n}\big(\mathfrak{p}(u^{j-h}_{1}),\ldots,\mathfrak{p}(u^{j-h}_{m+1})\text{ are distinct}\big)$
$\quad\cdot\mathbb{P}_{n}\big(\#\big\{\text{Anc}(u^{j-h-1}_{j},h_{r}):1\leq i\leq m\big\}=m\big).$
Therefore,
$\mathbb{P}_{n}\big(\tau(v^{j}_{l})>h+1,\,\widetilde{B}_{r}\big)=\left[1-\dfrac{m\sigma_{n}^{2}(j-h)}{q_{n}(j-h)}\big(1+o_{n}(1)\big)\right]\mathbb{P}_{n}\big(\tau(v^{j}_{l})>h,\,\widetilde{B}_{r}\big).$
This proves the lemma.
Proof 4.24 (Proof of Proposition 4.19 (2)).
The argument for part (2) is essentially the same as that for part (1), so we only give a sketch of the proof here.
Fix vertices $v^{j}_{l},\,v^{j^{\prime}}_{l^{\prime}}$ such that $h_{r}<j\leq j^{\prime}\leq h_{r-1}$, $1\leq l\leq q_{n}(j)-m$, and $1\leq l^{\prime}\leq q_{n}(j^{\prime})-m$. It suffices to show that, uniformly for any $0\leq k<j-h_{r}$ and $1\leq k^{\prime}\leq j^{\prime}-h_{r}$,
${\mathbb{P}}_{n}\big(\tau(v^{j}_{l})=\infty~\big|~\tau(v^{j^{\prime}}_{l^{\prime}})=k^{\prime},\,v^{j^{\prime}}_{l^{\prime}}\notin\text{Desc}(v^{j}_{l}),\,\widetilde{B}_{r}\big)$ (79)
$=$ $(1+o_{n}(1))\prod_{h=1}^{j-(j^{\prime}-k^{\prime})-2}\Bigg[1-\frac{(m+1)\sigma^{2}}{q_{n}(\bar{h}+j-h)}(1+o_{n}(1))\Bigg]$
$\quad\times\prod_{h=j-(j^{\prime}-k^{\prime})}^{j-k}\Bigg[1-\frac{m\sigma^{2}}{q_{n}(\bar{h}+j-h)}(1+o_{n}(1))\Bigg].$
Define
$p_{n}(h):={\mathbb{P}}_{n}\big(\tau(v^{j}_{l})>h+1~\big|~\tau(v^{j}_{l})>h,\,\tau(v^{j^{\prime}}_{l^{\prime}})=k^{\prime},\,v^{j^{\prime}}_{l^{\prime}}\notin\text{Desc}(v^{j}_{l}),\,\widetilde{B}_{r}\big).$
Under this conditioning, $j^{\prime}-k^{\prime}$ represents the height at which $v^{j^{\prime}}_{l^{\prime}}$ merges into $U_{i,r}$. By exchangeability, taking $h_{r-1}$ as the base height, we have:
  1. (i)
    If $j-h>j^{\prime}-k^{\prime}$, then the vector
    $\Big(\text{Anc}(u^{h_{r-1}}_{1},j-h),\ldots,\text{Anc}(u^{h_{r-1}}_{m},j-h),\text{Anc}(v^{j}_{l},j-h),\text{Anc}(v^{j^{\prime}}_{l^{\prime}},j-h)\Big),$
    together with the event
    $\Big\{\tau(v^{j}_{l})>h,\,\tau(v^{j^{\prime}}_{l^{\prime}})=k^{\prime},\,v^{j^{\prime}}_{l^{\prime}}\notin\text{Desc}(v^{j}_{l}),\,\widetilde{B}_{r}^{[j-h,h_{r-1}]}\Big\},$
    forms an (jh,m+2)(j-h,m+2)-uniform selection pair.
  2. (ii)
    If $j-h\leq j^{\prime}-k^{\prime}$, then the vector
    $\Big(\text{Anc}(u^{h_{r-1}}_{1},j-h),\ldots,\text{Anc}(u^{h_{r-1}}_{m},j-h),\text{Anc}(v^{j}_{l},j-h)\Big),$
    together with the event
    $\Big\{\tau(v^{j}_{l})>h,\,\tau(v^{j^{\prime}}_{l^{\prime}})=k^{\prime},\,v^{j^{\prime}}_{l^{\prime}}\notin\text{Desc}(v^{j}_{l}),\,\widetilde{B}_{r}^{[j-h,h_{r-1}]}\Big\},$
    forms an (jh,m+1)(j-h,m+1)-uniform selection pair.
Following the same reasoning as in the proof of Proposition 4.19 (1), we obtain:
  1. (i)
    Uniformly for $0\leq h\leq j-h_{r}-1$ with $j-h>j^{\prime}-k^{\prime}+1$,
    $q_{n}(h)$ $=\mathbb{P}_{n}\Big(\mathfrak{p}(u^{j-h}_{1}),\ldots,\mathfrak{p}(u^{j-h}_{m+2})\text{ are distinct}~\Big|~\mathfrak{p}(u^{j-h}_{1}),\ldots,\mathfrak{p}(u^{j-h}_{m+1})\text{ are distinct}\Big)$
[…]
  2. (ii)
    Uniformly for $0\leq h\leq j-h_{r}-1$ with $j-h\leq j^{\prime}-k^{\prime}$,
    $q_{n}(h)$ $=\mathbb{P}_{n}\Big(\mathfrak{p}(u^{j-h}_{1}),\ldots,\mathfrak{p}(u^{j-h}_{m+1})\text{ are distinct}~\Big|~\mathfrak{p}(u^{j-h}_{1}),\ldots,\mathfrak{p}(u^{j-h}_{m})\text{ are distinct}\Big)$
[…]
  3. (iii)
    For $h$ such that $j-h=j^{\prime}-k^{\prime}+1$,
    $q_{n}(h)$ $=\mathbb{P}_{n}\Big(\{\mathfrak{p}(u_{1}^{j-h})=\mathfrak{p}(u_{2}^{j-h})\}\cap\{\mathfrak{p}(u_{2}^{j-h}),\ldots,\mathfrak{p}(u_{m+1}^{j-h})\text{ are distinct}\}~\Big|~\text{same event}\Big)$
[…]
Combining these estimates, we conclude that
[…]
This completes the proof.
5 Proof of Theorem 1.1
In the previous section we established the tightness of $\widetilde{\mathscr{T}}^{n}$. Here we first prove, under (H1), (H2’) and (H2”), the finite-dimensional convergence of $\widetilde{\mathscr{T}}^{n}$ (Section 5.1); combined with tightness, this yields convergence of the contour functions. We then show in Section 5.2 that the discrepancy between the height and contour functions vanishes in the limit, so both processes converge to the same limiting process. Finally, we verify that (H2) implies (H2’) and (H2”) (Proposition 5.8), which completes the proof of Theorem 1.1.
5.1 Finite-dimensional convergence of $\widetilde{\mathscr{T}}^{n}$
Recall that $\mathcal{T}^{\ell,\sigma}$ is the rooted ordered real tree encoded by $W^{\ell^{\sigma}}_{\alpha^{-1}(I(\ell)t)}$, and that $\mathcal{T}^{\ell,\sigma}_{k}$ denotes its uniform $k$-point subtree, whose law is described in Theorem 2.20. Also recall that $\widetilde{\mathscr{T}}^{n}$ is the geometric tree obtained from $\mathscr{T}^{n}$ by assigning length $1/n$ to each edge. Let $(V_{n,1},\dots,V_{n,\#V(\mathscr{T}^{n})})$ be a uniform random ordering of $V(\mathscr{T}^{n})$, and let $\widetilde{\mathscr{T}}^{n}_{k}$ be the rooted ordered subtree of $\widetilde{\mathscr{T}}^{n}$ spanned by the root and $\{V_{n,1},\dots,V_{n,k}\}$. In this subsection we prove the following proposition on the finite-dimensional convergence of $\widetilde{\mathscr{T}}^{n}$.
Proposition 5.1.
Assume (H1) and (H2”). For every $k\geq 1$, $\widetilde{\mathscr{T}}^{n}_{k}\overset{d}{\longrightarrow}\mathcal{T}_{k}^{\ell,\sigma}$ in the sense of (6).
Proof 5.2.
The convergence in the sense of (6) is defined by the metric $d$ in (4). This metric is a function of the geometric structure of the tree, which is uniquely determined by the joint distribution of:
  • (1)
    the heights of the $k$ (ordered) leaves;
  • (2)
    the heights of all their associated branch points;
  • (3)
    the branching order of the tree.
Therefore, our goal is to show that the joint distribution of these components in $\widetilde{\mathscr{T}}^{n}_{k}$ converges to the corresponding joint distribution in $\mathcal{T}_{k}^{l}$. We break the proof into three parts.
(1) Convergence of leaf heights (”birth times”)
For any $\delta>0$, suppose $a_{n}\rightarrow a$ and $na_{n}$ are integers in $\delta n,h(q_{n})-\delta n]$. Then $n$ times the probability that the height of a uniform chosen vertex in $\mathscr{T}^{n}$ is $na_{n}$ is
$n\bigg(\sum_{s}q_{n}(s)\bigg)^{-1}q_{n}(na_{n})\longrightarrow\ell(a)/\int_{0}^{\infty}\ell(x)\,\mathrm{d}x.$ (80)
This resulting expression is precisely the probability density for the ”birth times” (i.e. heights) of the $k$ leaves in the limiting tree $\mathcal{T}_{k}^{l}$, as described in Theorem 2.20. By (H1), $n$ times the probability that the height of a uniform chosen vertex in $\mathscr{T}^{n}$ is $na_{n}$ in $0,\delta n]\cup[h(q_{n})-\delta n,h(q_{n})]$ goes to 0 as $\delta\rightarrow 0.$ Thus, the distribution of the $k$ leaf heights in $\widetilde{\mathscr{T}}_{k}^{n}$ converges to the distribution of leaf heights in $\mathcal{T}_{k}^{l}$.
(2) Convergence of branch point heights (”the coalescent process”)
Let $e_{n}(s)$ and $e(s)$ be the number of different branches at height $s$ of $\widetilde{\mathscr{T}}^{n}_{k}$ and $\mathcal{T}_{k}^{\ell}$ respectively. Denote by $a_{n,k}<\cdots<a_{n,2}<a_{n,1}$ the order statistics of the $k$ leaf heights of $\widetilde{\mathscr{T}}^{n}_{k}$. We first condition on the event that the $k$ leaf heights are in $\delta n,h(q_{n})-\delta n]$ for some $\delta$. Let us view $\widetilde{\mathscr{T}}^{n}_{k}$ as a coalescent process. First, we trace backward in time the lineage of the leaf at height $a_{n,1}$ until arriving at the height $a_{n,2}$. So $e_{n}(s)=1$ for $a_{n,2}<s\leq a_{n,,1}$. The ancestral at height $a_{n,2}$ of the first leaf is uniformly chosen among all the vertices at that height. So when $a_{n,3}<s\leq a_{n,2}$, $e_{n}(s)$ will behave like $X^{(2)}_{\lfloor ns\rfloor}(n,\lfloor na_{n,2}\rfloor)$. The coalescent process follows similarly below height $a_{n,3}$. By (1) and Proposition A.1, we can deduce that for every $\delta$ and $\delta<s<h(\ell)-\delta$,
$e_{n}(s)\rightarrow e(s).$ (81)
For the branch points near the boundary $h(q_{n})$, notice that the probability that branch points of $\mathscr{T}_{k}^{n}$ fall into $h(q_{n})-\delta,h(q_{n})]$ tends to 0 as $\delta\rightarrow 0$. For the branch points near the boundary 0, the difference of branch point heights of $\widetilde{\mathscr{T}}^{n}_{k}$ and $\mathcal{T}_{k}^{l}$ will tend to 0 as $\delta\rightarrow 0$. Thus, the joint distribution of the branch point heights in $\widetilde{\mathscr{T}}_{k}^{n}$ converges to that in $\mathcal{T}_{k}^{l}$.
(3) Convergence of branching order
$\mathscr{T}^{n}$ is defined by an exchangeable offspring distribution. This implies that the subtrees branching from any given parent are, in effect, placed in a uniformly random (lexicographic) order. This property directly corresponds to the construction of the limiting coalescent $\mathcal{T}_{k}^{l}$, where merging branches are randomly assigned a ”left” or ”right” position. Therefore, the order structure also converges.
Remark 5.3.
As is clear from the proof of Proposition 5.1, finite-dimensional convergence of the tree sequence is equivalent to convergence to Kingman’s coalescent in the sense of Proposition A.1.
Corollary 5.4.
Assume (H1), (H2’) and (H2”). Then
$\big(n^{-1}C_{\lfloor 2n^{2}t\rfloor}(\mathscr{T}^{n})\big)_{t\geq 0}\ \overset{d}{\longrightarrow}\ \big(W_{\alpha^{-1}(t)}^{\,\frac{4\ell}{\sigma^{2}}}\big)_{t\geq 0}\quad\text{in $D[0,\infty)$}.$ (82)
Proof $5.$
Theorems 4.1 and 5.1 establish tightness and finite-dimensional convergence of the sequence $(\mathscr{T}^{n})$ respectively. Recall that $\mathcal{T}^{\ell,\sigma}_{k}$ is the uniform $k$-point subtree induced by $W^{\ell^{\sigma}}_{\alpha^{-1}(I(\ell)t)}$, and note that
$\big(n^{-1}C_{k}(\mathscr{T}^{n})\big)_{0\leq k\leq 2(\#V(\mathscr{T}^{n})-1)}$
is exactly the contour function of $\widetilde{\mathscr{T}}^{n}$. Applying Theorem 2.6 yields
$\big(n^{-1}C_{\lfloor 2(\#V(\mathscr{T}^{n})-1)t\rfloor}(\mathscr{T}^{n})\big)_{0\leq t\leq 1}\ \overset{d}{\longrightarrow}\ \big(W_{\alpha^{-1}(I(\ell)t)}^{\,\frac{4\ell}{\sigma^{2}}}\big)_{0\leq t\leq 1}.$
Combining this with (2) gives (82).
5.2 Joint convergence of height and contour functions
In this subsection we prove the joint convergence of the (rescaled) contour and height functions under (H1), (H2’) and (H2”). Finally, we show that (H2) implies (H2’) and (H2”), so the same conclusion holds under (H1)–(H2), completing the proof of Theorem 1.1.
Theorem 5.6.
Assume (H1), (H2’) and (H2”). Then
$\Big(n^{-1}C_{\lfloor 2n^{2}t_{1}\rfloor}(\mathscr{T}^{n}),n^{-1}H_{\lfloor n^{2}t_{2}\rfloor}(\mathscr{T}^{n})\Big)_{0\leq t_{1},t_{2}<\infty}\overset{d}{\longrightarrow}\Big(W_{\alpha^{-1}(t_{1})}^{\ell^{\sigma}},W^{\ell^{\sigma}}_{\alpha^{-1}(t_{2})}\Big)_{0\leq t_{1},t_{2}<\infty}$ (83)
in the product space $\big(D[0,\infty)\big)^{2}$.
Proof 5.7.
Let $\rho=v^{n}_{0}<v^{n}_{1}<\cdots<v^{n}_{\#V(\mathscr{T}^{n})}$ be the enumeration of all vertices of $\mathscr{T}^{n}$ in lexicographic order. This is also the order in which the vertices are visited during the depth-first traversal of $\mathscr{T}^{n}$. For $i=0,\ldots,\#V(\mathscr{T}^{n})$, let $\tau^{n}_{i}$ denote the first hitting time of $v^{n}_{i}$ in this traversal.
Recall the definition of the contour function from Section 2.2.1. It is immediate that
$H_{i}(\mathscr{T}^{n})=C_{\tau^{n}_{i}}(\mathscr{T}^{n}).$
Moreover, by [marckert2003depth, Lemma 2], we have
$\tau^{n}_{i}+H_{i}(\mathscr{T}^{n})=2i.$
Hence, for $0\leq t\leq 1$,
$\left|\frac{C_{\lfloor 2n^{2}t\rfloor}(\mathscr{T}^{n})}{n}-\frac{H_{\lfloor n^{2}t\rfloor}(\mathscr{T}^{n})}{n}\right|$ $=\left|\frac{C_{\lfloor 2n^{2}t\rfloor}(\mathscr{T}^{n})}{n}-\frac{C_{\tau^{n}_{\lfloor n^{2}t\rfloor}}(\mathscr{T}^{n})}{n}\right|$
$=\left|\frac{C_{\lfloor 2n^{2}t\rfloor}(\mathscr{T}^{n})}{n}-\frac{C_{2\lfloor n^{2}t\rfloor-H_{\lfloor n^{2}t\rfloor}(\mathscr{T}^{n})}(\mathscr{T}^{n})}{n}\right|$
$\overset{(*)}{\leq}\sup_{\big(t-\frac{h(q_{n})+1}{2n^{2}}\big)\vee 0\leq s\leq t}\left|\frac{C_{\lfloor 2n^{2}t\rfloor}(\mathscr{T}^{n})}{n}-\frac{C_{\lfloor 2n^{2}s\rfloor}(\mathscr{T}^{n})}{n}\right|,$
where ()(*) follows from the inequality
$\lfloor 2n^{2}t\rfloor-\big(2\lfloor n^{2}t\rfloor-H_{\lfloor n^{2}t\rfloor}(\mathscr{T}^{n})\big)\leq h(q_{n})+1.$
Then, by the tightness of the rescaled contour functions
$\Big\{n^{-1}C_{\lfloor 2n^{2}\cdot\rfloor}(\mathscr{T}^{n}):n\geq 1\Big\},$
together with the fact that $h(q_{n})/n\to h(q)$ as $n\to\infty$, we obtain
$\sup_{t\geq 0}\Big|n^{-1}C_{\lfloor 2n^{2}t\rfloor}(\mathscr{T}^{n})-n^{-1}H_{\lfloor n^{2}t\rfloor}(\mathscr{T}^{n})\Big|\ \overset{d}{\longrightarrow}\ 0.$
Combining this with the convergence of the contour functions, we conclude the joint convergence.
Proposition 5.8.
The condition (H2) implies both conditions (H2’) and (H2”).
Proof 5.9.
[…]
$A_{i_{1}i_{2}}=A^{s}_{i_{1}i_{2}}:=\{\mathfrak{p}(u^{s}_{i_{1}})=\mathfrak{p}(u^{s}_{i_{2}})\}.$
Then applying the inclusion-exclusion principle, we get
${\mathbb{P}}_{n}\big(\#\{\mathfrak{p}(u^{s}_{i}):1\leq i\leq m\}\leq m-1\big)={\mathbb{P}}_{n}\Big(\bigcup_{1\leq i_{1}<i_{2}\leq m}A_{i_{1}i_{2}}\Big)$ (84)
$\geq$ $\binom{m}{2}{\mathbb{P}}_{n}(A_{12})-3\binom{m}{3}{\mathbb{P}}_{n}(A_{12}\cap A_{13})-6\binom{m}{4}{\mathbb{P}}_{n}(A_{12}\cap A_{34}).$
By Corollary 3.3 and (H2)H2(i), there exist constants $c_{i}=c_{i}(\delta)>0$ ($i=1,2,3$) such that for all $n$ sufficiently large and $\delta n\leq s\leq h(q_{n})-\delta n$,
${\mathbb{P}}_{n}(A_{12})$ $=\frac{q_{n}(s-1)}{(q_{n}(s))_{2}}\,{\mathbb{E}}_{n}\!\big[(\nu^{s-1}_{1})_{2}\big]\;\geq\;c_{1}\dfrac{\sigma^{2}}{n},$
${\mathbb{P}}_{n}(A_{12}\cap A_{13})$ $=\frac{q_{n}(s-1)}{(q_{n}(s))_{3}}\,{\mathbb{E}}_{n}\!\big[(\nu^{s-1}_{1})_{3}\big]\;\leq\;c_{2}\dfrac{1}{n(\log n)^{2+\varepsilon}},$
${\mathbb{P}}_{n}(A_{12}\cap A_{34})$ $\leq\frac{q_{n}(s-1)}{(q_{n}(s))_{4}}\,{\mathbb{E}}_{n}\!\big[(\nu^{s-1}_{1})_{4}\big]+\frac{(q_{n}(s-1))_{2}}{(q_{n}(s))_{4}}\,{\mathbb{E}}_{n}\!\big[(\nu^{s-1}_{1})_{2}(\nu^{s-1}_{2})_{2}\big]\overset{(*)}{\leq}c_{3}\dfrac{1}{n(\log n)^{2+\varepsilon}},$
where ()(*) uses (15) and (H2)H2(i). Substituting these into (84) with $m=(\log n)^{1+\varepsilon/4}$ yields (35).
We now prove (H2)H2(ii) $\Rightarrow$ (H2’). By (H2)H2(ii), for each $\delta>0$ there exist $N=N(\delta)$ and $K=K(\delta)\in\mathbb{N}_{+}$ such that
$\tilde{\sigma}^{2}:=\inf_{n\geq N,\,\delta n\leq s\leq h(q_{n})-\delta n}{\mathbb{E}}_{n}\big[\nu^{s}_{1}(\nu^{s}_{1}-1)\mathbf{1}\{\nu^{s}_{1}\leq K\}\big]>0.$ (85)
Define truncated coalescent events
$B_{i_{1}i_{2}}:=\{\mathfrak{p}(u^{s}_{i_{1}})=\mathfrak{p}(u^{s}_{i_{2}}),\,\xi(\mathfrak{p}(u^{s}_{i_{1}}))\leq K\},\qquad 1\leq i_{1}<i_{2}\leq m.$
Then
$\Big\{\#\{\mathfrak{p}(u^{s}_{i}):1\leq i\leq m\}\leq m-1\Big\}=\bigcup_{1\leq i_{1}<i_{2}\leq m}\big\{\mathfrak{p}(u^{s}_{i_{1}})=\mathfrak{p}(u^{s}_{i_{2}})\big\}\supset\bigcup_{1\leq i_{1}<i_{2}\leq m}B_{i_{1}i_{2}}.$ (86)
Applying the inclusion–exclusion principle, we get
${\mathbb{P}}_{n}\Big(\bigcup_{1\leq i_{1}<i_{2}\leq m}B_{i_{1}i_{2}}\Big)$ $\geq\binom{m}{2}{\mathbb{P}}_{n}(B_{12})-3\binom{m}{3}{\mathbb{P}}_{n}(B_{12}\cap B_{13})-6\binom{m}{4}{\mathbb{P}}_{n}(B_{12}\cap B_{34}).$
By Corollary 3.4, for all $n\geq N$,
${\mathbb{P}}_{n}(B_{12})$ $=\frac{q_{n}(s-1)}{(q_{n}(s))_{2}}\,{\mathbb{E}}_{n}\!\big[(\nu^{s-1}_{1})_{2}\mathbf{1}\{\nu^{s-1}_{1}\leq K\}\big]\;\geq\;\frac{q_{n}(s-1)}{(q_{n}(s))_{2}}\,\tilde{\sigma}^{2},$
${\mathbb{P}}_{n}(B_{12}\cap B_{13})$ $=\frac{q_{n}(s-1)}{(q_{n}(s))_{3}}\,{\mathbb{E}}_{n}\!\big[(\nu^{s-1}_{1})_{3}\mathbf{1}\{\nu^{s-1}_{1}\leq K\}\big]\;\leq\;\frac{q_{n}(s-1)}{(q_{n}(s))_{3}}\,K^{3},$
${\mathbb{P}}_{n}(B_{12}\cap B_{34})$ $\leq\frac{q_{n}(s-1)}{(q_{n}(s))_{4}}\,{\mathbb{E}}_{n}\!\big[(\nu^{s-1}_{1})_{4}\mathbf{1}\{\nu^{s-1}_{1}\leq K\}\big]$
$\quad+\frac{(q_{n}(s-1))_{2}}{(q_{n}(s))_{4}}\,{\mathbb{E}}_{n}\!\big[(\nu^{s-1}_{1})_{2}(\nu^{s-1}_{2})_{2}\mathbf{1}\{\nu^{s-1}_{1},\nu^{s-1}_{2}\leq K\}\big]$
$\leq 2\,\frac{(q_{n}(s-1))_{2}}{(q_{n}(s))_{4}}\,K^{4}.$
Hence, by (H1), there exist constants $c_{i}=c_{i}(\delta,\tilde{\sigma},K,N)>0$ $(i=1,2,3)$ such that, for all $n\geq N$, $\delta n\leq s\leq h(q_{n})-\delta n$ and $1\leq m\leq q_{n}(s)$,
${\mathbb{P}}_{n}\Big(\bigcup_{1\leq i<j\leq m}B_{ij}\Big)\;\geq\;c_{1}\frac{m^{2}}{n}-c_{2}\frac{m^{3}}{n^{2}}-c_{3}\frac{m^{4}}{n^{2}}.$
In particular, choosing $m=\lfloor(\log n)^{1+\varepsilon}\rfloor$ and using (86) yields (35).
Proof 5.10 (Proof of Theorem 1.1).
Combining Theorem 5.6 with Proposition 5.8 yields Theorem 1.1. This completes the proof.
Appendix A Convergence to Kingman’s coalescent
Fix an arbitrary height $\lfloor nh^{*}\rfloor\in[\delta n,h(q_{n})-\delta n]$. Consider the coalescent process as defined in Definition 25 with $h^{*}$ replaced by $\lfloor nh^{*}\rfloor$ there:
$X^{(k)}_{j}(n):=X^{(k)}_{j}(n,\lfloor nh^{*}\rfloor),\qquad 0\leq j\leq\lfloor nh^{*}\rfloor.$
Consider another continuous coalescent that starts with $k$ particles and each pair of clusters merges into one cluster with rate $\sigma(h^{*}-t)^{2}/\ell(h^{*}-t)$. Let $Y^{(k)}=(Y^{(k)}_{t}:0\leq t\leq h^{*})$ be the number of different clusters of the coalescent during time $0,h^{*}]$. We will show the following convergence result.
Proposition A.1.
Assume (H1) and (H2”). Then for any $k\geq 1$,
$\big(X^{(k)}_{\lfloor n(h^{*}-t)\rfloor}(n)\big)_{0\leq t\leq h^{*}-\delta}\overset{d}{\longrightarrow}(Y^{(k)}_{t})_{0\leq t\leq h^{*}-\delta}$
in the Skorokhod space $D[0,\infty)$.
Proof A.2.
Consider an arbitrary path $A(h^{*})$ with jump times $0=t_{0}<t_{1}<\cdots<t_{m}<t_{m+1}=h^{*}-\delta$ ($m<k$) and jump size 1-1 at each jump time. Denote by $I_{n}$ (resp. II) the probability that $X^{(k)}_{\lfloor n(h^{*}-\cdot)\rfloor}(n)$ (resp. $Y^{(k)}$) follows the path $A(h^{*})$. Then
$n^{m}I_{n}$ $=\prod_{i=0}^{m}\prod_{z\in(\lfloor n(h^{*}-t_{i+1})\rfloor+1,\lfloor n(h^{*}-t_{i})\rfloor]}{\mathbb{P}}_{n}\Big(\mathfrak{p}(u_{1}^{z}),\ldots,\mathfrak{p}(u_{k-i}^{z})~\text{are distinct}\Big)$
$\quad\cdot n^{m}\prod_{z\in\{\lfloor n(h^{*}-t_{i})\rfloor+1:i=1,\cdots,m\}}{\mathbb{P}}_{n}\Big(\mathfrak{p}(u_{1}^{z})=\mathfrak{p}(u_{2}^{z}),\mathfrak{p}(u_{2}^{z}),\ldots,\mathfrak{p}(u_{k-i+1}^{z})~\text{are distinct}\Big)$
$\longrightarrow\exp\left[-\sum_{i=0}^{m}\int_{h^{*}-t_{i+1}}^{h^{*}-t_{i}}\binom{k-i}{2}\frac{\sigma(s)^{2}}{\ell(s)}\,\mathrm{d}s\right]\prod_{i=1}^{m}\dfrac{\sigma(h^{*}-t_{i})^{2}}{\ell(h^{*}-t_{i})}=I,$
where the last approximation follows from Corollary 3.9 and (H1). We complete the proof by [billingsley2013convergence, Theorem 3.3].
Appendix B Near-sharpness of (H2’): a counterexample
Fix $\alpha>0$. For all $n\in{\mathbb{N}}$, denote by $r_{n}=\lfloor\frac{n}{(\log n)^{\alpha}}\rfloor$ and $p_{n}:=\frac{(\log n)^{2\alpha}}{n}$. Let $\nu=(\nu_{1},\cdots,\nu_{n})$ be a random vector distributed as follows:
  • With probability $1-p_{n}$, $\nu_{i}=1$ for all $1\leq i\leq n$;
[…]
    ${\mathbb{P}}(\nu_{i_{0}}=r_{n};\,\nu_{i_{m}}=0,\,\forall\,1\leq m\leq r_{n}-1;\,\nu_{j}=1,\forall\,j\notin\{i_{m}:0\leq m\leq r_{n}-1\})$
    $=\,$ $p_{n}\bigg[r_{n}\binom{n}{r_{n}}\bigg]^{-1}.$
Clearly, $\sum_{i=1}^{n}\nu_{i}=n$, and the random vector $\nu$ is exchangeable. Moreover, as $n\to+\infty$,
${\mathbb{E}}[\nu_{1}]=1,\quad{\mathbb{E}}[(\nu_{1})_{2}]\to 1,\quad{\mathbb{E}}[(\nu)_{3}]\sim\frac{n}{(\log n)^{\alpha}}.$
Let $(\nu^{s}:1\leq s\leq n)$ be independent and identically distributed as $\nu$. We consider the classical Cannings tree $\mathscr{T}^{n}$ associated with offspring $(\nu^{s}:1\leq s\leq n)$ as defined in Definition 2.1, i.e. $h(q_{n})=n+1$, $q_{n}(s)=n$ and vertices at generation $1\leq s\leq n$ has offspring $\nu^{s}$. As before, we assign length $1/n$ to each edge of $\mathscr{T}^{n}$ and obtain the geometric tree $\widetilde{\mathscr{T}}^{n}$. Since ${\mathbb{E}}[(\nu_{1})^{3}]/n\to 0$, the genealogy of finite vertices, after scaling, will converge to Kingman’s coalescent. However, the contour function of $\widetilde{\mathscr{T}}^{n}$ may not converge because of the following result:
Proposition B.1.
Suppose $0<\alpha<1$. Then $(\widetilde{\mathscr{T}}^{n})$ is not tight; in particular, the convergence stated in Theorem 1.1 fails.
Given a sequence of classical Cannings tree $(T^{n})$, we still denote by $\widetilde{T}^{n}$ the geometric tree obtained from $T^{n}$ with edge length replaced by $1/n$. Recall the coalescent process $\mathcal{X}_{j}$ and $X_{j}:=\#\mathcal{X}_{j}$ as defined in (25) with $h^{*}=n+1$ there, so $X_{1}$ is the number of ancestors at height $1$ of all individuals at height $n+1$. The idea of the proof of Proposition B.1 follows by the following observation: the tightness of trees of $(\widetilde{T}^{n})$ implies the coming-down-from-infinity property of $(\widetilde{T}^{n})$ (hence $(T^{n})$) introduced in Section 3.3.
Lemma B.2.
Suppose $(\widetilde{T}^{n})$ is tight. Then
$\lim_{M\to\infty}\limsup_{n\to\infty}{\mathbb{P}}(X_{1}>M)=0$
Proof B.3.
Recall the definition of $\Delta(n,k)$ in (5). Then that $(\widetilde{T}^{n})$ is tight means for any $\delta>0$,
$\lim_{k\to\infty}\limsup_{n\to\infty}{\mathbb{P}}(\max_{v\in V(T^{n})}d_{T^{n}}(v,\{V_{n,1},\cdots,V_{n,k}\})>\delta n)=0,$
where $V_{n,1},\cdots,V_{n,k}$ are $k$ vertices uniformly selected from $V(T^{n})$. We observe that on the event $\{\max_{v\in V(T^{n})}d(v,\{V_{n,1},\cdots,V_{n,k}\})\leq\frac{1}{2}n\}$, any vertex at height $n$ must share a common ancestor at height $\lfloor\frac{1}{2}n\rfloor$ with some $V_{n,l},\,1\leq l\leq k$. Since the number of ancestors at the first generation of $\{V_{n,1},\cdots,V_{n,k}\}$ is no greater than $k$, we deduce that $X_{1}\leq k$ on this event. Therefore for all $\varepsilon>0$ and then all $k$ large enough,
$\limsup_{n\to\infty}{\mathbb{P}}(X_{1}>k)\leq\limsup_{n\to\infty}{\mathbb{P}}\Big(\max_{v\in V(T^{n})}d(v,\{V_{n,1},\cdots,V_{n,k}\})>\frac{1}{2}n\Big)<\varepsilon.$
We complete the proof.
Proof B.4 (Proof of Proposition B.1).
In view of Lemma B.2, it suffices to show that the coming-down-from-infinity property fails for $(\mathscr{T}^{n})$, i.e. there exists some $\varepsilon_{0}>0$, such that for all $k\geq 1$,
$\limsup_{n\to\infty}{\mathbb{P}}_{n}(X_{n}>k)\geq\varepsilon_{0}.$
Let $S:=\#\{1\leq s\leq n:\exists\,1\leq i\leq n,\ \nu^{s}_{i}\neq 1\}$, $h_{1}>\cdots>h_{S}$ be the generation where the offspring is non-trivial and $m_{i}$ be the number such that $\nu^{h_{i}}_{m_{i}}=r_{n}$ for $1\leq i\leq S$. By the definition of $\nu$ and the independence of $(\nu^{s})$ across $s$, we obtain that conditionally on $S$, $\{h_{i}:1\leq i\leq S\}$ is uniformly chosen in $\{1,2,\cdots,n\}$, and $(m_{i}:1\leq i\leq S)$ is a sequence of i.i.d. random variables distributed as the uniform distribution on $\{1,2,\cdots,n\}$. We then define $k_{0}=k$ and for all $1\leq i\leq S$,
$k_{i}:=\inf\{j<m_{i}:\#\{1\leq l\leq j:\nu^{h_{i}}_{l}=1\}=k_{i-1}\},$
with $\inf\emptyset=+\infty$ by convention. We observe that the sequence $(k_{i}:0\leq i\leq S)$ is increasing, and $\{k_{S}<\infty\}\subseteq\{X_{1}\geq k\}$. Indeed, recall the construction of $\mathscr{T}^{n}$ and that $m_{i}$ satisfies $\nu^{h_{i}}_{m_{i}}=r_{n}$. Then on the event $\{k_{S}<\infty\}$, the first $k$ vertices at height $n+1$ do not share any common ancestors at any height $h_{i}$ (hence also at any height $s\in[1,n]$), implying that $X_{1}\geq k$. Therefore, it suffices to prove that there exists some $\varepsilon_{0}>0$, such that for all $k\geq 1$,
$\limsup_{n\to\infty}{\mathbb{P}}_{n}(k_{S}<\infty)\geq\varepsilon_{0}.$ (87)
We divide the proof of (87) into several steps.
Step 1. Note that $S\sim\text{Binomial}~\left(n,\tfrac{(\log n)^{2\alpha}}{n}\right)$. Therefore there exists some $c_{1}>0$ such that for all $n$ sufficiently large,
${\mathbb{P}}_{n}(S\geq 3(\log n)^{2\alpha})\leq\exp(-c_{1}(\log n)^{2\alpha}),$ (88)
by the Chernoff bound. Let $m_{*}=\min\{m_{i}:1\leq i\leq S\}$. We prove that for all $\varepsilon>0$ small, there exists some $c_{2}>0$, such that for all $n$ large enough,
${\mathbb{P}}_{n}\bigg(m_{*}>\frac{c_{2}n}{(\log n)^{2\alpha}},\,S<3(\log n)^{2\alpha}\bigg)\geq 1-\varepsilon.$ (89)
Indeed, let $(U_{i}:i\geq 1)$ be a sequence of i.i.d random variables with the law of uniform distribution on $\{1,\cdots,n\}$. We derive that
${\mathbb{P}}\bigg(m_{*}\leq\frac{c_{2}n}{(\log n)^{2\alpha}},\,S<3(\log n)^{2\alpha}\bigg)\leq{\mathbb{P}}\bigg(\min\{U_{i}:1\leq i\leq\lfloor 3(\log n)^{2\alpha}\rfloor\}\leq\frac{c_{2}n}{(\log n)^{2\alpha}}\bigg)$
$=\,$ $1-\bigg(1-\frac{c_{2}}{(\log n)^{2\alpha}}\bigg)^{\lfloor 3(\log n)^{2\alpha}\rfloor}\leq 4c_{2}$
for all $n$ large enough. Combining with (88) and taking $c_{2}=\varepsilon/8$, we deduce that (89) holds for all $n$ sufficiently large.
Step 2. Define the conditional probability measure
$\widetilde{{\mathbb{P}}}_{n}(\cdot):={\mathbb{P}}_{n}\bigg(\cdot\bigg|\,m_{*}>\frac{c_{2}n}{(\log n)^{2\alpha}},\,S<3(\log n)^{2\alpha}\bigg).$
We now prove that there exists some $c_{3}>0$, such that for all $k,i\geq 1$ and $n$ large enough, on the event $\{S>i,\,k_{i}<\frac{c_{2}n}{2(\log n)^{2\alpha}}\}$,
$\widetilde{{\mathbb{P}}}_{n}\bigg(k_{i+1}\leq k_{i}\Big(1+\frac{2}{(\log n)^{\alpha}}\Big)\,\bigg|\,k_{i}\bigg)\geq 1-\exp\Big(-\frac{c_{3}k_{i}}{(\log n)^{2\alpha}}\Big).$ (90)
Conditionally on $h_{1},\cdots,h_{S}$ and the event $\{m_{*}>\frac{c_{2}n}{(\log n)^{2\alpha}},\,S<3(\log n)^{2\alpha}\}$, the value of $k_{i}$ only depends on $(\nu^{s}:s\geq h_{i})$ hence independent of $(\nu^{s}:s\leq h_{i}-1)$. Furthermore, under the same condition and on the event $\{S>i,\,k_{i}<\frac{c_{2}n}{2(\log n)^{2\alpha}}\}$, $k_{i}<m_{*}$ for all $n$ sufficiently large, and therefore for $k_{i}\leq l<m_{*}$, $k_{i+1}\leq l$ if and only if there are no less than $k_{i}$ of $(\nu^{h_{i+1}}_{j}:1\leq j\leq l)$ taking the value 1. Let $Y_{l}$ be the Hypergeometric random variable $H(n-1,n-r_{n},l)$. Then we derive that on the event $\{k_{i}<\frac{c_{2}n}{2(\log n)^{2\alpha}}\}$,
$\widetilde{{\mathbb{P}}}_{n}\bigg(k_{i+1}\leq k_{i}\Big(1+\frac{2}{(\log n)^{\alpha}}\Big)\,\bigg|\,k_{i},\,i<S\bigg)={\mathbb{P}}(Y_{\lfloor m(1+\frac{2}{(\log n)^{\alpha}})\rfloor}\geq m)\,\big|_{m=k_{i}}.$
Since ${\mathbb{E}}[Y_{l}]=l(n-r_{n})/(n-1)\approx l(1-\frac{1}{(\log n)^{\alpha}})$, we obtain by [Serfling, Corollary 1.1] that for all $n$ large enough,
${\mathbb{P}}(Y_{\lfloor m(1+\frac{2}{(\log n)^{\alpha}})\rfloor}<m)\leq\exp\bigg(-\frac{3(\frac{m}{(\log n)^{\alpha}})^{2}}{\lfloor m(1+\frac{2}{(\log n)^{\alpha}})\rfloor}\bigg)\leq\exp\Big(-\frac{c_{3}m}{(\log n)^{2\alpha}}\Big).$
We deduce the desired inequality by taking $c_{3}$ large enough.
Step 3. Let $i_{*}:=\inf\{1\leq i\leq S:k_{i}>\sqrt{n}\}$. We now prove that
$\widetilde{{\mathbb{P}}}_{n}\bigg(k_{i_{*}}\leq\sqrt{n}\Big(1+\frac{2}{(\log n)^{\alpha}}\Big)\,\bigg|\,i_{*}\leq S\bigg)\geq 1-\exp\Big(-\frac{c_{3}\sqrt{n}}{(\log n)^{2\alpha}}\Big).$ (91)
By conditioning on $S$ and the total probability formula, we have
$\widetilde{{\mathbb{P}}}_{n}\bigg(k_{i_{*}}\leq\sqrt{n}\Big(1+\frac{2}{(\log n)^{\alpha}}\Big)\,\bigg|\,i_{*}\leq S\bigg)$ (92)
$=$ $\frac{\sum_{i=1}^{\lfloor 3(\log n)^{2\alpha}\rfloor}\widetilde{{\mathbb{P}}}_{n}(\sqrt{n}<k_{i}\leq\sqrt{n}(1+\frac{2}{(\log n)^{\alpha}}),\,k_{i-1}\leq\sqrt{n},\,i\leq S)}{\sum_{i=1}^{\lfloor 3(\log n)^{2\alpha}\rfloor}\widetilde{{\mathbb{P}}}_{n}(k_{i}>\sqrt{n},\,k_{i-1}\leq\sqrt{n},\,i\leq S)}.$
Conditionally on $h_{1},\cdots,h_{S}$ and the event $\{m_{*}>\frac{c_{2}n}{(\log n)^{2\alpha}}\}$, we have for all $n$ large enough and $\sqrt{n}\leq l\leq\sqrt{n}(1+\frac{2}{(\log n)^{\alpha}})$, on the event $\{S\geq i,\,k_{i-1}\leq\sqrt{n}\}$, $k_{i}=\ell$ if and only if $\nu^{h_{i}}_{l}=1$ and there are $k_{i-1}-1$ of $(\nu^{h_{i+1}}_{j}:1\leq j\leq l-1)$ taking the value 1. We deduce that on the event $\{S\geq i,\,k_{i-1}\leq\sqrt{n}\}$,
$\widetilde{{\mathbb{P}}}_{n}(k_{i}=l\,|\,k_{i-1},\,i\leq S)=\frac{\binom{l-1}{k_{i-1}-1}\binom{n-l}{n-k_{i-1}-r_{n}}}{\binom{n}{n-r_{n}}}=:b_{l}.$
It follows that
$\frac{b_{l+1}}{b_{l}}=\frac{l}{l-k_{i-1}+1}\cdot\frac{k_{i-1}+r_{n}-l}{n-l},$
which is decreasing on $l$. Since $k_{i-1}\leq\sqrt{n}$, it follows that for all $l\geq l_{n}:=\sqrt{n}(1+\frac{3}{2(\log n)^{\alpha}})$, $\frac{b_{l+1}}{b_{l}}\leq\frac{2}{3}+o_{n}(1)$ and hence
$\widetilde{{\mathbb{P}}}_{n}\bigg(k_{i}\geq\sqrt{n}\Big(1+\frac{2}{(\log n)^{\alpha}}\Big)\,\Big|\,k_{i-1},\,i\leq S\bigg)\leq b_{l_{n}}\sum_{l\geq\lfloor\sqrt{n}(1+\frac{2}{(\log n)^{\alpha}})\rfloor}(\frac{4}{5})^{l-l_{n}}.$
Hence for all $1\leq i\leq S$,
$\frac{\widetilde{{\mathbb{P}}}_{n}\big(k_{i}\geq\sqrt{n}(1+\frac{2}{(\log n)^{\alpha}})\,|\,k_{i-1},\,i\leq S\big)}{\widetilde{{\mathbb{P}}}_{n}(k_{i}\geq\sqrt{n}\,|\,k_{i-1},\,i\leq S)}\leq\sum_{i\geq\frac{\sqrt{n}}{2(\log n)^{\alpha}}}(\frac{4}{5})^{i}\leq\exp\bigg(-\frac{c_{3}\sqrt{n}}{(\log n)^{2\alpha}}\bigg).$
Going back to (92) and conditioning on $k_{i-1}$ and $i<S$ in each part of the sum, we deduce (91).
Step 4. On the event $\{m_{*}>\frac{c_{2}n}{(\log n)^{2\alpha}},\,S<3(\log n)^{2\alpha}\}$, if $i_{*}=\infty$, then $k_{S}\leq\sqrt{n}<m_{*}$ for all $1\leq i\leq S$ for $n$ large enough. On the other hand, by (90), we deduce that
$\widetilde{{\mathbb{P}}}_{n}\bigg(k_{i+1}\leq k_{i}\Big(1+\frac{2}{(\log n)^{\alpha}}\Big)\,\bigg|\,S>i,\,\sqrt{n}<k_{i}<\frac{c_{2}n}{2(\log n)^{2\alpha}}\bigg)\geq 1-\exp\Big(-\frac{c_{3}\sqrt{n}}{(\log n)^{2\alpha}}\Big).$
We observe that for any $\eta>0$, $(1+\frac{2}{(\log n)^{\alpha}})^{3(\log n)^{2\alpha}}$ is of order $\exp(6(\log n)^{\alpha})$ which is smaller than $n^{\eta}$ for all $n$ sufficiently large, thanks to the assumption $\alpha<1$. It follows that if $i_{*}<S<3(\log n)^{2\alpha},\,k_{i_{*}}\leq\sqrt{n}(1+\frac{2}{(\log n)^{\alpha}})$ and $k_{j+1}\leq k_{j}(1+\frac{2}{(\log n)^{\alpha}})$ for all $i_{*}\leq j\leq i-1$, then $k_{i}\leq\sqrt{n}(1+\frac{2}{(\log n)^{\alpha}})^{i+1-i_{*}}$ which is of course smaller than $\frac{c_{2}n}{2(\log n)^{2\alpha}}$ for all $n$ sufficiently large. Let $A_{i}:=\{k_{i}\leq\sqrt{n}(1+\frac{2}{(\log n)^{\alpha}})^{i+1-i_{*}}\}$. Then we derive that
$\widetilde{{\mathbb{P}}}_{n}\bigg(\exists\,i_{*}\leq i<S,\,k_{i+1}>k_{i}\Big(1+\frac{2}{(\log n)^{\alpha}}\Big)\,\bigg|\,i_{*}<S,\,k_{i_{*}}\leq\sqrt{n}\Big(1+\frac{2}{(\log n)^{\alpha}}\Big)\bigg)$
$\leq\,$ $\widetilde{{\mathbb{E}}}_{n}\bigg[\sum_{j=i_{*}}^{S-1}\widetilde{{\mathbb{P}}}_{n}\bigg(k_{j+1}>k_{j}\Big(1+\frac{2}{(\log n)^{\alpha}}\Big),A_{j}\,\bigg|\,i_{*}<S,\,k_{i_{*}}\leq\sqrt{n}\Big(1+\frac{2}{(\log n)^{\alpha}}\Big)\bigg)\bigg]$
$\leq\,$ $\sum_{i=1}^{\lfloor 3(\log n)^{2\alpha}\rfloor-1}\widetilde{{\mathbb{P}}}_{n}\bigg(k_{i+1}>k_{i}\Big(1+\frac{2}{(\log n)^{\alpha}}\Big)\,\bigg|\,S>i,\,\sqrt{n}<k_{i}<\frac{c_{2}n}{2(\log n)^{2\alpha}}\bigg)$
$\leq\,$ $[3(\log n)^{2\alpha}-1]\exp\Big(-\frac{c_{3}\sqrt{n}}{(\log n)^{2\alpha}}\Big),$
where the third inequality follows since $(\nu^{s}:s\geq h_{i})$ is independent of $(\nu^{s}:1\leq s\leq h_{i}-1)$ for all $i\geq i_{*}$ conditionally on $S,h_{1},\cdots,h_{S},i_{*}$. But if $i_{*}<S,\,k_{i_{*}}\leq\sqrt{n}\Big(1+\frac{2}{(\log n)^{\alpha}}\Big)$ and $k_{i+1}\leq k_{i}\Big(1+\frac{2}{(\log n)^{\alpha}}\Big)$ for all $i_{*}\leq i<S$, then $A_{S}$ occurs, meaning $k_{S}\leq\sqrt{n}(1+\frac{2}{(\log n)^{\alpha}})^{S+1-i_{*}}<m_{*}$. We conclude that
$\widetilde{{\mathbb{P}}}_{n}(k_{S}<m_{*})$
$\geq\$$\widetilde{{\mathbb{P}}}_{n}(i_{*}=\infty)+\widetilde{{\mathbb{P}}}_{n}\bigg(i_{*}=S,\,k_{i_{*}}\leq\sqrt{n}\Big(1+\frac{2}{(\log n)^{\alpha}}\Big)\bigg)$
$+\widetilde{{\mathbb{P}}}_{n}\bigg(i_{*}<S,\,k_{i_{*}}\leq\sqrt{n}\Big(1+\frac{2}{(\log n)^{\alpha}}\Big),\,\forall\,i_{*}\leq i<S,\,k_{i+1}\leq k_{i}\Big(1+\frac{2}{(\log n)^{\alpha}}\Big)\bigg)$
$\geq\$$1-3(\log n)^{2\alpha}\exp\Big(-\frac{c_{3}\sqrt{n}}{(\log n)^{2\alpha}}\Big).$
It together with (89) implies that for all $k\geq 1$ and $0<\varepsilon<1$, ${\mathbb{P}}(k_{S}<\infty)>1-\varepsilon$ for all $n$ large enough. We complete the proof.